### v2

# MD 20742, USA

Abstract.
We study heavy tailed random variables obtained by sampling an observable with a power singularity along an orbit of an expanding map on the circle. In the regime where the observables have infinite variance, we show that the set of limit points of the ergodic sum process, rescaled by $N^{1/s}(\operatorname{ln}N)^{\alpha}$, where $s$ is the index of limiting stable law, equals to the set of piecewise constant increasing functions with finitely many jumps. The proof relies on the multiple Borel-Cantelli Lemma proven in Dolgopyat, Fayad and Liu [J. Mod. Dyn. 18 (2022) 209–289].
1. Introduction
Let $X_{k}$ be a sequence of independent and identically distributed random variables with mean 0 and variance 1. Let $S_{n}=\sum_{k=1}^{n}X_{k}.$ Consider a function $W_{N}:[0,1]\to\mathbb{R}$ defined by $W_{N}\left(\frac{n}{N}\right)=\frac{S_{n}}{\sqrt{N}}$ at rational points $\frac{n}{N}$ and by linear interpolation in between. The functional central limit theorem states that the distribution of $W_{N}(\cdot)$ converges to the standard Wiener measure as $N\to\infty$. Since Wiener measure has full support, it follows (see e.g. [41]) that with probability $1$ for any continuous function $h$ on $[0,1]$, there is a sequence $N_{j}$ such that $W_{N_{j}}$ converges uniformly to $h.$
It is natural to ask whether one can obtain a smaller set of limits by considering a larger normalization (so that the renormalized function is close to 00 most of the time). Strassen [42] answered this question by showing that if one takes
${\widetilde{W}}_{N}\left(\frac{n}{N}\right)=\frac{S_{n}}{\sqrt{2N\operatorname{ln}\operatorname{ln}N}},$
then with probability $1$, $h$ is a limit point of $\{{\widetilde{W}}_{N}\}$ if and only if $h$ is an absolutely continuous function such that $h(0)=0$ and $\int_{0}^{1}(h^{\prime}(x))^{2}{\rm{d}}x\leq 1.$ This result is referred to as the functional form of law of the iterated logarithm (FLIL) in the literature [16, 39]. FLIL is a generalization of Khintchine-Kolmogoroff’s original form [28, 29] of the law of the iterated logarithm (LIL) and Hartman–Wintner’s LIL [17] for random walks. The method used in [42] for proving FLIL is known as Strassen’s almost sure invariance principle (ASIP). ASIP was further investigated for martingales (see e.g. [16, 20, 43]), which states that a martingale satisfying certain moment conditions is with probability $1$ close to a Brownian motion. Consequently, FLIL holding for Brownian motion is still valid for martingales.
ASIP can be extended to dynamical systems [39] by approximating weakly dependent random variables by martingales. This method has been applied to show ASIP for various dynamical systems (see e.g. [7, 8, 13, 15, 18, 19, 21, 30, 31, 33, 36, 38, 40]). Specifically, these systems include uniformly hyperbolic systems [10], Lorenz attractors [22], non-uniformly hyperbolic systems [32, 34, 35], partially hyperbolic systems [11], dispersing billiards [3, 5] and compact group extensions [14]. However, there are still many examples of dynamical systems where the limit distributions of ergodic sums are non-Gaussian. For these systems, the above approach is not applicable and ASIP does not hold. It is therefore of interest to develop a new method for finding a suitable normalizing sequence $L_{N}$ so that the function ${\tilde{S}}_{N}$ defined by ${\tilde{S}}_{N}(\frac{n}{N})=\frac{S_{n}}{L_{N}}$ has a non-trivial set of limit points.
In this paper, we address the above problem in a simple setting. Let $S_{n}(y)=\sum_{k=1}^{n}\phi(f^{k}y)$ where $f$ is an expanding map of the circle $\mathbb{S}$ and $\phi$ has a power singularity of the type $|y-x|^{-1/s}$ for fixed $x\in\mathbb{S}.$ Assuming that $0<s<2,$ $s\neq 1$ and $x$ is a Lebesgue typical point, we show that a reasonable choice of $L_{N}$ is $L_{N}=N^{1/s}(\operatorname{ln}N)^{\alpha}.$ If $\alpha$ satisfies ${[(r+1)s]}^{-1}<\alpha\leq{(rs)}^{-1},$ then the set of limit points consists of piecewise constant increasing functions starting at 00 with at most $r$ jumps.
Before formulating our result precisely, let us mention two related topics.
The first topic is Laws of Large Numbers for trimmed sums (see [1, 23, 24, 25, 27, 37]). Trimming is an important tool in our analysis since it allows us to discard all the summands in $S_{N}$ except for $r$ largest ones. However, this is only part of our analysis, since the contribution of the retained terms needs to be analyzed in more detail than in the papers mentioned above. We also note that unlike the above mentioned papers, our results apply not only to sums with infinite mean but also to the sums with infinite variance.
The second topic is the application of the classical Borel-Cantelli Lemma to estimate the rate of convergence in the Law of Large Numbers (see [4, 6, 9, 26] and references wherein). This topic is closer to the subject of our paper. However, the papers mentioned above deal with a single $N$ at a time, while we obtain the functional version of the deviation result. Fundamental to our approach is the multiple Borel-Cantelli Lemma proposed recently in [12], which provides a criterion to determine when several rare events occur at the same time scale.
In Section 2, we present the main theorem and sketch its proof using the multiple Borel-Cantelli Lemma. In Section 3, we focus on a key lemma which allows us to control the trimmed sums. The proof in Section 3 uses the reflection principle for the processes under consideration. For completeness, we include some standard proofs as appendices.
2. An analogue of LIL for Smooth expanding maps
2.1. The main result.
Let $f$ be a $C^{2}$ expanding map of the circle $\mathbb{S}.$ Here we say $f$ is expanding if there exists $\lambda>1$ such that $|f^{\prime}(x)|>\lambda$ for any $x\in\mathbb{S}.$ Then there is a unique $f$-invariant measure $\mu$ that is equivalent to Lebesgue measure [44]. Fix $x,$ and let $\phi$ be an unbounded function of the form
$\phi(y)=a|y-x|^{-1/s}+\psi(y),$
where $\psi$ is a $C^{1}$ function and $s>0.$ Let
$X_{n}(y)=\phi(f^{n}y).$
Denote $\|\cdot\|_{\infty}$ the uniform norm and $L=\|\psi\|_{\infty}.$ Then $\mu(X_{n}>t)=\mu(\phi(y)>t)$ and
(2.1) $\mu\left(|y-x|<\left(\frac{a}{t+L}\right)^{s}\right)\leq\mu(\phi(y)>t)\leq\mu\left(|y-x|\leq\left(\frac{a}{t-L}\right)^{s}\right)$
for $t>L.$ It follows that
(2.2) $\mu(X_{n}>t)\sim a_{s}t^{-s}$
for sufficiently large $t$ and $a_{s}=\rho_{0}a^{s}$, where $\rho_{0}$ is the density of $\mu$ at $x.$ For a positive $\alpha$, set
$W_{N}(t)=\frac{\left(\sum_{n\leq Nt}X_{n}\right)-a_{N}}{N^{1/s}(\operatorname{ln}N)^{\alpha}},$
where $a_{N}=0$ if $s\leq 1$ and $a_{N}=Nt\int\phi(y){\rm{d}}\mu(y)$ if $s>1.$
We regard $W_{N}(\cdot)$ as an element of the space ${\mathbb{D}}[0,1]$ of functions from $[0,1]$ to [,+][$\infty$,$\infty$][ - ∞ , + ∞ ] that are right continuous and have left limits, where we endow [,+][$\infty$,$\infty$][ - ∞ , + ∞ ] with the distance $d(z_{1},z_{2})=\left|\tan^{-1}(z_{1})-\tan^{-1}(z_{2})\right|.$ There are several possible topologies on ${\mathbb{D}}[0,1]$, the most relevant for us is Skorokhod’s $J_{1}$ topology. Let $\Lambda$ be the space of order preserving homeomorphisms on $[0,1].$ Then $J_{1}$ topology is defined using the distance
$d_{J_{1}}(h_{1},h_{2})=\inf_{\lambda\in\Lambda}\max\left\{\|\lambda-Id\|_{\infty},\,\max_{t\in[0,1]}d(h_{1}(t),h_{2}(\lambda(t)))\right\}.$
Two functions are $J_{1}$ close, if their graphs are close with respect to the Hausdorff distance.
In our analysis we shall use several facts about $J_{1}$ topology. First, if two functions are close with respect to the supremum norm, then they are $J_{1}$ close. Second, if one takes ${\mathbb{H}}_{r}$ to be the space of step functions with (at most) $r$ jumps, then $W\in{\mathbb{H}}_{r}$ can be written in the form
$W(t)=\sum_{m=1}^{\bar{r}}c_{m}{\mathbf{1}}_{t\leq t_{m}}(t),$
where $\bar{r}\leq r$ and $\{c_{m}\}_{m=1}^{r}$ and $\{t_{m}\}_{m=1}^{r}$ are $\bar{r}$–tuples such that $c_{m}$’s are positive and $t_{m}$’s are increasing. Two elements $W_{1}$ and $W_{2}$ of ${\mathbb{H}}_{r}$ with exactly $r$ jumps are $J_{1}$ close provided that $\{c_{m}^{1}\}_{m=1}^{r}$ is close to $\{c_{m}^{2}\}_{m=1}^{r}$ and $\{t_{m}^{1}\}_{m=1}^{r}$ is close to $\{t_{m}^{2}\}_{m=1}^{r}$. Third, $J_{1}$ topology allows us to ignore small jumps.
We also need the following definition. We call $x$ Diophantine if there exists $\epsilon=\epsilon(x)>0$ and $\rho_{0}=\rho_{0}(x)>0$ such that for all positive $\rho\leq\rho_{0}$ and all positive integer $k\leq\epsilon|\operatorname{ln}\rho|,$ $f^{-k}B(x,\rho)\cap B(x,\rho)=\emptyset.$ Due to Lemma 5 in [2], $\mu-$almost every $x$ is Diophantine.
Theorem 2.1.
Suppose that either
(i) $10<s<$ 1 and $x$ is non-periodic or
(ii) $21\leq s<$ 2 and $x$ is Diophantine.
Let $\alpha$ be such that $\frac{1}{(r+1)s}<\alpha\leq\frac{1}{rs}$ for some $r\in{\mathbb{N}}.$ Then for almost every $y$ the following holds: the set of $J_{1}$ limit points of $\{W_{N}(t)\}_{N=1}^{\infty}$ is equal to ${\mathbb{H}}_{r}.$
2.2. Multiple Borel-Cantelli Lemma.
The proof of Theorem 2.1 requires control on the number of times $k\leq n$ such that the events $\left\{y:\,c_{j}-\varepsilon\leq X_{k}(y)\leq c_{j}+\varepsilon\right\}$ occur for some $c_{j}>0$ and $0<\varepsilon<c_{j}.$ If several such events occur for some $n$ we also need to control the occurrence time of these events. To this end, we shall use a generalization of the multiple Borel-Cantelli Lemma proven in [12], those statement is recalled in this subsection.
Consider a sequence $\{\rho_{n}\}_{n=1}^{\infty}$ and a family of events $\{\Omega_{\rho_{n}}\}_{n=1}^{\infty}.$ Suppose that there exist ${\hat{\varepsilon}}>0$ and a decomposition of $\Omega_{\rho_{n}}$ for each $\rho_{n}$ such that
$\Omega_{\rho_{n}}=\bigcup_{i=1}^{p}\Omega_{\rho_{n},i}\quad\text{and}\quad\mu(\Omega_{\rho_{n},i})\geq{\hat{\varepsilon}}\mu(\Omega_{\rho_{n}}).$
We denote by
(2.3) $\Omega^{k}_{\rho_{n}}=f^{-k}(\Omega_{\rho_{n}}),\,\,\,\Omega^{k}_{\rho_{n},i}=f^{-k}(\Omega_{\rho_{n},i}),\,\,\,\sigma_{\rho_{n}}=\mu(\Omega_{\rho_{n}})\,\,\,{\text{and}}\,\,\,\sigma_{\rho_{n},i}=\mu(\Omega_{\rho_{n},i}).$
Fix $r\in\mathbb{N^{*}},$ for each $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r},$ we will characterize whether with probability 1, for infinitely many $n,$ the events $\Omega_{\rho_{n},i_{j}}^{k_{j}}$ happen for all $1\leq j\leq r.$
For an arbitrary $r$-tuple $0<k_{1}<\dots<k_{r}\leq n,$ we take the separation indices
$\operatorname{Sep}_{n}(k_{1},\dots,k_{r})=\text{Card}\left\{j\in\{0,\dots r-1\}:k_{j+1}-k_{j}\geq s(n)\right\},\quad k_{0}:=0;$
$\widehat{{\rm Sep}}_{n}(k_{1},\dots,k_{r})=\text{Card}\left\{j\in\{0,\dots r-1\}:k_{j+1}-k_{j}\geq\hat{s}(n)\right\},\quad k_{0}:=0$
for functions $s:{\mathbb{N}}\rightarrow{\mathbb{N}},$ $s(n)\leq(\operatorname{ln}n)^{2}$ and $\hat{s}:{\mathbb{N}}\rightarrow{\mathbb{N}},$ $\varepsilon n\leq\hat{s}(n)<n(1-q)/{(2r)}$ for some $0<q<1,$ $0<\varepsilon<(1-q)/{(2r)},$ respectively.
Suppose the function $\sigma.$ and $\sigma_{\cdot,i}$ defined in (2.3) are increasing and there is a sequence ${\varepsilon}_{n}\to 0$ satisfying some of the mixing conditions below.
  • $(M1)_{r}$
    If $0<k_{1}<k_{2}<\dots<k_{r}\leq n$ are such that $\operatorname{Sep}_{n}(k_{1},\dots,k_{r})=r$ and for each $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r},$
    $\left(\prod_{j=1}^{r}\sigma_{\rho_{n},i_{j}}\right)(1-{\varepsilon}_{n})\leq\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{n},i_{j}}\right)\leq\left(\prod_{j=1}^{r}\sigma_{\rho_{n},i_{j}}\right)(1+{\varepsilon}_{n});$
  • $(M2)_{r}$
    There exists $K>0$ such that if $0<k_{1}<k_{2}<\cdots<k_{r}\leq n$ are such that $\operatorname{Sep}_{n}(k_{1},\dots,k_{r})=m<r$, then
    $\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{n}}\right)\leq{\frac{K\sigma_{{\rho}_{n}}^{m}}{(\operatorname{ln}n)^{100r}}};$
  • $(M3)_{r}$
    If $0<k_{1}<\dots<k_{r}<l_{1}<\dots<l_{r}$ are such that $2^{i}<k_{\alpha}\leq 2^{i+1},2^{j}<l_{\beta}\leq 2^{j+1}$ for $1\leq\alpha,\beta\leq r$, $j-i\geq b$ for some constant $b\geq 2,$
    $\widehat{{\rm Sep}}_{2^{i+1}}(k_{1},\ldots,k_{r})=r,\quad\widehat{{\rm Sep}}_{2^{j+1}}(l_{1},\ldots,l_{r})=r,\quad l_{1}-k_{r}\geq\hat{s}(2^{j+1})$
    and for each $(i_{1},i_{2},\dots,i_{r}),\,(j_{1},j_{2},\dots,j_{r})\in\{1,\dots,p\}^{r},$
    $\mu\left(\left[\bigcap_{\alpha=1}^{r}\Omega^{k_{\alpha}}_{{\rho}_{2^{i}},i_{\alpha}}\right]\bigcap\left[\bigcap_{\beta=1}^{r}\Omega^{l_{\beta}}_{{\rho}_{2^{j}},j_{\beta}}\right]\right)\leq\left(\prod_{\alpha=1}^{r}\sigma_{\rho_{2^{i}},i_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\sigma_{\rho_{2^{j}},j_{\beta}}\right)(1+{\varepsilon}_{i}).$
We consider the events
$A_{*}^{n}=\{\exists\,0<k_{1}<k_{2}<\dots<k_{r}\leq n,\,\,\operatorname{Sep}(k_{1},\dots k_{r})=r,\,\,\,\Omega^{k_{j}}_{{\rho}_{n}}\text{ happens}\}$
and
$A_{I_{1},\dots,I_{r},i_{1},\dots i_{r}}^{n}=\{\exists\,k_{j},\,\,k_{j}/n\in I_{j},\,\Omega^{k_{j}}_{{\rho}_{n},i_{j}}\text{ happens}\}.$
We also let
$(M4)_{*}\quad\quad\lim_{n\to\infty}\frac{\mu(A_{*}^{n})}{n^{r}\sigma_{\rho_{n}}^{r}}=1.$
$(M4)_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}\quad\quad\lim_{n\to\infty}\frac{\mu(A_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}^{n})}{n^{r}\Pi_{j=1}^{r}\left(\sigma_{\rho_{n},i_{j}}|I_{j}|\right)}=1.$
Lemma 2.2.
Given a family of events $\{\Omega_{\rho_{n}}\}_{\rho_{n}>0}$ and their decomposition $\Omega_{\rho_{n}}=\bigcup_{i=1}^{p}\Omega_{\rho_{n},i}$, define
${\mathbf{S}}_{r}=\sum_{j=1}^{\infty}\left(2^{j}{\sigma_{{\rho}_{2^{j}}}}\right)^{r}.$
We assume that $\rho_{n_{1}}\leq\rho_{n_{2}}$ for $n_{1}\geq n_{2}.$
Suppose ${\mathbf{S}}_{r}<\infty$ and $\Omega_{{\rho}_{n_{1}}}\subset\Omega_{{\rho}_{n_{2}}}$ for $n_{1}\geq n_{2}:$
(a) If $(M1)_{r}$ and $(M2)_{r}$ are satisfied, then with probability $1$, we have that for large $n,$ the events $\Omega_{\rho_{n}}^{k}$ happen for at most $(r-1)$ times.
(b) If $(M4)_{*}$ holds, then with probability $1$, $A^{n}_{*}$ does not occur for large $n.$
Suppose ${\mathbf{S}}_{r}=\infty.$ For any $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r}$ and any intervals $I_{1},I_{2}\dots I_{r}\subset[0,1]:$
If
(c) $(M1)_{k}$, $(M2)_{k}$ for $k=r+1,\dots,2r$ as well as $(M3)_{r}$ are satisfied and $\hat{s}(n)/n<\min_{0\leq i,\,j\leq r,\,\,i\neq j}d(I_{i},I_{j}),$ where $I_{0}=0;$ or
(d) $(M1)_{r},$ $(M3)_{r}$ are satisfied and $(M4)_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}$ holds for $\hat{s}(n)/n<\min_{0\leq i,\,j\leq r,\,\,i\neq j}d(I_{i},I_{j}),$ where $I_{0}=0,$ then with probability $1$, there are infinitely many $n$ such that for some $k_{1}(n),k_{2}(n)\dots k_{r}(n)$ with $k_{j}(n)/n\in I_{j},$ the events $\Omega^{k_{j}}_{\rho_{n},i_{j}}$ occur.
Proof.
Part (a) is Theorem 2.6 in [12]. We consider the case ${\mathbf{S}}_{r}<\infty$ and $(M4)_{*}$ holds. Suppose $\{\delta_{n}\}$ is a sequence such that $\rho_{n}=\delta_{2n}$ and
$A_{*,2}^{n}=\{\exists\,0<k_{1}<k_{2}<\dots<k_{r}\leq 2n,\,\,\operatorname{Sep}(k_{1},\dots k_{r})=r,\,\,\,\Omega^{k_{j}}_{{\rho}_{n}}\text{ happens}\}.$
We obtain that
$\lim_{n\to\infty}\frac{\mu(A_{*,2}^{n})}{(2n)^{r}\sigma_{\delta_{2n}}^{r}}=1$
using $(M4)_{*}.$ Since $\sigma_{\rho_{n}}=\mu(\Omega_{\rho_{n}})=\mu(\Omega_{\delta_{2n}})=\sigma_{\delta_{2n}},$
$\lim_{n\to\infty}\frac{\mu(A_{*,2}^{n})}{n^{r}\sigma_{\rho_{n}}^{r}}=2^{r}.$
By taking $n=2^{m},$ we have that $\mu(A_{*,2}^{2^{m}})=O\left(2^{rm}{\sigma_{{\rho}_{2^{m}}}^{r}}\right)$ and
$\sum_{m=1}^{\infty}\mu(A_{*,2}^{2^{m}})=O\left({\mathbf{S}}_{r}\right)<\infty.$
It implies that $A_{*,2}^{2^{m}}$ does not occurs for infinitely many $m$’s with probability 1 using the classical Borel-Cantelli Lemma. Then Part (b) follows from the fact that $A_{*}^{n}\subset A_{*,2}^{2^{m}}$ when $2^{m}\leq n\leq 2^{m+1}.$
Part (c) is Theorem 2.12 in [12]. Finally, part (d) holds because in the proof of part (c) in [12], $(M1)_{k}$ and $(M2)_{k}$ for $k=r+1,\dots,2r$ are only used to verify $(M4)_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}$. $\square$
We now turn to this paper’s problem to consider the following targets:
$\Omega_{\rho_{N}}=\{y:\phi(y)>\rho_{N}\}$
and
$\Omega_{\rho_{N},j}=\left\{y:\frac{\phi(y)}{\rho_{N}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}]\right\}$
for $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{t},$ $t\in\mathbb{R}$ and $c_{j}>{\varepsilon}$ for any $j.$ Take $s(N)=K\operatorname{ln}N$ with $K$ being sufficiently large and $\hat{s}(N)=2{\varepsilon}N.$
Proposition 2.3.
(a) For any $r\in\mathbb{N},$ $\Omega_{\rho_{N}}$ and $\Omega_{\rho_{N},j}$ satisfy $(M1)_{r}$ and $(M3)_{r}.$
(b) If $x$ is Diophantine then $(M2)_{r}$ holds.
(c) If $x$ is non-periodic then $(M4)_{*}$ and $(M4)_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}$ are satisfied.
The proof of Proposition 2.3 is similar to Proposition 3.9 in [12] with some modifications, which is included in Appendix B.
Next, we will recall the exponential mixing property for expanding maps, which is useful not only in the proof Proposition 2.3 but also in the next section. Denote by $BV$ the space of functions with bounded variation on $\mathbb{S}$ and
${\mathbb{E}}\psi=\int\psi{\rm{d}}\mu\,\,\text{for}\,\,\psi\in BV.$
By Proposition 3.8 of [44], there are constants $C,$ $\theta<1$ such that
(2.4) $\left|{\mathbb{E}}\left(\psi_{1}(\psi_{2}\circ f^{n})\right)-({\mathbb{E}}\psi_{1})(E\psi_{2})\right|\leq C\|\psi_{1}\|_{BV}\|\psi_{2}\|_{L^{1}}\theta^{n},\,n\in\mathbb{N}$
for any $\psi_{1},\,\psi_{2}\in BV,$ where $\|\psi\|_{BV}:=\|\psi\|_{L^{1}}+\text{Var}(\psi)$ and $\text{Var}(\psi)$ is the variation of $\psi.$
In the following, we use $C$ to denote a constant that may change from line to line but that will not depend on $N.$
Lemma 2.4.
[…]
  1. (i)
    If $\varphi_{j}\geq 0$ for $1\leq j\leq q,$ we have for any $k_{1}\leq k_{2}\leq\cdots\leq k_{q},$ $k_{i}\in\mathbb{N},$
    (2.5) $\left(\prod_{j=1}^{q-1}\left(\|\varphi_{j}\|_{L^{1}}-C\|\varphi_{j}\|_{BV}\theta^{k_{j+1}-k_{j}}\right)\right)\|\varphi_{q}\|_{L^{1}}\leq\left\|\prod_{j=1}^{q}\varphi_{j}\circ f^{k_{j}}\right\|_{L^{1}}$
    $\leq\left(\prod_{j=1}^{q-1}\left(\|\varphi_{j}\|_{L^{1}}+C\|\varphi_{j}\|_{BV}\theta^{k_{j+1}-k_{j}}\right)\right)\|\varphi_{q}\|_{L^{1}}$
  2. (ii)
    If for each $j,$ $k_{j+1}-k_{j}\geq k,$ then
    (2.6) $\left|{\mathbb{E}}\left(\prod_{j=1}^{q}(\varphi_{j}\circ f^{k_{j}})\right)-\prod_{j=1}^{q}{\mathbb{E}}\varphi_{j}\right|\leq C\theta^{k}\prod_{j=1}^{q}\|\varphi_{j}\|_{BV}.$
Lemma 2.4 could be proved by the arguments of Proposition 6.2 in [12] with $\|\cdot\|_{Lip}$ replaced by $\|\cdot\|_{BV}.$ However, in the present case a simpler argument is available and we provide it in Appendix A for the convenience of the reader.
2.3. Proof of Theorem 2.1
Fix a small $\delta>0$ and let
$S_{n}^{\prime}=\sum_{1\leq k\leq n}\left(X_{k}{\mathbf{1}}_{\{X_{k}\geq N^{1/s}(\operatorname{ln}N)^{\delta}\}}-b_{N}^{\prime}\right),\quad S_{n}^{\prime\prime}=\sum_{1\leq k\leq n}\left(X_{k}{\mathbf{1}}_{\{X_{k}<N^{1/s}(\operatorname{ln}N)^{\delta}\}}-b_{N}^{\prime\prime}\right),$
where $b_{N}^{\prime}=b_{N}^{\prime\prime}=0$ if $s<1$ and $b_{N}^{\prime},b_{N}^{\prime\prime}$ are chosen to make the expectations of the LHS’s zero if $s>1.$
Lemma 2.5.
If $\delta<\alpha$ and either $10<s<$ 1 or $21<s<$ 2 and $x$ is Diophanitne then
[…]
The proof of Lemma 2.5 is the main part of this paper and is contained in the next section. Given Lemma 2.5, we proceed to show Theorem 2.1.
Let
$W_{N}^{\prime}(t)=\frac{S_{Nt}^{\prime}}{N^{1/s}(\operatorname{ln}N)^{\alpha}}.$
By Lemma 2.5, the limit points of $W_{N}(\cdot)$ and $W_{N}^{\prime}(\cdot)$ are the same. Consider the targets $\Omega_{\rho_{N}}=\{y:\phi(y)>\rho_{N}\}$ for $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{\delta}$ and suppose that $x$ is non-periodic. By Proposition 2.3 and Lemma 2.2(b), for large $N$ there are at most $\left[(\delta s)^{-1}\right]$ intervals $I_{j}$ of length $2s(N)$ such that all numbers $n$ such that $X_{n}>N^{1/s}(\operatorname{ln}N)^{\delta}$ are contained in one of those intervals. Moreover taking ${\bar{\alpha}}$ such that $((r+1)s)^{-1}<{\bar{\alpha}}<\alpha,$ we see that among those intervals there are at most $r$ which contain $n^{\prime}$s with $X_{n}\geq N^{1/s}(\operatorname{ln}N)^{{\bar{\alpha}}}.$ We claim that the contribution of the remaining intervals is negligible for a.e. x. Indeed suppose there is an interval $I$ of length $2s(N)$ and ${\bar{n}}\in I$ such that
(2.7) $\max_{n\in I}X_{n}=X_{\bar{n}}\leq N^{1/s}(\operatorname{ln}N)^{\bar{\alpha}}\;\;\mathrm{and}\;\;\sum_{n\in I}X_{n}\geq{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}.$
Let
$I^{\prime}=\{n\in I:|n-{\bar{n}}|\leq(\operatorname{ln}N)^{(\alpha-{\bar{\alpha}})/2}\},\quad I^{\prime\prime}=I\setminus I^{\prime},\quad Y_{I}=\max_{n\in I^{\prime\prime}}X_{n}.$
Then
$\sum_{n\in I}X_{n}\leq\sum_{n\in I^{\prime}}X_{n}+\sum_{n\in I^{\prime\prime}}X_{n}\leq N^{1/s}(\operatorname{ln}N)^{(\alpha+{\bar{\alpha}})/2}+2Y_{I}s(N).$
By the definition of $I$ we have that $Y_{I}\geq{\bar{\varepsilon}}N^{1/s}(\operatorname{ln}N)^{\alpha-1}$ for some ${\bar{\varepsilon}}={\bar{\varepsilon}}({\varepsilon})>0.$ Accordingly if the interval $I$ satisfying (2.7), then there are two moments ${\bar{n}},{\hat{n}}\in[1,N]$ such that
(2.8) $(\operatorname{ln}N)^{(\alpha-{\bar{\alpha}})/2}\leq|{\bar{n}}-{\hat{n}}|\leq 2s(N),\quad\min\left\{X_{\bar{n}},\,X_{\hat{n}}\right\}>{\bar{\varepsilon}}N^{1/s}(\operatorname{ln}N)^{\alpha-1}.$
Thus the lemma below shows that the contribution of the intervals satisfying (2.7) is negligible. Its proof is available in Appendix B.
Lemma 2.6.
For a.e. $y$, (2.8) happens only for finitely many $N.$
It then follows that the limit points of $W_{N}^{\prime}(\cdot)$ have at most $r$ positive jumps. That is, all limit points are in ${\mathbb{H}}_{r}.$
To see that all points in ${\mathbb{H}}_{r}$ are realized, fix $r$-tuples $(c_{1},\dots,c_{r})$ and $(t_{1},\dots t_{r})$ such that $c_{j}>0$ for all $j$ and fix $0<{\varepsilon}<c_{j}.$ Consider the events
$\Omega_{\rho_{N},j}=\left\{y:\frac{\phi(y)}{\rho_{N}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}]\right\}.$
Taking $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{\alpha}$ and applying Proposition 2.3 and Lemma 2.2(d), we see that with probability $1$ there is a subsequence $N_{k}$ and times $n_{1}(k),n_{2}(k),\dots,n_{r}(k)$ such that
$\frac{n_{j}(k)}{N_{k}}\in[t_{j}-{\varepsilon},t_{j}+{\varepsilon}]\quad\text{and}\quad\frac{X_{n_{j}(k)}(y)}{N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}].$
It follows that for each $W\in{\mathbb{H}}_{r}$, $W$ belongs to the limit set with probability $1$ by taking a countable set of ${\varepsilon}_{l}$ converging to $0$ Recall the fact that two elements $W_{1}$ and $W_{2}$ of ${\mathbb{H}}_{r}$ with exactly $r$ jumps are $J_{1}$ close if and only if $\{c_{m}^{1}\}_{m=1}^{r}$ is close to $\{c_{m}^{2}\}_{m=1}^{r}$ and $\{t_{m}^{1}\}_{m=1}^{r}$ is close to $\{t_{m}^{2}\}_{m=1}^{r}$. Taking a countable dense subset in ${\mathbb{H}}_{r}$ completes the proof.
3. Proof of Lemma 2.5.
3.1. Splitting the sum
The proof of Lemma 2.5 will be given in the next three subsections. Fix a large number $D$ satisfying
(3.1) $D(1-s)+\alpha>1$ $\quad\text{if}\quad s<1;$
(3.2) $D(2-s)+2\alpha>3$ $\quad\text{if}\quad 1\leq s<2.$
and split $S_{N}^{\prime\prime}=\bar{S}_{N}+\bar{\bar{S}}_{N}$ where $\bar{S}_{n}=\sum_{k=1}^{n}\bar{X}_{k}$ and $\bar{\bar{S}}_{n}=\sum_{k=1}^{n}\bar{\bar{X}}_{k}$ for
$\bar{X}_{k}=X_{k}{\mathbf{1}}_{\{X_{k}<N^{1/s}(\operatorname{ln}N)^{-D}\}}-\bar{b}_{N}\,\,\text{and}\,\,\bar{\bar{X}}_{k}=X_{k}{\mathbf{1}}_{\{N^{1/s}(\operatorname{ln}N)^{-D}\leq X_{k}<N^{1/s}(\operatorname{ln}N)^{\delta}\}}-\bar{\bar{b}}_{N},$
where, as before, $\bar{b}_{N}=\bar{\bar{b}}_{N}=0$ for $s\leq 1$ and they are chosen to make the expectations of LHS’s zero for $s>1.$
We will show that
[…]
and
[…]
In § 3.2 we consider the case $s<1$ where the argument is much simpler since the sums ${\bar{S}}_{n}$ and ${\bar{{\bar{S}}}}_{n}$ are almost monotone. The case $21\leq s<$ 2 is more involved and requires the use of the reflection principle (see equations (3.14) and (3.23)). In § 3.3 and § 3.4 we prove (3.3) and (3.4) respectively for $21\leq s<$ 2 .
3.2. The case of infinite expectation.
Here we prove (3.3) and (3.4) in case $s<1.$
We begin with (3.3). By (2.2) we have
$E\left|\bar{S}_{N}\right|\leq N\int_{0}^{N^{1/s}(\operatorname{ln}N)^{-D}}\frac{a_{s}{\rm{d}}t}{t^{s}}+LN\leq CN^{1/s}(\operatorname{ln}N)^{-D(1-s)}.$
Hence
$E\left|\frac{\bar{S}_{N}}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\right|\leq\frac{C}{(\operatorname{ln}N)^{D(1-s)+\alpha}}.$
Take $N_{k}=2^{k}$, then for each $\varepsilon>0$,
(3.5) $\sum_{k\geq 1}\mu\left(\left|\frac{\bar{S}_{N_{k}}}{N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}}\right|>\varepsilon\right)\leq\sum_{k\geq 1}\frac{C}{\varepsilon k^{D(1-s)+\alpha}}<\infty$
where the convergence follows from (3.1). Hence Borel-Cantelli Lemma gives
[…]
Next, consider $2^{k}<N\leq 2^{k+1}.$ Since $|\bar{S}_{N}|\leq|\bar{S}_{N_{k+1}}|+2LN_{k+1}$ and $s>1$, (3.6) gives
[…]
proving (3.3).
Next, we prove (3.4). Choose $D$ such that $D=q\delta$ where $q$ is a positive integer. For $j=0,-1,\cdots,-q,$ let
$X_{k,j}=X_{k}{\mathbf{1}}_{\{X_{k}\in(N^{1/s}(\operatorname{ln}N)^{j\delta},N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}]\}},\quad{\bar{{\bar{S}}}}_{n,j}=\sum_{k=1}^{n}X_{k,j},$
where ${\mathbf{1}}.$ is the indicator function. It suffices to show that for each $j$,
(3.8) $\max_{1\leq n\leq N}\frac{|{\bar{{\bar{S}}}}_{n,j}|}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\to 0.$
Since $|{\bar{{\bar{S}}}}_{n,j}|\leq|{\bar{{\bar{S}}}}_{N,j}|+2LN$ it suffices to prove that
(3.9) $\frac{|{\bar{{\bar{S}}}}_{N,j}|}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\to 0.$
Next, we claim that for each $m$ there is a constant $C_{m}$ such that
(3.10) $E\left({\bar{{\bar{S}}}}_{N,j}^{m}\right)\leq C_{m}N^{m/s}(\operatorname{ln}N)^{\delta m}.$
Let ${\hat{B}}_{N,j}=\left\{y:\phi(y)\in\left(N^{1/s}(\operatorname{ln}N)^{j\delta},\;N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}\right]\right\},$ $j=0,-1,\cdots,-q.$ Note that
$\frac{{\bar{{\bar{S}}}}_{N,j}}{N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}}\leq\sum_{n=1}^{N}{\mathbf{1}}_{{\hat{B}}_{N,j}}(f^{n}y).$
Thus
(3.11) $E\left(\left(\frac{\bar{\bar{S}}_{N,j}}{N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}}\right)^{m}\right)\leq\sum_{1\leq k_{1}\leq\cdots\leq k_{m}\leq N}\mu\left(\bigcap_{j=1}^{m}\left\{f^{k_{j}}y\in{\hat{B}}_{N,j}\right\}\right).$
The next result is proven in Appendix B.
Lemma 3.1.
There is a constant $C$ such that
$\sum_{1\leq k_{1}\leq\cdots\leq k_{m}\leq N}\mu\left(\bigcap_{j=1}^{m}\left\{f^{k_{j}}y\in{\hat{B}}_{N,j}\right\}\right)\leq C\left(N\mu\left({\hat{B}}_{N,j}\right)\right)^{m}.$
Combining the lemma with (2.2) we see that (3.11) is less than $C(\operatorname{ln}N)^{-j\delta sm}.$ Since $j\leq 0$ and $s<1,$ (3.10) follows.
By Markov inequality
$\mu\left(|{\bar{{\bar{S}}}}_{N,j}|\geq{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\frac{C}{{\varepsilon}^{m}(\operatorname{ln}N)^{(\alpha-\delta)m}}$
for even $m.$ Taking $m$ such that $m(\alpha-\delta)>1$ and arguing as in the proof of (3.7) we obtain
[…]
Since $j$ is arbitrary, we obtain (3.4) for $s<1.$
3.3. Controlling small values of $X$ for $21\leq s<$ 2
Here we prove (3.3) in the case $21\leq s<$ 2 .
Given $y\in\mathbb{S}$, let $C_{n}(y)$ be the continuity component of $\phi 1_{\phi\leq N^{1/s}(\operatorname{ln}N)^{-D}}\circ f^{k},$ $k\leq n$ containing $y$ and $f^{n}|_{C_{n}(y)}$ be a diffeomorphism. Let $I_{n}(y)=f^{n}C_{n}(y).$ We say that $C_{n}(y)$ is short if the length of $I_{n}(y)$ is less than $N^{-1}(\operatorname{ln}N)^{-100}$ and we say that it is long otherwise. Fix $\varepsilon>0,$ we say that $C_{n}$ is extremal if there exists $z\in C_{n}$ such that $|{\bar{S}}_{n}(z)|\geq{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}$ and for each $k<n,$ $|{\bar{S}}_{k}(z)|<{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}$ on $C_{n}.$ Note that different extremal components are disjoint. In addition,
(3.13) $|{\bar{S}}_{n}(y)|\geq\frac{3}{4}{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}\quad\text{for all }y\text{ in }C_{n}.$
Indeed, fix a large $K$ and write
$\left|{\bar{S}}_{n}(y)-{\bar{S}}_{n}(z)\right|\leq\sum_{k=1}^{n-K\operatorname{ln}N}|{\bar{X}}_{k}(y)-{\bar{X}}_{k}(z)|+|{\bar{S}}_{K\operatorname{ln}N}(f^{n-K\operatorname{ln}N}y)|+|{\bar{S}}_{K\operatorname{ln}N}(f^{n-K\operatorname{ln}N}z)|,$
then the last two terms are smaller than $2KN^{1/s}(\operatorname{ln}N)^{1-D}$ due to the cutoff in the definition of ${\bar{S}}_{n}.$ To estimate the first term, we note that since $d(f^{n}y,f^{n}z)\leq 1$ and $f$ is expanding then for $k\leq n-K\operatorname{ln}N,$ $d(f^{k}y,f^{k}z)\leq\theta^{K\operatorname{ln}N}.$ Since $\phi{\mathbf{1}}_{\{\phi<N^{1/s}(\operatorname{ln}N)^{-D}\}}$ is Lipschitz on its continuity components with Lipschitz constant less than $CN^{1+1/s}(\operatorname{ln}N)^{-(1+s)D}$, it follows that if $K$ is sufficiently large then $|{\bar{X}}_{k}(y)-{\bar{X}}_{k}(z)|\leq N^{-100}$, proving (3.13).
Let ${\mathcal{E}}_{N}$ be the set of points which belong to some extremal component for $n\leq N.$
Lemma 3.2.
If $D$ is large enough then $\mu({\mathcal{E}}_{N})\leq{C}{(\operatorname{ln}N)^{-2}}.$
Lemma 3.2 implies (3.3). In fact, combining this lemma with Borel-Cantelli Lemma we conclude that with probability $1$, for large $k$, $|{\bar{S}}_{n}|<{\varepsilon}N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}$ for $n\leq N_{k},$ $N_{k}=2^{k}.$ Since ${\varepsilon}$ is arbitrary, (3.3) follows.
Proof of Lemma 3.2.
Note that $\phi{\mathbf{1}}_{\{\phi<N^{1/s}(\operatorname{ln}N)^{-D}\}}$ is only discontinuous at two points $z_{-,N}$ and $z_{+,N}$ where $\phi(z_{\pm,N})=N^{1/s}(\operatorname{ln}N)^{-D}.$ Therefore, if $y$ belongs to a short component, there exists $k\leq n$ such that
$d\left(f^{k}y,\{z_{-,N},z_{+,N}\}\right)\leq N^{-1}(\operatorname{ln}N)^{-100}.$
Since for a fixed $k,$ $\mu\left(\left\{y:\,d\left(f^{k}y,\{z_{-,N},z_{+,N}\}\right)\leq N^{-1}(\operatorname{ln}N)^{-100}\right\}\right)=O\left(N^{-1}(\operatorname{ln}N)^{-100}\right),$ we obtain that the measure of points which belong to short components is $O\left((\operatorname{ln}N)^{-100}\right).$ It remains to estimate the measure of points which belong to long extreme components. Let $C_{n}$ be one such component and $I_{n}=f^{n}C_{n}.$ We will prove two estimates:
(3.14) $\mu\left(y\in C_{n}:|{\bar{S}}_{N}(y)|\geq\frac{{\varepsilon}}{2}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\geq C\mu(C_{n})$
for some $C>0$ independent of $N,$ $C_{n}$ and
(3.15) $\mu\left(y\in\mathbb{S}:|{\bar{S}}_{N}(y)|\geq\frac{{\varepsilon}}{2}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\frac{C}{(\operatorname{ln}N)^{2}}$
for some $C>0$ independent of $N.$
To prove (3.15) we use Chebyshev inequality. By (2.2) we have
(3.16) ${\rm Var}({\bar{X}}_{n})\leq CN^{2/s-1}(\operatorname{ln}N)^{{-D(2-s)}}.$
We now estimate
(3.17) $\left|\text{Cov}({\bar{X}}_{n},{\bar{X}}_{m})\right|\leq\begin{cases}O\left(N^{2/s-1}(\operatorname{ln}N)^{{-D(2-s)}}\right)&\text{if }|m-n|<K\operatorname{ln}N,\\ N^{-2}&\text{if }|m-n|\geq K\operatorname{ln}N,\end{cases}$
where the first bound follows from (3.16) and Cauchy-Schwartz inequality and the second bound follows from (2.4).
Summing over all $m,n$ we get
(3.18) ${\rm Var}({\bar{S}}_{N})\leq C\left(N^{2/s}(\operatorname{ln}N)^{{1-D(2-s)}}+1\right).$
Since $D(2-s)+2\alpha>3$ due to (3.2), (3.15) follows by Chebyshev inequality.
The proof of (3.14) follows similar ideas. Namely, in view of (3.13), it suffices to show that there exists $C<1$ such that
(3.19) $\mu\left(y\in C_{n}:\left|\sum_{k=n+1}^{N}{\bar{X}}_{k}(y)\right|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq C\mu(C_{n}).$
Since $f^{n}$ has bounded distortion property, (3.19) follows from
(3.20) $\mu\left(z\in I_{n}:|{\bar{S}}_{N-n}(z)|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq C\mu(I_{n}).$
Split ${\bar{S}}_{N-n}={\bar{S}}_{K\operatorname{ln}N}+({\bar{S}}_{N-n}-{\bar{S}}_{K\operatorname{ln}N}).$ The first term is bounded by
(3.21) $O\left(N^{1/s}(\operatorname{ln}N)^{1-D}\right).$
Note that
${\mathbb{E}}\left(({\bar{S}}_{N-n}-{\bar{S}}_{K\operatorname{ln}N}){\mathbf{1}}_{I_{n}}\right)^{2}=\sum_{K\operatorname{ln}N<k_{1}\leq k_{2}\leq N-n}{\mathbb{E}}\left(\bar{X}_{k_{1}}\bar{X}_{k_{2}}{\mathbf{1}}_{I_{n}}\right).$
Recall that ${\bar{X}}_{k_{1}}=\phi_{N}\circ f^{k_{1}}$ for $\phi_{N}=\phi{\mathbf{1}}_{\{\phi<N^{1/s}(\operatorname{ln}N)^{-D}\}}.$ Thus denoting $p=K\operatorname{ln}N/2,$ we obtain
${\mathbb{E}}\left(\bar{X}_{k_{1}}\bar{X}_{k_{2}}{\mathbf{1}}_{I_{n}}\right)={\mathbb{E}}\left((\phi_{N}\circ f^{k_{1}-p})(\phi_{N}\circ f^{k_{2}-p})\right)\mu(I_{n})+O\left(\left\|(\phi_{N}\circ f^{k_{1}-p})(\phi_{N}\circ f^{k_{2}-p})\right\|_{L_{1}}\eta_{n,p}\right)$
by (2.4), where
$|\eta_{n,p}|\leq C\|I_{n}\|_{BV}\theta^{p}\leq C\theta^{p}.$
If $K$ and hence $p$ is large enough then $|\eta_{n,p}|\leq N^{-(2/s+100)}$ and so
$\sum_{K\operatorname{ln}N\leq k_{1}\leq k_{2}\leq N-n}\left\|(\phi_{N}\circ f^{k_{1}-p})(\phi_{N}\circ f^{k_{2}-p})\right\|_{L_{1}}\eta_{n,p}=O\left(N^{-98}\right).$
On the other hand, similarly to (3.18) we get
$\sum_{K\operatorname{ln}N\leq k_{1}\leq k_{2}\leq N-n}{\mathbb{E}}\left((\phi_{N}\circ f^{k_{1}-p})(\phi_{N}\circ f^{k_{2}-p})\right)\mu(I_{n})\leq C\left(N^{2/s}(\operatorname{ln}N)^{{1-D(2-s)}}+1\right)\mu(I_{n}).$
It follows that
$\int_{I_{n}}\left(\sum_{k=K\operatorname{ln}N}^{N-n}{\bar{X}}_{k}(z)\right)^{2}{\rm{d}}\mu(z)\leq C\mu(I_{n})\left[N^{2/s}(\operatorname{ln}N)^{1-D(2-s)}+1\right].$
Hence
(3.22) $\mu\left(z\in I_{n}:\left|\sum_{k=K\operatorname{ln}N}^{N-n}{\bar{X}}_{k}(z)\right|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\frac{16C\mu(I_{n})}{{\varepsilon}^{2}}(\operatorname{ln}N)^{1-D(2-s)-2\alpha}.$
Combining (3.21) and (3.22) we obtain (3.14). $\square$
3.4. Controlling intermediate values of $X$ for $21\leq s<$ 2
Here we prove (3.4) in case $21\leq s<$ 2 .
The proof of (3.4) for $s\geq 1$ will proceed similarly to §3.3. Namely we define extremal components similarly to §3.3 by replacing $\phi{\mathbf{1}}_{\{\phi<N^{1/s}(\operatorname{ln}N)^{-D}\}}$ with $\phi{\mathbf{1}}_{\{N^{1/s}(\operatorname{ln}N)^{-D}\leq\phi<N^{1/s}(\operatorname{ln}N)^{\delta}\}}$ and ${\bar{S}}_{n}$ with ${\bar{{\bar{S}}}}_{n}.$ We need to obtain the following analogues of (3.14) and (3.15) for long extremal components $C_{n}$:
(3.23) $\mu\left(y\in C_{n}:|{\bar{{\bar{S}}}}_{N}(y)|>\frac{{\varepsilon}}{2}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\geq C\mu(C_{n})$
for some $C$ independent of $N,$ $C_{n}$ and
(3.24) $\mu\left(y\in\mathbb{S}:|{\bar{{\bar{S}}}}_{N}(y)|\geq\frac{{\varepsilon}}{2}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\frac{C}{(\operatorname{ln}N)^{2}}$
for some $C$ independent of $N.$
In addition, we claim that (3.13) still holds for almost every $y$ in $C_{n},$ even though the proof needs to be modified slightly. Namely it is not true that each summand in ${\bar{{\bar{S}}}}_{K\operatorname{ln}N}(y)$ is smaller than $N^{1/s}(\operatorname{ln}N)^{-D}.$ However, we claim that for almost every $y,$
[…]
for $N$ large enough. If (3.25) is not satisfied, then
(3.26) $\quad\quad\exists\,{\bar{n}},\,{\hat{n}}\leq K\operatorname{ln}N,\,\,|{\bar{n}}-{\hat{n}}|\geq(\operatorname{ln}N)^{(\alpha-\delta)/2},\,\min\left\{{\bar{{\bar{X}}}}_{\bar{n}},\,{\bar{{\bar{X}}}}_{\hat{n}}\right\}>{\bar{\varepsilon}}N^{1/s}(\operatorname{ln}N)^{-D}.$
The same argument of Lemma 2.6 shows that for almost every $y,$ (3.26) occurs for finitely many $N.$ Since the value of each term which is larger than $N^{1/s}(\operatorname{ln}N)^{-D}$ is smaller than $N^{1/s}(\operatorname{ln}N)^{\delta},$ we conclude that for almost every $y,$
(3.27) $\left|{\bar{{\bar{S}}}}_{K\operatorname{ln}N}(y)\right|\leq CN^{1/s}(\operatorname{ln}N)^{(\delta+\alpha)/2},$
which is sufficient to obtain (3.13).
Equation (3.24) follows from the Markov inequality and the following estimate.
Lemma 3.3.
Suppose $x$ is Diophantine. For each $m,$ we have
(3.28) $E\left({\bar{{\bar{S}}}}_{N}^{m}\right)\leq C_{m}N^{m/s}(\operatorname{ln}N)^{\delta m}.$
Proof.
Recall that we take
$\bar{\bar{X}}_{n}=X_{n}{\mathbf{1}}_{\{N^{1/s}(\operatorname{ln}N)^{-D}\leq X_{n}<N^{1/s}(\operatorname{ln}N)^{\delta}\}}-\bar{\bar{b}}_{N},\quad\bar{\bar{b}}_{N}=E\left(\phi{\mathbf{1}}_{\{N^{1/s}(\operatorname{ln}N)^{-D}\leq\phi<N^{1/s}(\operatorname{ln}N)^{\delta}\}}\right)$
for $21<s<$ 2 and
$\bar{\bar{X}}_{n}=X_{n}{\mathbf{1}}_{\{N^{1/s}(\operatorname{ln}N)^{-D}\leq X_{n}<N^{1/s}(\operatorname{ln}N)^{\delta}\}}$
for $s=1.$
Thus
(3.29) $E\left({\bar{{\bar{S}}}}_{N}^{m}\right)=\sum_{k=1}^{m}\sum_{1\leq n_{1}<\dots<n_{k}\leq N}\sum_{m_{1}+\dots+m_{k}=m}c_{m_{1},\dots,m_{k}}E(\bar{\bar{X}}_{n_{1}}^{m_{1}}\dots\bar{\bar{X}}_{n_{k}}^{m_{k}}),$
where $c_{m_{1},\dots,m_{r}}$ are combinatorial factors.
Note that by (2.2)
(3.30) ${\mathbb{E}}\left|\bar{\bar{X}}_{0}\right|^{l}\leq\left\{\begin{array}[]{ll}CN^{l/s-1}(\operatorname{ln}N)^{(s-l)D},&l<s,\\ CN^{l/s-1}(\operatorname{ln}N)^{(l-s)\delta},&l>s,\\ C\operatorname{ln}\operatorname{ln}N,&l=s,\\ \end{array}\right.$
and
$\left\|\bar{\bar{X}}^{l}_{0}\right\|_{BV}\leq CN^{l/s}(\operatorname{ln}N)^{l\delta}.$
(i) Suppose that $\operatorname{Sep}(n_{1},\dots,n_{k})=k.$ Then we use the multiple mixing bound of Lemma 2.4(ii). We take $K$ in the definition of separation condition so large that the error term in Lemma 2.4(ii) is $O(N^{-k}).$ There are two further subcases.
(a) $\min_{j}m_{j}>1.$ Using the second estimate in (3.30) we see that
$\prod_{j}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)\leq CN^{m/s-k}(\operatorname{ln}N)^{m\delta},$
while the error term is much smaller. Summing over all $n_{1},\dots,n_{k}$ we obtain this case gives the leading contribution of $CN^{m/s}(\operatorname{ln}N)^{m\delta}.$
(b) $\min_{j}m_{j}=1.$ If $1<s<2,$ then
$\prod_{j}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)=0,$
so each term is $O(N^{-k})$ and the total contribution of all terms is $O(1).$
If $s=1,$
$\prod_{j}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)$ $=$ $\left(\prod_{m_{j}\geq 2}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)\right)\left(\prod_{m_{j}=1}{\mathbb{E}}\bar{\bar{X}}_{n_{j}}\right)$
$\leq$ $C\left(\prod_{m_{j}\geq 2}N^{m_{j}-1}\left(\operatorname{ln}N\right)^{(m_{j}-1)\delta}\right)\left(\operatorname{ln}\operatorname{ln}N\right)^{m-k}\leq CN^{m-k}\left(\operatorname{ln}N\right)^{(m-k)\delta}\left(\operatorname{ln}\operatorname{ln}N\right)^{m}$
for $k<m$ and
$\prod_{j}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)=\prod_{j}{\mathbb{E}}\bar{\bar{X}}_{n_{j}}\leq C\left(\operatorname{ln}\operatorname{ln}N\right)^{m}$
for $k=m.$ It follows that the total contribution of all terms is bounded by $CN^{m}(\operatorname{ln}N)^{m\delta}$ when $\min_{j}m_{j}=1$ and $s=1.$
(ii) Suppose that $\operatorname{Sep}(n_{1},\dots,n_{k})=\bar{k}<k.$ Given that $x$ is Diophantine, we consider the case where
$\min_{j}\{n_{j+1}-n_{j}\}\geq\epsilon\operatorname{ln}N$
for some $\epsilon>0.$ Otherwise, the contribution to the sum is $0$ By (3.30),
${\mathbb{E}}\left|\bar{\bar{X}}_{n_{j}}\right|^{m_{j}}\leq CN^{m_{j}/s-1}(\operatorname{ln}N)^{m_{j}D}.$
In addition using Lemma 2.4(i), we see that $E(\bar{\bar{X}}_{n_{1}}^{m_{1}}\dots\bar{\bar{X}}_{n_{k}}^{m_{k}})$ is $O\left(N^{(m/s)-\bar{k}-\eta}(\operatorname{ln}N)^{Dm}\right),$ where $\eta=(k-\bar{k})\epsilon|\operatorname{ln}\theta|.$ Since the number of such terms in (3.29) is $O\left(N^{\bar{k}}(\operatorname{ln}N)^{k-\bar{k}}\right),$ we conclude that the contribution of these non-separated terms is $O\left(N^{m/s-\eta}(\operatorname{ln}N)^{(D+1)m}\right)$, which is much smaller than the RHS of (3.28). $\square$
The proof of (3.23) is similar to the proof of (3.14). Namely, as in Section 3.3, it suffices to have the following analogue of (3.20):
(3.31) $\mu\left(z\in I_{n}:|{\bar{{\bar{S}}}}_{N-n}(z)|>\frac{{\varepsilon}}{2}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq C\mu(I_{n}),$
for some $C>0.$ Split ${\bar{{\bar{S}}}}_{N-n}={\bar{{\bar{S}}}}_{K\operatorname{ln}N}+\left({\bar{{\bar{S}}}}_{N-n}-{\bar{{\bar{S}}}}_{K\operatorname{ln}N}\right).$ The first term is estimated by (3.27). For the second term arguing in the same way as in the proof of (3.20) but replacing (3.18) by (3.28), we obtain
${\mathbb{E}}\left(\left({\bar{{\bar{S}}}}_{N-n}-{\bar{{\bar{S}}}}_{K\operatorname{ln}N}\right)^{m}{\mathbf{1}}_{I_{n}}\right)\leq CN^{m/s}(\operatorname{ln}N)^{\delta m}\mu(I_{n}).$
Therefore,
(3.32) $\mu\left(z\in I_{n}:\left|{\bar{{\bar{S}}}}_{N-n}(z)-{\bar{{\bar{S}}}}_{K\operatorname{ln}N}(z)\right|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)$
$\leq\frac{16C\mu(I_{n})}{{\varepsilon}^{2}}(\operatorname{ln}N)^{-m(\alpha-\delta)}.$
Then we get (3.31) by combining (3.27) with (3.32).
Appendix A Exponential mixing of all orders for expanding maps
Proof of Lemma 2.4.
We prove part(i) by induction. When $q=2$ the statement reduces to (2.4). For $q>2$ we apply (2.4) with $\psi_{1}=\varphi_{1},$ $\psi_{2}=\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}-k_{2}})$ and obtain
$\left\|\prod_{j=1}^{q}\left(\varphi_{j}\circ f^{k_{j}}\right)\right\|_{L^{1}}=\|\varphi_{1}\|_{L^{1}}\|\psi_{2}\|_{L^{1}}+O\left(\theta^{k_{2}-k_{1}}\|\varphi_{1}\|_{BV}\|\psi_{2}\|_{L^{1}}\right),$
By induction
$\|\psi_{2}\|_{L^{1}}\leq\left(\prod_{j=2}^{q-1}\left(||\varphi_{j}||_{L^{1}}+\theta^{k_{j}-k_{j-1}}\|\varphi_{j}\|_{BV}\right)\right)\|\varphi_{q}\|_{L^{1}}.$
It follows that
$\left\|\prod_{j=1}^{q}\left(\varphi_{j}\circ f^{k_{j}}\right)\right\|_{L^{1}}\leq\left(\prod_{j=1}^{q-1}\left(\|\varphi_{j}\|_{L^{1}}+C\|\varphi_{j}\|_{BV}\theta^{k_{j+1}-k_{j}}\right)\right)\|\varphi_{q}\|_{L^{1}}.$
Proceeding in a similar way, we can prove that
$\left\|\prod_{j=1}^{q}\left(\varphi_{j}\circ f^{k_{j}}\right)\right\|_{L^{1}}\geq\left(\prod_{j=1}^{q-1}\left(\|\varphi_{j}\|_{L^{1}}-C\|\varphi_{j}\|_{BV}\theta^{k_{j+1}-k_{j}}\right)\right)\|\varphi_{q}\|_{L^{1}}.$
Now we turn to the proof of part (ii). When $q=2$ the statement also reduces to (2.4). Applying (2.4) with $\psi_{1}=\varphi_{1},$ $\psi_{2}=\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}-k_{2}}),$ we get
$\left|{\mathbb{E}}\left(\prod_{j=1}^{q}(\varphi_{j}\circ f^{k_{j}})\right)-\left({\mathbb{E}}\varphi_{1}\right){\mathbb{E}}\left(\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}})\right)\right|$
$\leq$ $C\theta^{k}\|\varphi_{1}\|_{BV}\left\|\left(\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}})\right)\right\|_{L_{1}}\leq C\theta^{k}\|\varphi_{1}\|_{BV}\left\|\left(\prod_{j=2}^{q}(|\varphi_{j}|\circ f^{k_{j}})\right)\right\|_{L_{1}}\leq C\theta^{k}\prod_{j=1}^{q}\|\varphi_{j}\|_{BV},$
where the last inequality is due to part (i) of the lemma, which we have already proved. Applying inductive estimate to ${\mathbb{E}}\left(\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}})\right)$ we obtain (2.6). $\square$
Appendix B Good targets for expanding maps
Proof of Proposition 2.3.
We prove for $\Omega_{\rho_{N},j},$ and the proof for $\Omega_{\rho_{N}}$ follows in the same approach. Denote $B(x,r_{1},r_{2})=\{y\in\mathbb{S}:r_{1}\leq|y-x|\leq r_{2}\},$ and $\tilde{A}^{+}_{\rho_{N},j},$ $\tilde{A}^{-}_{\rho_{N},j}$ to be the characteristic functions of
$B\left(x,\left(\frac{a}{\rho_{N}(c_{j}+\varepsilon)+L}\right)^{s},\left(\frac{a}{\rho_{N}(c_{j}-\varepsilon)-L}\right)^{s}\right)$
and
$B\left(x,\left(\frac{a}{\rho_{N}(c_{j}+\varepsilon)-L}\right)^{s},\left(\frac{a}{\rho_{N}(c_{j}-\varepsilon)+L}\right)^{s}\right),$
respectively. Then
$\tilde{A}^{-}_{\rho_{N},j}\leq{\mathbf{1}}_{\Omega_{\rho_{N},j}}\leq\tilde{A}^{+}_{\rho_{N},j},\quad E\tilde{A}^{+}_{\rho_{N},j}-E\tilde{A}^{-}_{\rho_{N},j}\leq C\rho_{N}^{-(s+1)}\text{ and }\|\tilde{A}^{*}_{\rho_{N},j}\|_{BV}\leq 2\mu(\Omega_{\rho_{N},j})+4,$
[…]
First we prove $(M1)_{r}.$ If $0<k_{1}<k_{2}<\dots<k_{r}\leq N$ are such that $k_{i+1}-k_{i}\geq K\operatorname{ln}N,$ where $K$ is a sufficiently large constant, and each $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r},$
$\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right)\leq{\mathbb{E}}\left(\prod_{j=1}^{r}\tilde{A}^{+}_{\rho_{N},i_{j}}\circ f^{k_{j}}\right)\leq$
$\left(\prod_{j=1}^{r-1}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{j}}+C\theta^{K\operatorname{ln}N}(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{j}}+4)\right)\right)({\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{r}})\leq\left(\prod_{j=1}^{r}\sigma_{\rho_{N},i_{j}}\right)(1+{\varepsilon}_{N})$
for ${\varepsilon}_{N}\rightarrow 0,$ which yields the RHS of $(M1)_{r}.$ The LHS is proved similarly replacing $\tilde{A}^{+}_{\rho_{N},i_{j}}$ by $\tilde{A}^{-}_{\rho_{N},i_{j}}$.
Second we show $(M2)_{r}$ assuming that $x$ is Diophanitine. Let $\operatorname{Sep}(k_{1},\dots k_{r})=m<r.$ Since $x$ is Diophantine, there exists $\epsilon>0$ such that $B(x,c\rho_{N}^{-s})\cap f^{-k}B(x,c\rho_{N}^{-s})=\emptyset$ for $k\leq\epsilon\operatorname{ln}N$ and $c=2a^{s}/c_{j}^{s}.$ Thus we assume that $k_{j+1}-k_{j}\geq\epsilon\operatorname{ln}N$ for all $j.$ Then
$\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right)$$\leq$ $C\left(\prod_{k_{j+1}-k_{j}>K\operatorname{ln}N}\sigma_{\rho_{N},i_{j}}\right)(\theta^{\epsilon\operatorname{ln}N})^{r-m}(1+{\varepsilon}_{N})\leq{\frac{C\sigma_{{\rho}_{N}}^{m}}{(\operatorname{ln}N)^{100r}}}.$
Third we prove $(M3)_{r}.$ If $k_{1}<\dots<k_{r}<l_{1}<\dots<l_{r}$ are such that $2^{i}<k_{\alpha}\leq 2^{i+1},2^{j}<l_{\beta}\leq 2^{j+1}$ for $1\leq\alpha,\beta\leq r$, then for each $(i_{1},i_{2},\dots,i_{r}),\,(j_{1},j_{2}\dots j_{r})\in\{1,\dots,p\}^{r},$
$\mu\left(\left[\bigcap_{\alpha=1}^{r}\Omega^{k_{\alpha}}_{{\rho}_{2^{i}},i_{\alpha}}\right]\bigcap\left[\bigcap_{\beta=1}^{r}\Omega^{l_{\beta}}_{{\rho}_{2^{j}},j_{\beta}}\right]\right)\leq{\mathbb{E}}\left(\left(\prod_{\alpha=1}^{r}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}\circ f^{k_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}\circ f^{k_{\beta}}\right)\right)$
$\leq\left(\prod_{\alpha=1}^{r}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}+C\theta^{Ki}(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}+4)\right)\right)\left(\prod_{\beta=1}^{r-1}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}+C\theta^{Kj}(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}+4)\right)\right)({\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{r}})$
$\leq\left(\prod_{\alpha=1}^{r}\sigma_{\rho_{2^{i}},i_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\sigma_{\rho_{2^{j}},j_{\beta}}\right)(1+{\varepsilon}_{i}).$
Finally we prove $(M4)_{*}$ assuming that $x$ is non-periodic. The proof of $(M4)_{I_{1},\dots I_{r},i_{1},\dots,i_{r}}$ is similar and will be left to the reader. We denote by
$\tau_{*}^{N}=\sum_{\overset{0<k_{1}<k_{2}<\dots<k_{r}\leq N}{{\operatorname{Sep}_{n}}(k_{1},\dots k_{r})=r}}\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N}}\right).$
Using $(M1)_{r},$ which has been proved, we have $\lim_{N\rightarrow\infty}\tau_{*}^{N}/\left(N\sigma_{\rho_{N}}\right)^{r}=1.$ Therefore, it suffices to show that
$\lim_{N\rightarrow\infty}\frac{\mu(A_{*}^{N})}{\tau_{*}^{N}}=1.$
By Boole’s inequality $\mu(A_{*}^{N})\leq\tau_{*}^{N}.$ For the lower bound we note that
(B.1) $\mu\left(A_{*}^{N}\right)\geq\tau_{*,b}^{N}-\sum^{*}\mu(\Omega^{k_{1}}_{{\rho}_{N}}\cap\dots\cap\Omega^{k_{r}}_{{\rho}_{N}}\cap\Omega^{k_{r+1}}_{{\rho}_{N}}),$
where the sum in $\sum^{*}$ is taken over all tuples satisfying $\operatorname{Sep}(k_{1},\dots k_{r})=r.$ We now consider several cases depending on the size of $p=\min_{j=1,\dots,r}|k_{r+1}-k_{j}|$.
(i) Since $x$ is non-periodic for each fixed $p_{0},$ we have that $\Omega_{{\rho}_{N}}\cap f^{-p}\Omega_{{\rho}_{N}}=\emptyset$ for all $p\leq p_{0}$ provided that $n$ is large enough. Hence the contribution of terms $p\leq p_{0}$ is zero.
(ii) Next we suppose that $p_{0}<p\leq\varepsilon\operatorname{ln}N.$ Precisely we may assume that $k_{r+1}=k_{\bar{j}}+p$ for some ${\bar{j}}\leq r$ (the case where $k_{r+1}=k_{\bar{j}}-p$ is similar). Define $\bar{A}_{{\rho}_{N}}^{+}$ to be the characteristic function of $B\left(x,\left(\frac{a}{\rho_{N}-L}\right)^{s}\right).$ Then
$\bar{A}^{+}_{\rho_{N}}\geq{\mathbf{1}}_{\Omega_{\rho_{N}}},\,{\mathbb{E}}\bar{A}^{+}_{\rho_{N}}-\mu(\Omega_{\rho_{N}})\leq C\rho_{N}^{-(s+1)}\text{ and }\|\bar{A}^{+}_{\rho_{N}}\|_{BV}\leq 2\mu(\Omega_{\rho_{N}})+4.$
For $j=1,\dots,r,$ define
${\hat{A}}_{j}=\begin{cases}\bar{A}_{{\rho}_{N}}^{+}&\text{if }j\neq{\bar{j}},\\ \bar{A}_{{\rho}_{N}}^{+}(\bar{A}_{{\rho}_{N}}^{+}\circ f^{p})&\text{if }j={\bar{j}}.\end{cases}$
If we denote by $M=\max_{x\in\mathbb{S}}|f^{\prime}(x)|,$ then
$\mu\left(\prod_{j=1}^{r+1}\Omega_{{\rho}_{N}}^{k_{j}}\right)\leq{\mathbb{E}}\left(\prod_{j=1}^{r}{\hat{A}}_{j}\circ f^{k_{j}}\right)$
$\leq$ $\left(\prod_{j\neq{\bar{j}}}\left({\mathbb{E}}{\hat{A}}_{j}+C\|\bar{A}_{{\rho}_{N}}^{+}\|_{BV}\theta^{K\operatorname{ln}N}\right)\right)\left({\mathbb{E}}{\hat{A}}_{\bar{j}}+C\|\bar{A}_{{\rho}_{N}}^{+}\|_{BV}^{2}M^{p}\theta^{K\operatorname{ln}N}\right)\leq\prod_{j=1}^{r}\left({\mathbb{E}}{\hat{A}}_{j}+n^{-10}\right)$
if $s(n)=K\operatorname{ln}N$ and $K$ is large enough. Noting that
${\mathbb{E}}({\hat{A}}_{\bar{j}})={\mathbb{E}}\left(\bar{A}_{{\rho}_{N}}^{+}(\bar{A}_{{\rho}_{N}}^{+}\circ f^{p})\right)\leq C\sigma_{{\rho}_{N}}\theta^{p}$
by (2.4), we have that each term with $p$ as in the case (ii) is at most $C\sigma_{{\rho}_{N}}^{r}\theta^{p}.$ Summing over all $r$ tuples and over all $p\in(p_{0},\varepsilon\operatorname{ln}N]$ yields a total contribution of $CN^{r}\sigma_{\rho_{N}}^{r}\theta^{p_{0}}$ for some constant $C$ independent of $p_{0}.$
(iii) The terms where ${\varepsilon}\operatorname{ln}N<p\leq K\operatorname{ln}N$ are $O(\sigma_{\rho_{N}}^{r+\eta})$ for some $\eta=\eta({\varepsilon})>0.$ Thus the total contribution of those terms are $O(N^{r}(\operatorname{ln}N)\;\sigma_{\rho_{N}}^{r+\eta}),$ which is also negligible.
(iv) The terms where $p>K\operatorname{ln}N$ for sufficiently large $K$ is $O(\sigma_{\rho_{N}}^{r+1})$ similarly to the proof of $(M1)_{r}.$ Thus the total contribution of those terms are $O((N\sigma_{\rho_{N}})^{r+1})$ which is negligible compared to $\tau_{*}^{N}.$
Summing the estimates of cases (i)–(iii) above we see that the second term in the RHS of (B.1) is negligible comparing to the first, which proves $(M4)_{*}.$square
Proof of Lemma 2.6.
Let
$H_{N,{\bar{\varepsilon}}}=\left\{y:\exists\,{\bar{n}},\,{\hat{n}}\leq N,\,\,(\operatorname{ln}N)^{(\alpha-{\bar{\alpha}})/2}\leq|{\bar{n}}-{\hat{n}}|\leq 2s(N),\,\,\min\{X_{\bar{n}},\,X_{\hat{n}}\}>{\bar{\varepsilon}}N^{1/s}(\operatorname{ln}N)^{\alpha-1}\right\},$
and
$\tilde{H}_{k,{\bar{\varepsilon}}}=\left\{y:\exists\,{\bar{n}},\,{\hat{n}}\leq 2^{k+1},\,\,k^{(\alpha-{\bar{\alpha}})/2}\leq|{\bar{n}}-{\hat{n}}|\leq 2s(2^{k+1}),\,\,\min\{X_{\bar{n}},\,X_{\hat{n}}\}>{\bar{\varepsilon}}2^{k/s}k^{\alpha-1}\right\}.$
Since $H_{N,{\bar{\varepsilon}}}\subset\tilde{H}_{k,{\bar{\varepsilon}}}$ for $2^{k}\leq N\leq 2^{k+1},$ it suffices to show that $\tilde{H}_{k,{\bar{\varepsilon}}}$ happens only for finitely many $k.$ Using (2.4) and similar method in the proof of Proposition 2.3 we conclude that
$\mu(\tilde{H}_{k,{\bar{\varepsilon}}})\leq\frac{C2^{k+1}(k+1)}{2^{k}k^{s(\alpha-1)}}\theta^{k(\alpha-{\bar{\alpha}})/2}=2Ck^{s(1-\alpha)+1}2^{k(\alpha-{\bar{\alpha}})\operatorname{ln}\theta/2}.$
Hence $\sum_{k}\mu(\tilde{H}_{k,{\bar{\varepsilon}}})<\infty$ and the result follows from the classical Borel-Cantelli Lemma. $\square$
Proof of Lemma 3.1.
Let $p=\min_{1\leq j\leq m-1}\{k_{j+1}-k_{j}\}.$ Since $x$ is non-periodic, for each fixed $p_{0},$ we have that ${\hat{B}}_{N,j}\cap f^{-p}{\hat{B}}_{N,j}=\emptyset$ for all $p\leq p_{0}$ provided that $N$ is large enough because $x$ is non-periodic. Thus the contribution of the terms with tuples satisfying $p\leq p_{0}$ is zero for large $N.$
We suppose that $\operatorname{Sep}(k_{1},\dots k_{m})=l<m.$ If $p_{0}<p\leq\varepsilon\operatorname{ln}N,$ Arguing as the case $(M4)_{*}$(ii) in the proof of Proposition 2.3, we conclude that each term is smaller than $C\mu({\hat{B}}_{N,j})^{l}\theta^{p}$ for fixed $p.$ Since the total number of terms is $O\left(N^{l}p^{m-l}\right),$ we have that total contribution of the terms with $p={\hat{p}}$ is $O\left(N^{l}{\hat{p}}^{m-l}\mu({\hat{B}}_{N,j})^{l}\theta^{\hat{p}}\right).$ The sum over tuples with ${\hat{p}}\in[p_{0},{\varepsilon}\operatorname{ln}N]$ gives a total contribution of $CN^{l}{p_{0}}^{m-l}\mu({\hat{B}}_{N,j})^{l}\theta^{p_{0}}$ for fixed $l$ and $p_{0}$ large enough (independent of $N$).
On the other hand if $\varepsilon\operatorname{ln}N<p\leq K\operatorname{ln}N,$ then similarly to the proof of the case $(M4)_{*}$(iii) we obtain that
$\mu\left(\bigcap_{j=1}^{m}\left\{f^{n_{j}}y\in{\hat{B}}_{N,j}\right\}\right)\leq C\mu({\hat{B}}_{N,j})^{l+\eta}.$
Summing over all $r$ tuples and over all $p\in(\varepsilon\operatorname{ln}N,K\operatorname{ln}N]$ yields a total contribution that is not greater than $C\left(N\mu({\hat{B}}_{N,j})\right)^{l}\mu({\hat{B}}_{N,j})^{\eta}\left(\operatorname{ln}N\right)^{m-l}.$
Finally, if $\operatorname{Sep}(k_{1},\dots k_{m})=m,$ similarly to the case $(M1)_{r}$ in the proof of Proposition 2.3, we obtain that
$\mu\left(\bigcap_{j=1}^{m}\left\{f^{k_{j}}y\in{\hat{B}}_{N,j}\right\}\right)\leq C\mu({\hat{B}}_{N,j})^{m}.$
Summing over all tuples $(k_{1},\dots k_{m})$ and using that $(N\mu({\hat{B}}_{N,j}))^{l}\leq C(N\mu({\hat{B}}_{N,j}))^{m}$ for $l<m$, we obtain the result. $\square$

### v3

@@
 Abstract.
-We study heavy tailed random variables obtained by sampling an observable with a power singularity along an orbit of an expanding map on the circle.
-In the regime where the observables have infinite variance, we show that the set of limit points of the ergodic sum process, rescaled by $N^{1/s}(\operatorname{ln}N)^{\alpha}$, where $s$ is the index of limiting stable law, equals to the set of piecewise constant increasing functions with finitely many jumps.
-The proof relies on the multiple Borel-Cantelli Lemma proven in Dolgopyat, Fayad and Liu [J.
-Mod.
-Dyn.
-18 (2022) 209–289].
+We establish functional limit theorems for ergodic sums of observables with power singularities for expanding circle maps.
+In the regime where the observables have infinite variance, we show that when rescaled by $N^{1/s}(\operatorname{ln}N)^{\alpha}$, the partial sum process has limit points consisting precisely of increasing piecewise constant functions with finitely many jumps.
+Our approach combines trimming techniques with a multiple Borel-Cantelli argument.
+It provides a functional law of the iterated logarithm for heavy-tailed processes where classical almost sure invariance principles do not apply.
 1. Introduction
-Let $X_{k}$ be a sequence of independent and identically distributed random variables with mean 0 and variance 1.
-Let $S_{n}=\sum_{k=1}^{n}X_{k}.$ Consider a function $W_{N}:[0,1]\to\mathbb{R}$ defined by $W_{N}\left(\frac{n}{N}\right)=\frac{S_{n}}{\sqrt{N}}$ at rational points $\frac{n}{N}$ and by linear interpolation in between.
-The functional central limit theorem states that the distribution of $W_{N}(\cdot)$ converges to the standard Wiener measure as $N\to\infty$.
-Since Wiener measure has full support, it follows (see e.g.
-[41]) that with probability $1$ for any continuous function $h$ on $[0,1]$, there is a sequence $N_{j}$ such that $W_{N_{j}}$ converges uniformly to $h.$
-It is natural to ask whether one can obtain a smaller set of limits by considering a larger normalization (so that the renormalized function is close to 00 most of the time).
-Strassen [42] answered this question by showing that if one takes
+Let $\{X_{k}\}_{k=1}^{\infty}$ be a sequence of independent and identically distributed random variables with mean 0 and variance 1.
+Define the partial sum $S_{n}=\sum_{k=1}^{n}X_{k}$ and consider the function $W_{N}:[0,1]\to\mathbb{R}$ given by
+$W_{N}\left(\frac{n}{N}\right)=\frac{S_{n}}{\sqrt{N}}$
+at rational points $n/N,$ extended by linear interpolation elsewhere.
+The functional central limit theorem states that $W_{N}(\cdot)$ converges in distribution to the standard Wiener measure as $N\to\infty$.
+Since Wiener measure has full support on $C([0,1])$, it follows that almost surely, for any continuous $h\in C([0,1])$, there exists a subsequence $N_{j}$ such that $W_{N_{j}}$ converges uniformly to $h$ (see e.g.
+[28]).
+It is natural to ask whether one can obtain a smaller set of limits by considering a larger normalization. Strassen [29] answered this question by defining
 ${\widetilde{W}}_{N}\left(\frac{n}{N}\right)=\frac{S_{n}}{\sqrt{2N\operatorname{ln}\operatorname{ln}N}},$
-then with probability $1$, $h$ is a limit point of $\{{\widetilde{W}}_{N}\}$ if and only if $h$ is an absolutely continuous function such that $h(0)=0$ and $\int_{0}^{1}(h^{\prime}(x))^{2}{\rm{d}}x\leq 1.$ This result is referred to as the functional form of law of the iterated logarithm (FLIL) in the literature [16, 39].
-FLIL is a generalization of Khintchine-Kolmogoroff’s original form [28, 29] of the law of the iterated logarithm (LIL) and Hartman–Wintner’s LIL [17] for random walks.
-The method used in [42] for proving FLIL is known as Strassen’s almost sure invariance principle (ASIP).
-ASIP was further investigated for martingales (see e.g.
-[16, 20, 43]), which states that a martingale satisfying certain moment conditions is with probability $1$ close to a Brownian motion.
-Consequently, FLIL holding for Brownian motion is still valid for martingales.
-ASIP can be extended to dynamical systems [39] by approximating weakly dependent random variables by martingales.
-This method has been applied to show ASIP for various dynamical systems (see e.g.
-[7, 8, 13, 15, 18, 19, 21, 30, 31, 33, 36, 38, 40]).
-Specifically, these systems include uniformly hyperbolic systems [10], Lorenz attractors [22], non-uniformly hyperbolic systems [32, 34, 35], partially hyperbolic systems [11], dispersing billiards [3, 5] and compact group extensions [14].
+and proving that almost surely, the limit points of $\{\widetilde{W}_{N}\}$ are precisely the absolutely continuous functions $h$ with $h(0)=0$ and $\int_{0}^{1}|h^{\prime}(x)|^{2}{\rm{d}}x\leq 1.$ This result is referred to as the functional form of law of the iterated logarithm (FLIL) in the literature [12, 27].
+FLIL generalizes the classical LIL for random variables [13, 21, 22].
+Strassen’s proof relies on the almost sure invariance principle (ASIP), which establishes that, under suitable conditions, a martingale is almost surely close to a Brownian motion.
+Consequently, FLIL extends to martingales satisfying appropriate moment conditions [12, 30].
+The ASIP framework further extends to dynamical systems via martingale approximations for weakly dependent processes [27], establishing FLIL for systems including uniformly hyperbolic systems [8], Lorenz attractors [15], non-uniformly hyperbolic systems [14, 23, 24, 25], partially hyperbolic systems [9], dispersing billiards [3, 5] and compact group extensions [11].
 However, there are still many examples of dynamical systems where the limit distributions of ergodic sums are non-Gaussian.
-For these systems, the above approach is not applicable and ASIP does not hold.
-It is therefore of interest to develop a new method for finding a suitable normalizing sequence $L_{N}$ so that the function ${\tilde{S}}_{N}$ defined by ${\tilde{S}}_{N}(\frac{n}{N})=\frac{S_{n}}{L_{N}}$ has a non-trivial set of limit points.
-In this paper, we address the above problem in a simple setting.
-Let $S_{n}(y)=\sum_{k=1}^{n}\phi(f^{k}y)$ where $f$ is an expanding map of the circle $\mathbb{S}$ and $\phi$ has a power singularity of the type $|y-x|^{-1/s}$ for fixed $x\in\mathbb{S}.$ Assuming that $0<s<2,$ $s\neq 1$ and $x$ is a Lebesgue typical point, we show that a reasonable choice of $L_{N}$ is $L_{N}=N^{1/s}(\operatorname{ln}N)^{\alpha}.$ If $\alpha$ satisfies ${[(r+1)s]}^{-1}<\alpha\leq{(rs)}^{-1},$ then the set of limit points consists of piecewise constant increasing functions starting at 00 with at most $r$ jumps.
+For these systems, the aforementioned approach is not applicable, and ASIP does not hold.
+It is therefore of interest to develop a new method for determining a suitable normalizing sequence $L_{N}$ such that the function
+${\tilde{S}}_{N}\left(\frac{n}{N}\right)=\frac{S_{n}}{L_{N}}$
+has a non-trivial set of limit points.
+In this paper, we address the above problem in a simple setting. Let
+$S_{n}(y)=\sum_{k=1}^{n}\phi(f^{k}y),$
+where $f$ is an expanding circle map $\mathbb{S},$ and $\phi$ has a power singularity of the form $|y-x|^{-1/s}$ for a fixed point $x\in\mathbb{S}.$ Assuming that $0<s<2$ and $x$ is a Lebesgue typical point, we show that a natural normalizing sequence is
+$L_{N}=N^{1/s}(\operatorname{ln}N)^{\alpha}.$
+If $\alpha$ satisfies
+${[(r+1)s]}^{-1}<\alpha\leq{(rs)}^{-1},$
+then the set of limit points consists of piecewise constant increasing functions starting at 0 with at most $r$ jumps.
 Before formulating our result precisely, let us mention two related topics.
-The first topic is Laws of Large Numbers for trimmed sums (see [1, 23, 24, 25, 27, 37]).
-Trimming is an important tool in our analysis since it allows us to discard all the summands in $S_{N}$ except for $r$ largest ones.
-However, this is only part of our analysis, since the contribution of the retained terms needs to be analyzed in more detail than in the papers mentioned above.
-We also note that unlike the above mentioned papers, our results apply not only to sums with infinite mean but also to the sums with infinite variance.
-The second topic is the application of the classical Borel-Cantelli Lemma to estimate the rate of convergence in the Law of Large Numbers (see [4, 6, 9, 26] and references wherein).
+The first topic is laws of large numbers for trimmed sums (see [1, 17, 18, 20, 26]).
+Trimming is an important tool in our analysis, as it allows us to discard all summands in $S_{N}$ except for the $r$ largest ones.
+However, this is only part of our analysis, as the contribution of the retained terms requires more detailed analysis than in the papers mentioned above.
+Additionally, unlike the studies cited above, our results apply not only to sums with infinite mean but also to sums with infinite variance.
+The second topic involves the application of the classical Borel-Cantelli lemma to estimate the rate of convergence in laws of large numbers (see [4, 6, 7, 19] and references wherein).
 This topic is closer to the subject of our paper.
-However, the papers mentioned above deal with a single $N$ at a time, while we obtain the functional version of the deviation result.
-Fundamental to our approach is the multiple Borel-Cantelli Lemma proposed recently in [12], which provides a criterion to determine when several rare events occur at the same time scale.
-In Section 2, we present the main theorem and sketch its proof using the multiple Borel-Cantelli Lemma.
-In Section 3, we focus on a key lemma which allows us to control the trimmed sums.
-The proof in Section 3 uses the reflection principle for the processes under consideration.
-For completeness, we include some standard proofs as appendices.
+However, while the papers mentioned above deal with a single $N$ at a time, we obtain the functional version of the deviation result.
+Fundamental to our approach is the multiple Borel-Cantelli lemma proposed recently in [10], which provides a criterion for determining when multiple rare events occur at the same time scale.
+The paper is structured as follows.
+Section 2 states the main theorem and outlines its proof via the multiple Borel-Cantelli lemma.
+Section 3 proves a key lemma controlling trimmed sums using reflection principles.
+Standard technical results are deferred to appendices.
 2. An analogue of LIL for Smooth expanding maps
 2.1. The main result.
-Let $f$ be a $C^{2}$ expanding map of the circle $\mathbb{S}.$ Here we say $f$ is expanding if there exists $\lambda>1$ such that $|f^{\prime}(x)|>\lambda$ for any $x\in\mathbb{S}.$ Then there is a unique $f$-invariant measure $\mu$ that is equivalent to Lebesgue measure [44].
-Fix $x,$ and let $\phi$ be an unbounded function of the form
-$\phi(y)=a|y-x|^{-1/s}+\psi(y),$
-where $\psi$ is a $C^{1}$ function and $s>0.$ Let
-$X_{n}(y)=\phi(f^{n}y).$
-Denote $\|\cdot\|_{\infty}$ the uniform norm and $L=\|\psi\|_{\infty}.$ Then $\mu(X_{n}>t)=\mu(\phi(y)>t)$ and
-(2.1) $\mu\left(|y-x|<\left(\frac{a}{t+L}\right)^{s}\right)\leq\mu(\phi(y)>t)\leq\mu\left(|y-x|\leq\left(\frac{a}{t-L}\right)^{s}\right)$
-for $t>L.$ It follows that
-(2.2) $\mu(X_{n}>t)\sim a_{s}t^{-s}$
-for sufficiently large $t$ and $a_{s}=\rho_{0}a^{s}$, where $\rho_{0}$ is the density of $\mu$ at $x.$ For a positive $\alpha$, set
-$W_{N}(t)=\frac{\left(\sum_{n\leq Nt}X_{n}\right)-a_{N}}{N^{1/s}(\operatorname{ln}N)^{\alpha}},$
-where $a_{N}=0$ if $s\leq 1$ and $a_{N}=Nt\int\phi(y){\rm{d}}\mu(y)$ if $s>1.$
-We regard $W_{N}(\cdot)$ as an element of the space ${\mathbb{D}}[0,1]$ of functions from $[0,1]$ to [,+][$\infty$,$\infty$][ - ∞ , + ∞ ] that are right continuous and have left limits, where we endow [,+][$\infty$,$\infty$][ - ∞ , + ∞ ] with the distance $d(z_{1},z_{2})=\left|\tan^{-1}(z_{1})-\tan^{-1}(z_{2})\right|.$ There are several possible topologies on ${\mathbb{D}}[0,1]$, the most relevant for us is Skorokhod’s $J_{1}$ topology.
-Let $\Lambda$ be the space of order preserving homeomorphisms on $[0,1].$ Then $J_{1}$ topology is defined using the distance
-$d_{J_{1}}(h_{1},h_{2})=\inf_{\lambda\in\Lambda}\max\left\{\|\lambda-Id\|_{\infty},\,\max_{t\in[0,1]}d(h_{1}(t),h_{2}(\lambda(t)))\right\}.$
-Two functions are $J_{1}$ close, if their graphs are close with respect to the Hausdorff distance.
-In our analysis we shall use several facts about $J_{1}$ topology.
-First, if two functions are close with respect to the supremum norm, then they are $J_{1}$ close.
-Second, if one takes ${\mathbb{H}}_{r}$ to be the space of step functions with (at most) $r$ jumps, then $W\in{\mathbb{H}}_{r}$ can be written in the form
+Let $f$ be a $C^{2}$ expanding map of the circle $\mathbb{S}.$ Here $f$ is called expanding if there exists $\gamma>1$ such that $|f^{\prime}(x)|>\gamma$ for all $x\in\mathbb{S}.$ It is known that $f$ admits a unique invariant measure $\mu$ that is absolutely continuous with respect to the Lebesgue measure [31].
+Fix $x\in\mathbb{S},$ and consider an unbounded observable $\phi$ of the form
+$\phi(y)=\frac{a}{|y-x|^{1/s}}+\psi(y),$
+where $\psi$ is a $\psi\in C^{1}(\mathbb{S}),$ $a>0,$ and $s>0.$ Define the sequence $X_{k}(y)=\phi(f^{k}y).$ Then
+(2.1) $\mu\left(\left\{y:X_{k}(y)>t\right\}\right)\sim a_{s}t^{-s}$
+for sufficiently large $t,$ where $a_{s}=\rho_{0}a^{s}$ and $\rho_{0}$ denotes the density of $\mu$ at $x.$ Now for a positive parameter $\alpha,$ define the process
+(2.2) $W_{N}(t)=\frac{\left(\sum_{k\leq Nt}X_{k}\right)-a_{N}}{N^{1/s}(\operatorname{ln}N)^{\alpha}},$
+with centering term $a_{N}=0$ if $s\leq 1,$ and $a_{N}=Nt\int\phi(y){\rm{d}}\mu(y)$ if $s>1.$
+We regard $W_{N}(\cdot)$ as a sequence of random elements in the space ${\mathbb{D}}[0,1]$ of càdlàg functions from $[0,1]$ to ¯=[$\overline{\mathbb{R}}$$\infty$$\infty$], equipped with the metric
+$d(z_{1},z_{2})=\left|\tan^{-1}(z_{1})-\tan^{-1}(z_{2})\right|.$
+While several topologies exist on ${\mathbb{D}}[0,1]$, the Skorokhod $J_{1}$ topology is most relevant for our analysis.
+Let $\Lambda$ denote the set of strictly increasing continuous functions $\lambda:[0,1]\to[0,1]$ with $\lambda(0)=0$ and $\lambda(1)=1$.
+The $J_{1}$ topology is defined [16] using the distance
+$d_{J_{1}}(h_{1},h_{2})=\inf_{\lambda\in\Lambda}\left\{\|h_{1}\circ\lambda-h_{2}\|_{\infty}+\|\lambda-id\|_{\infty}\right\},$
+where $\|\cdot\|_{\infty}$ denotes the uniform norm.
+Key properties of the $J_{1}$ topology that are employed in our analysis include:
+  1. (i)
+    Convergence in uniform norm implies convergence in $J_{1}$ topology.
+  2. (ii)
+    Let ${\mathbb{H}}_{r}$ be the space of step functions with at most $r$ positive jumps, then $W\in{\mathbb{H}}_{r}$ can be written as
     $W(t)=\sum_{m=1}^{\bar{r}}c_{m}{\mathbf{1}}_{t\leq t_{m}}(t),$
-where $\bar{r}\leq r$ and $\{c_{m}\}_{m=1}^{r}$ and $\{t_{m}\}_{m=1}^{r}$ are $\bar{r}$–tuples such that $c_{m}$’s are positive and $t_{m}$’s are increasing.
-Two elements $W_{1}$ and $W_{2}$ of ${\mathbb{H}}_{r}$ with exactly $r$ jumps are $J_{1}$ close provided that $\{c_{m}^{1}\}_{m=1}^{r}$ is close to $\{c_{m}^{2}\}_{m=1}^{r}$ and $\{t_{m}^{1}\}_{m=1}^{r}$ is close to $\{t_{m}^{2}\}_{m=1}^{r}$.
-Third, $J_{1}$ topology allows us to ignore small jumps.
-We also need the following definition.
-We call $x$ Diophantine if there exists $\epsilon=\epsilon(x)>0$ and $\rho_{0}=\rho_{0}(x)>0$ such that for all positive $\rho\leq\rho_{0}$ and all positive integer $k\leq\epsilon|\operatorname{ln}\rho|,$ $f^{-k}B(x,\rho)\cap B(x,\rho)=\emptyset.$ Due to Lemma 5 in [2], $\mu-$almost every $x$ is Diophantine.
+    where $\bar{r}\leq r$, the coefficients $\{c_{m}\}_{m=1}^{\bar{r}}$ represent positive amplitudes, and the points $\{t_{m}\}_{m=1}^{\bar{r}}$ are strictly increasing jump times.
+Two elements $W_{1},W_{2}\in{\mathbb{H}}_{r}$ with exactly $r$ jumps are $J_{1}$-close if their respective amplitudes and jump times are close.
+  3. (iii)
+    The $J_{1}$ topology allows us to ignore small jumps in the limit.
+To present our results, we introduce the notion of Diophantine points.
+A point $x$ is Diophantine if there exist constants $\epsilon=\epsilon(x)>0$ and $\rho_{0}=\rho_{0}(x)>0$ such that for all $\rho\leq\rho_{0}$ and all positive integers $k\leq\epsilon|\operatorname{ln}\rho|,$
+$f^{-k}B(x,\rho)\cap B(x,\rho)=\emptyset.$
+This notion of Diophantine points characterizes a specific type of slow recurrence behavior.
+Rooted in Diophantine approximation, which restricts how well real numbers can be approximated by rational ones, this property effectively prevents dense orbit clustering in dynamical systems.
+Unlike periodic points, Diophantine points are quantitatively non-periodic.
+Furthermore, as stated in Lemma 5 of [2], $\mu$-almost every $x$ is Diophantine.
+With this definition established, we now proceed to formally state our main theorem.
 Theorem 2.1.
-Suppose that either
-(i) $10<s<$ 1 and $x$ is non-periodic or
-(ii) $21\leq s<$ 2 and $x$ is Diophantine.
-Let $\alpha$ be such that $\frac{1}{(r+1)s}<\alpha\leq\frac{1}{rs}$ for some $r\in{\mathbb{N}}.$ Then for almost every $y$ the following holds: the set of $J_{1}$ limit points of $\{W_{N}(t)\}_{N=1}^{\infty}$ is equal to ${\mathbb{H}}_{r}.$
+Under the above hypotheses, assume that either
+(i) $0<s<1,$ or
+(ii) $1\leq s<2$ and $x$ is Diophantine.
+Let $\alpha$ satisfy
+$\frac{1}{(r+1)s}<\alpha\leq\frac{1}{rs}$
+for some $r\in{\mathbb{N}}.$ Then for almost every $y,$ the following holds: the set of $J_{1}$ limit points of $\{W_{N}(t)\}_{N=1}^{\infty}$ is equal to ${\mathbb{H}}_{r}.$
 2.2. Multiple Borel-Cantelli lemma.
-The proof of Theorem 2.1 requires control on the number of times $k\leq n$ such that the events $\left\{y:\,c_{j}-\varepsilon\leq X_{k}(y)\leq c_{j}+\varepsilon\right\}$ occur for some $c_{j}>0$ and $0<\varepsilon<c_{j}.$ If several such events occur for some $n$ we also need to control the occurrence time of these events.
-To this end, we shall use a generalization of the multiple Borel-Cantelli Lemma proven in [12], those statement is recalled in this subsection.
-Consider a sequence $\{\rho_{n}\}_{n=1}^{\infty}$ and a family of events $\{\Omega_{\rho_{n}}\}_{n=1}^{\infty}.$ Suppose that there exist ${\hat{\varepsilon}}>0$ and a decomposition of $\Omega_{\rho_{n}}$ for each $\rho_{n}$ such that
+Given sequences of positive numbers $\{\rho_{N}\}$ and $\{c_{j}\},$ and some $\varepsilon<c_{j}.$ The proof of Theorem 2.1 requires controlling the number of times $k\leq N$ such that the following events occur:
+$\left\{y:X_{k}(y)>\rho_{N}\right\}$
+and
+$\left\{y:\,c_{j}-\varepsilon\leq X_{k}(y)\leq c_{j}+\varepsilon\right\}.$
+Furthermore, if multiple such events occur for some $k,$ controlling their timing is essential.
+To address this,we employ an extension of the multiple Borel–Cantelli lemma, generalizing Theorem 2.12 in [10].
+Throughout, we denotes $C$ a constant depending on the fixed number $r$ of the occurring events, which may vary across lines but remains independent of the time scale $N,$ the events $\Omega_{\rho_{N}},$ and the iteration order of $f,$ etc.
+Consider a probability space $(\Omega,\mathcal{F},\mu)$ and a family of events $\{\Omega_{\rho_{N}}\}_{N=1}^{\infty}.$ Suppose there exists a constant $\hat{\varepsilon}>0$ and a decomposition of $\Omega_{\rho_{N}}$ such that
 $\Omega_{\rho_{N}}=\bigcup_{i=1}^{p}\Omega_{\rho_{N},i}\quad\text{and}\quad\mu(\Omega_{\rho_{N},i})\geq\hat{\varepsilon}\mu(\Omega_{\rho_{N}}).$
-We denote by
+We introduce the following notation:
 (2.3) $\Omega^{k}_{\rho_{N}}=f^{-k}(\Omega_{\rho_{N}}),\,\,\,\Omega^{k}_{\rho_{N},i}=f^{-k}(\Omega_{\rho_{N},i}),\,\,\,\sigma_{\rho_{N}}=\mu(\Omega_{\rho_{N}})\,\,\,{\text{and}}\,\,\,\sigma_{\rho_{N},i}=\mu(\Omega_{\rho_{N},i}).$
-Fix $r\in\mathbb{N^{*}},$ for each $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r},$ we will characterize whether with probability 1, for infinitely many $n,$ the events $\Omega_{\rho_{n},i_{j}}^{k_{j}}$ happen for all $1\leq j\leq r.$
-For an arbitrary $r$-tuple $0<k_{1}<\dots<k_{r}\leq n,$ we take the separation indices
-$\operatorname{Sep}_{n}(k_{1},\dots,k_{r})=\text{Card}\left\{j\in\{0,\dots r-1\}:k_{j+1}-k_{j}\geq s(n)\right\},\quad k_{0}:=0;$
-$\widehat{{\rm Sep}}_{n}(k_{1},\dots,k_{r})=\text{Card}\left\{j\in\{0,\dots r-1\}:k_{j+1}-k_{j}\geq\hat{s}(n)\right\},\quad k_{0}:=0$
-for functions $s:{\mathbb{N}}\rightarrow{\mathbb{N}},$ $s(n)\leq(\operatorname{ln}n)^{2}$ and $\hat{s}:{\mathbb{N}}\rightarrow{\mathbb{N}},$ $\varepsilon n\leq\hat{s}(n)<n(1-q)/{(2r)}$ for some $0<q<1,$ $0<\varepsilon<(1-q)/{(2r)},$ respectively.
-Suppose the function $\sigma.$ and $\sigma_{\cdot,i}$ defined in (2.3) are increasing and there is a sequence ${\varepsilon}_{n}\to 0$ satisfying some of the mixing conditions below.
+For a fixed $r\in\mathbb{N^{*}},$ and each tuple $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r},$ we characterize whether, with probability 1, the events $\left\{\Omega_{\rho_{N},i_{j}}^{k_{j}}\right\}$ occur for all $1\leq j\leq r$ for infinitely many $N.$
+For an $r$-tuple $0<k_{1}<\dots<k_{r}\leq N,$ we take the separation indices
+$\operatorname{Sep}_{N}(k_{1},\dots,k_{r})=\sharp\left\{j:0\leq j\leq r-1,\,k_{j+1}-k_{j}\geq s(N)\right\},\,\,\,k_{0}:=0,$
+$\widehat{{\rm Sep}}_{N}(k_{1},\dots,k_{r})=\sharp\left\{j:0\leq j\leq r-1,\,k_{j+1}-k_{j}\geq\hat{s}(N)\right\},\,\,\,k_{0}:=0,$
+where $s:\mathbb{N}\to\mathbb{N}$ satisfies
+$s(N)\leq C(\operatorname{ln}N)^{2},$
+and $\hat{s}$$\mathbb{N}$$\rightarrow{\mathbb{N}}$ satisfies
+$\varepsilon N\leq\hat{s}(N)<qN/{(2r)}$
+for some constants $0<q<1$ and $0<\varepsilon<q/{(2r)}.$
+Assume there exists a sequence ${\varepsilon}_{N}\to 0$ satisfying the following conditions, which quantify mixing and non-clustering for $\left\{\Omega^{k}_{{\rho}_{N}}\right\}$ and $\left\{\Omega^{k}_{{\rho}_{N},i}\right\}$ under appropriate separations.
   • $(M1)_{r}$
-    If $0<k_{1}<k_{2}<\dots<k_{r}\leq n$ are such that $\operatorname{Sep}_{n}(k_{1},\dots,k_{r})=r$ and for each $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r},$
-    $\left(\prod_{j=1}^{r}\sigma_{\rho_{n},i_{j}}\right)(1-{\varepsilon}_{n})\leq\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{n},i_{j}}\right)\leq\left(\prod_{j=1}^{r}\sigma_{\rho_{n},i_{j}}\right)(1+{\varepsilon}_{n});$
+    For $0<k_{1}<k_{2}<\dots<k_{r}\leq N$ with $\operatorname{Sep}_{N}(k_{1},\dots,k_{r})=r$ and all $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r},$
+    $\left(\prod_{j=1}^{r}\sigma_{\rho_{N},i_{j}}\right)\left(1-{\varepsilon}_{N}\right)\leq\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right)\leq\left(\prod_{j=1}^{r}\sigma_{\rho_{N},i_{j}}\right)\left(1+{\varepsilon}_{N}\right).$
   • $(M2)_{r}$
-    There exists $K>0$ such that if $0<k_{1}<k_{2}<\cdots<k_{r}\leq n$ are such that $\operatorname{Sep}_{n}(k_{1},\dots,k_{r})=m<r$, then
-    $\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{n}}\right)\leq{\frac{K\sigma_{{\rho}_{n}}^{m}}{(\operatorname{ln}n)^{100r}}};$
+    For $0<k_{1}<k_{2}<\cdots<k_{r}\leq N$ with $\operatorname{Sep}_{N}(k_{1},\dots,k_{r})=m<r$,
+    $\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N}}\right)\leq{\frac{C\sigma_{{\rho}_{N}}^{m}}{(\operatorname{ln}N)^{100r}}}.$
   • $(M3)_{r}$
-    If $0<k_{1}<\dots<k_{r}<l_{1}<\dots<l_{r}$ are such that $2^{i}<k_{\alpha}\leq 2^{i+1},2^{j}<l_{\beta}\leq 2^{j+1}$ for $1\leq\alpha,\beta\leq r$, $j-i\geq b$ for some constant $b\geq 2,$
+    For $0<k_{1}<\dots<k_{r}<l_{1}<\dots<l_{r}$ where $2^{i}<k_{\alpha}\leq 2^{i+1},$ $2^{j}<l_{\beta}\leq 2^{j+1}$ ($1\leq\alpha,\beta\leq r$), $j-i\geq b$ ($b\geq 2$),
     $\widehat{{\rm Sep}}_{2^{i+1}}(k_{1},\ldots,k_{r})=r,\quad\widehat{{\rm Sep}}_{2^{j+1}}(l_{1},\ldots,l_{r})=r,\quad l_{1}-k_{r}\geq\hat{s}(2^{j+1})$
     and for each $(i_{1},i_{2},\dots,i_{r}),\,(j_{1},j_{2},\dots,j_{r})\in\{1,\dots,p\}^{r},$
-    $\mu\left(\left[\bigcap_{\alpha=1}^{r}\Omega^{k_{\alpha}}_{{\rho}_{2^{i}},i_{\alpha}}\right]\bigcap\left[\bigcap_{\beta=1}^{r}\Omega^{l_{\beta}}_{{\rho}_{2^{j}},j_{\beta}}\right]\right)\leq\left(\prod_{\alpha=1}^{r}\sigma_{\rho_{2^{i}},i_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\sigma_{\rho_{2^{j}},j_{\beta}}\right)(1+{\varepsilon}_{i}).$
-We consider the events
-$A_{*}^{n}=\{\exists\,0<k_{1}<k_{2}<\dots<k_{r}\leq n,\,\,\operatorname{Sep}(k_{1},\dots k_{r})=r,\,\,\,\Omega^{k_{j}}_{{\rho}_{n}}\text{ happens}\}$
+    $\mu\left(\left(\bigcap_{\alpha=1}^{r}\Omega^{k_{\alpha}}_{{\rho}_{2^{i}},i_{\alpha}}\right)\bigcap\left(\bigcap_{\beta=1}^{r}\Omega^{l_{\beta}}_{{\rho}_{2^{j}},j_{\beta}}\right)\right)\leq\left(\prod_{\alpha=1}^{r}\sigma_{\rho_{2^{i}},i_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\sigma_{\rho_{2^{j}},j_{\beta}}\right)\left(1+{\varepsilon}_{N}\right).$
+Define the events
+$A_{*}^{N}=\left\{\omega\in\Omega:\,\exists\,0<k_{1}<k_{2}<\dots<k_{r}\leq N\,\,\text{such that}\,\,\operatorname{Sep}_{N}(k_{1},\dots,k_{r})=r\,\,\text{and}\,\,\omega\in\Omega^{k_{j}}_{{\rho}_{N}}\right\},$
+$A_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}=\left\{\omega\in\Omega:\,\exists\,k_{j}\,\,\text{such that}\,\,k_{j}/N\in I_{j}\,\,\text{and }\,\,\omega\in\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right\},$
 and
-$A_{I_{1},\dots,I_{r},i_{1},\dots i_{r}}^{n}=\{\exists\,k_{j},\,\,k_{j}/n\in I_{j},\,\Omega^{k_{j}}_{{\rho}_{n},i_{j}}\text{ happens}\}.$
-We also let
-$(M4)_{*}\quad\quad\lim_{n\to\infty}\frac{\mu(A_{*}^{n})}{n^{r}\sigma_{\rho_{n}}^{r}}=1.$
-$(M4)_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}\quad\quad\lim_{n\to\infty}\frac{\mu(A_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}^{n})}{n^{r}\Pi_{j=1}^{r}\left(\sigma_{\rho_{n},i_{j}}|I_{j}|\right)}=1.$
+$H_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}=\bigcap_{M=1}^{\infty}\bigcup_{N=M}^{\infty}A_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}.$
+Additionally, assume
+$(M4)_{*}\quad\quad\limsup_{N\to\infty}\frac{\mu\left(A_{*}^{N}\right)}{N^{r}\sigma_{\rho_{N}}^{r}}\leq C$
+and
+$(M4)_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}\quad\quad\lim_{N\to\infty}\frac{\mu\left(A_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}^{N}\right)}{N^{r}\Pi_{j=1}^{r}\left(\sigma_{\rho_{N},i_{j}}|I_{j}|\right)}=1.$
 Lemma 2.2.
-Given a family of events $\{\Omega_{\rho_{n}}\}_{\rho_{n}>0}$ and their decomposition $\Omega_{\rho_{n}}=\bigcup_{i=1}^{p}\Omega_{\rho_{n},i}$, define
+Given a family of events $\{\Omega_{\rho_{N}}\}_{N=1}^{\infty}$ with decompositions $\Omega_{\rho_{N}}=\bigcup_{i=1}^{p}\Omega_{\rho_{N},i}$, define
 ${\mathbf{S}}_{r}=\sum_{j=1}^{\infty}\left(2^{j}{\sigma_{{\rho}_{2^{j}}}}\right)^{r}.$
-We assume that $\rho_{n_{1}}\leq\rho_{n_{2}}$ for $n_{1}\geq n_{2}.$
-Suppose ${\mathbf{S}}_{r}<\infty$ and $\Omega_{{\rho}_{n_{1}}}\subset\Omega_{{\rho}_{n_{2}}}$ for $n_{1}\geq n_{2}:$
-(a) If $(M1)_{r}$ and $(M2)_{r}$ are satisfied, then with probability $1$, we have that for large $n,$ the events $\Omega_{\rho_{n}}^{k}$ happen for at most $(r-1)$ times.
-(b) If $(M4)_{*}$ holds, then with probability $1$, $A^{n}_{*}$ does not occur for large $n.$
-Suppose ${\mathbf{S}}_{r}=\infty.$ For any $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r}$ and any intervals $I_{1},I_{2}\dots I_{r}\subset[0,1]:$
-If
-(c) $(M1)_{k}$, $(M2)_{k}$ for $k=r+1,\dots,2r$ as well as $(M3)_{r}$ are satisfied and $\hat{s}(n)/n<\min_{0\leq i,\,j\leq r,\,\,i\neq j}d(I_{i},I_{j}),$ where $I_{0}=0;$ or
-(d) $(M1)_{r},$ $(M3)_{r}$ are satisfied and $(M4)_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}$ holds for $\hat{s}(n)/n<\min_{0\leq i,\,j\leq r,\,\,i\neq j}d(I_{i},I_{j}),$ where $I_{0}=0,$ then with probability $1$, there are infinitely many $n$ such that for some $k_{1}(n),k_{2}(n)\dots k_{r}(n)$ with $k_{j}(n)/n\in I_{j},$ the events $\Omega^{k_{j}}_{\rho_{n},i_{j}}$ occur.
-Proof.
-Part (a) is Theorem 2.6 in [12]. We consider the case ${\mathbf{S}}_{r}<\infty$ and $(M4)_{*}$ holds. Suppose $\{\delta_{n}\}$ is a sequence such that $\rho_{n}=\delta_{2n}$ and
-$A_{*,2}^{n}=\{\exists\,0<k_{1}<k_{2}<\dots<k_{r}\leq 2n,\,\,\operatorname{Sep}(k_{1},\dots k_{r})=r,\,\,\,\Omega^{k_{j}}_{{\rho}_{n}}\text{ happens}\}.$
-We obtain that
-$\lim_{n\to\infty}\frac{\mu(A_{*,2}^{n})}{(2n)^{r}\sigma_{\delta_{2n}}^{r}}=1$
-using $(M4)_{*}.$ Since $\sigma_{\rho_{n}}=\mu(\Omega_{\rho_{n}})=\mu(\Omega_{\delta_{2n}})=\sigma_{\delta_{2n}},$
-$\lim_{n\to\infty}\frac{\mu(A_{*,2}^{n})}{n^{r}\sigma_{\rho_{n}}^{r}}=2^{r}.$
-By taking $n=2^{m},$ we have that $\mu(A_{*,2}^{2^{m}})=O\left(2^{rm}{\sigma_{{\rho}_{2^{m}}}^{r}}\right)$ and
-$\sum_{m=1}^{\infty}\mu(A_{*,2}^{2^{m}})=O\left({\mathbf{S}}_{r}\right)<\infty.$
-It implies that $A_{*,2}^{2^{m}}$ does not occurs for infinitely many $m$’s with probability 1 using the classical Borel-Cantelli Lemma.
+Assume $\rho_{N_{1}}\leq\rho_{N_{2}}$ and $\Omega_{{\rho}_{N_{1}}}\subset\Omega_{{\rho}_{N_{2}}}$ for $N_{1}\geq N_{2}.$
+Case 1: ${\mathbf{S}}_{r}<\infty.$
+  • (a)
+    If $(M1)_{r}$ and $(M2)_{r}$ hold, then with probability $1,$ the events $\{\Omega_{\rho_{N}}^{k}\}_{k=1}^{N}$ occur at most (r1)(r-1) times for all sufficiently large $N.$
+  • (b)
+    If $(M4)_{*}$ holds, then with probability 1, the event $A^{N}_{*}$ does not occur for all sufficiently large $N.$
+Case 2: ${\mathbf{S}}_{r}=\infty.$
+For any $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r}$ and any intervals $I_{1},I_{2}\dots I_{r}\subset[0,1],$
+  • (c)
+    If $(M1)_{k}$ and $(M2)_{k}$ hold for $k=r,r+1,\dots,2r,$ and $(M3)_{r}$ is satisfied under the condition $\hat{s}(N)/N<\min_{0\leq i<j\leq r}d(I_{i},I_{j})$ where $I_{0}=0,$ then $\mu\left(H_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}\right)=1.$
+  • (d)
+    If $(M3)_{r}$ is satisfied under the condition $\hat{s}(N)/N<\min_{0\leq i<j\leq r}d(I_{i},I_{j})$ where $I_{0}=0,$ and $(M4)_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}$ holds, then $\mu\left(H_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}\right)=1.$
+Proof of Lemma 2.2.
+Part (a) is Theorem 2.6 in [10].
+We now consider the case where ${\mathbf{S}}_{r}<\infty$ and $(M4)_{*}$ holds.
+Suppose $\{\delta_{N}\}$ is a sequence of positive numbers such that $\rho_{N}=\delta_{2N}.$ Define the event
+$A_{*,2}^{N}=\left\{\omega\in\Omega:\,\exists\,0<k_{1}<k_{2}<\dots<k_{r}\leq 2N,\,\operatorname{Sep}_{2N}(k_{1},\dots k_{r})=r\,\,\text{such that}\,\,\omega\in\Omega^{k_{j}}_{{\rho}_{N}}\right\}.$
+By $(M4)_{*},$ $\limsup_{N\to\infty}\frac{\mu\left(A_{*,2}^{N}\right)}{(2N)^{r}\sigma_{\delta_{2N}}^{r}}\leq C.$ Since $\sigma_{\rho_{N}}\!\!=\!\!\mu\left(\Omega_{\rho_{N}}\right)\!\!=\!\!\mu\left(\Omega_{\delta_{2N}}\right)\!\!=\!\!\sigma_{\delta_{2N}},$ we have
+$\limsup_{N\to\infty}\frac{\mu\left(A_{*,2}^{N}\right)}{N^{r}\sigma_{\rho_{N}}^{r}}\leq C.$
+By taking $N=2^{M},$ we find that $\mu\left(A_{*,2}^{2^{M}}\right)\leq C2^{rM}{\sigma_{{\rho}_{2^{M}}}^{r}}$ and thus
+$\sum_{M=1}^{\infty}\mu\left(A_{*,2}^{2^{M}}\right)\leq C{\mathbf{S}}_{r}<\infty.$
+By the classical Borel–Cantelli lemma, $A_{*,2}^{2^{M}}$ does not occurs for infinitely many $M$’s with probability 1.
 Then Part (b) follows from the fact that $A_{*}^{N}\subset A_{*,2}^{2^{M}}$ when $2^{M}\leq N\leq 2^{M+1}.$
 Part (c) is Theorem 2.12 in [10].
-Finally, part (d) holds because in the proof of part (c) in [12], $(M1)_{k}$ and $(M2)_{k}$ for $k=r+1,\dots,2r$ are only used to verify $(M4)_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}$.
-$\square$
-We now turn to this paper’s problem to consider the following targets:
-$\Omega_{\rho_{N}}=\{y:\phi(y)>\rho_{N}\}$
-and
-$\Omega_{\rho_{N},j}=\left\{y:\frac{\phi(y)}{\rho_{N}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}]\right\}$
-for $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{t},$ $t\in\mathbb{R}$ and $c_{j}>{\varepsilon}$ for any $j.$ Take $s(N)=K\operatorname{ln}N$ with $K$ being sufficiently large and $\hat{s}(N)=2{\varepsilon}N.$
+Finally, part (d) holds because in the proof of part (c) in [10], $(M1)_{k}$ and $(M2)_{k}$ for $k=r,r+1,\dots,2r$ are only used to verify $(M4)_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}$.
+∎
+We now address the central problem of this paper by analyzing the following targets:
+(2.4) $\Omega_{\rho_{N}}=\left\{y:\phi(y)>C\rho_{N}\right\}$
+with $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{t}$ where $t>0$ and $C>0;$
+(2.5) $\Omega_{\rho_{N},j}=\left\{y:\frac{\phi(y)}{\rho_{N}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}]\right\}$
+with $c_{j}>0$, $0<{\varepsilon}<c_{j}$ for non-periodic points, and $0<{\varepsilon}<\left(\gamma^{q/s}-1\right)c_{j}/\left(\gamma^{q/s}+1\right)$ for periodic points of period $q$. Also let
+(2.6) $\tilde{\Omega}_{\rho_{N},j}=\Omega_{\rho_{N},j}^{c}\mathrel{{$\bigcap$}}f^{-1}\Omega_{\rho_{N},j}^{c}\mathrel{{$\bigcap$}}\cdots\mathrel{{$\bigcap$}}f^{-(p_{0}-1)}\Omega_{\rho_{N},j}^{c}\mathrel{{$\bigcap$}}f^{-p_{0}}\Omega_{\rho_{N},j}$
+where $p_{0}=\operatorname{ln}\operatorname{ln}N.$
+Additionally, we set $s(N)=K\operatorname{ln}N$, where $K$ is chosen to be sufficiently large, and set $\hat{s}(N)=2\varepsilon N$.
 Proposition 2.3.
-(a) For any $r\in\mathbb{N},$ $\Omega_{\rho_{N}}$ and $\Omega_{\rho_{N},j}$ satisfy $(M1)_{r}$ and $(M3)_{r}.$
-(b) If $x$ is Diophantine then $(M2)_{r}$ holds.
-(c) If $x$ is non-periodic then $(M4)_{*}$ and $(M4)_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}$ are satisfied.
-The proof of Proposition 2.3 is similar to Proposition 3.9 in [12] with some modifications, which is included in Appendix B.
-Next, we will recall the exponential mixing property for expanding maps, which is useful not only in the proof Proposition 2.3 but also in the next section.
-Denote by $BV$ the space of functions with bounded variation on $\mathbb{S}$ and
-${\mathbb{E}}\psi=\int\psi{\rm{d}}\mu\,\,\text{for}\,\,\psi\in BV.$
-By Proposition 3.8 of [44], there are constants $C,$ $\theta<1$ such that
-(2.4) $\left|{\mathbb{E}}\left(\psi_{1}(\psi_{2}\circ f^{n})\right)-({\mathbb{E}}\psi_{1})(E\psi_{2})\right|\leq C\|\psi_{1}\|_{BV}\|\psi_{2}\|_{L^{1}}\theta^{n},\,n\in\mathbb{N}$
-for any $\psi_{1},\,\psi_{2}\in BV,$ where $\|\psi\|_{BV}:=\|\psi\|_{L^{1}}+\text{Var}(\psi)$ and $\text{Var}(\psi)$ is the variation of $\psi.$
-In the following, we use $C$ to denote a constant that may change from line to line but that will not depend on $N.$
+  • (a)
+    The events $\left\{\Omega_{\rho_{N}}\right\}$ and $\left\{\Omega_{\rho_{N},j}\right\}$ satisfy conditions $(M1)_{r}$ and $(M3)_{r}$ for all $r\in\mathbb{N}.$
+  • (b)
+    If $x$ is Diophantine then condition $(M2)_{r}$ holds for any $r\in\mathbb{N}.$
+  • (c)
+    Condition $(M4)_{*}$ is satisfied for $\left\{\Omega_{\rho_{N}}\right\}.$
+  • (d)
+    Suppose that $\min_{i\neq j}d(I_{i},I_{j})>{\varepsilon}$ for some ${\varepsilon}>0.$ If $x$ is not a periodic point, condition $(M4)_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}$ holds for the events $\left\{\Omega_{\rho_{N},j}\right\}.$ Moreover, if $x$ is periodic, the same condition holds when $\left\{\Omega_{\rho_{N},j}\right\}$ is replaced by the events $\left\{\tilde{\Omega}_{\rho_{N},j}\right\}$ in the definition of $A_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}.$
+The proof of Proposition 2.3 follows a similar structure to that of Proposition 3.9 in [10] with some modifications, which is provided in Appendix B.
+Next, we will recall the exponential mixing property for expanding maps, which will be useful not only in the proof Proposition 2.3 but also in the next section.
+We denote $BV$ to be the space of functions with bounded variation on $\mathbb{S},$ and ${\mathbb{E}}\psi=\int\psi{\rm{d}}\mu\,\,\,\text{for}\,\,\,\psi\in L^{1}(\mu).$ By Proposition 3.8 of [31], there exist constants $C>0,$ $\theta<1$ such that for all $\psi_{1},\,\psi_{2}\in BV,$
+(2.7) $\left|{\mathbb{E}}\left(\psi_{1}(\psi_{2}\circ f^{n})\right)-({\mathbb{E}}\psi_{1})(E\psi_{2})\right|\leq C\|\psi_{1}\|_{BV}\|\psi_{2}\|_{L^{1}}\theta^{n},\,n\in\mathbb{N},$
+where $\|\psi\|_{BV}:=\|\psi\|_{L^{1}}+\text{Var}(\psi),$ and $\text{Var}(\psi)$ denotes the variation of $\psi.$
 Lemma 2.4.
-[…]
+Suppose $\varphi_{j}\in BV$ for $1\leq j\leq q.$
   1. (i)
-    If $\varphi_{j}\geq 0$ for $1\leq j\leq q,$ we have for any $k_{1}\leq k_{2}\leq\cdots\leq k_{q},$ $k_{i}\in\mathbb{N},$
+    If $\varphi_{j}\geq 0$ for $1\leq j\leq q,$ then for any $k_{1}\leq k_{2}\leq\cdots\leq k_{q}$ with $k_{i}\in\mathbb{N},$ we have
     (2.8) $\left(\prod_{j=1}^{q-1}\left(\|\varphi_{j}\|_{L^{1}}-C\|\varphi_{j}\|_{BV}\theta^{k_{j+1}-k_{j}}\right)\right)\|\varphi_{q}\|_{L^{1}}\leq\left\|\prod_{j=1}^{q}\varphi_{j}\circ f^{k_{j}}\right\|_{L^{1}}$
@@
     If for each $j,$ $k_{j+1}-k_{j}\geq k,$ then
-    (2.6) $\left|{\mathbb{E}}\left(\prod_{j=1}^{q}(\varphi_{j}\circ f^{k_{j}})\right)-\prod_{j=1}^{q}{\mathbb{E}}\varphi_{j}\right|\leq C\theta^{k}\prod_{j=1}^{q}\|\varphi_{j}\|_{BV}.$
+    (2.9) $\left|{\mathbb{E}}\left(\prod_{j=1}^{q}\varphi_{j}\circ f^{k_{j}}\right)-\prod_{j=1}^{q}{\mathbb{E}}\varphi_{j}\right|\leq C\theta^{k}\prod_{j=1}^{q}\|\varphi_{j}\|_{BV}.$
 Lemma 2.4 could be proved by the arguments of Proposition 6.2 in [10] with $\|\cdot\|_{Lip}$ replaced by $\|\cdot\|_{BV}.$ However, in the present case a simpler argument is available and we provide it in Appendix A for the convenience of the reader.
@@
 $S_{n}^{\prime}=\sum_{1\leq k\leq n}\left(X_{k}{\mathbf{1}}_{\{X_{k}\geq N^{1/s}(\operatorname{ln}N)^{\delta}\}}-b_{N}^{\prime}\right),\quad S_{n}^{\prime\prime}=\sum_{1\leq k\leq n}\left(X_{k}{\mathbf{1}}_{\{X_{k}<N^{1/s}(\operatorname{ln}N)^{\delta}\}}-b_{N}^{\prime\prime}\right),$
-where $b_{N}^{\prime}=b_{N}^{\prime\prime}=0$ if $s<1$ and $b_{N}^{\prime},b_{N}^{\prime\prime}$ are chosen to make the expectations of the LHS’s zero if $s>1.$
+where $b_{N}^{\prime}=b_{N}^{\prime\prime}=0$ if $s\leq 1$ and $b_{N}^{\prime},b_{N}^{\prime\prime}$ are chosen to so that the expectations of the left-hand sides are zero when $s>1.$
 Lemma 2.5.
-If $\delta<\alpha$ and either $10<s<$ 1 or $21<s<$ 2 and $x$ is Diophanitne then
-[…]
+If $\delta<\alpha$ and either $0<s<1$ or $1\leq s<2$ and $x$ is Diophanitne then
+$\max_{1\leq n\leq N}\frac{\left|S_{n}^{\prime\prime}\right|}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\rightarrow 0\quad\text{a.s.}.$
 The proof of Lemma 2.5 is the main part of this paper and is contained in the next section. Given Lemma 2.5, we proceed to show Theorem 2.1.
@@
 $W_{N}^{\prime}(t)=\frac{S_{Nt}^{\prime}}{N^{1/s}(\operatorname{ln}N)^{\alpha}}.$
-By Lemma 2.5, the limit points of $W_{N}(\cdot)$ and $W_{N}^{\prime}(\cdot)$ are the same.
-Consider the targets $\Omega_{\rho_{N}}=\{y:\phi(y)>\rho_{N}\}$ for $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{\delta}$ and suppose that $x$ is non-periodic.
-By Proposition 2.3 and Lemma 2.2(b), for large $N$ there are at most $\left[(\delta s)^{-1}\right]$ intervals $I_{j}$ of length $2s(N)$ such that all numbers $n$ such that $X_{n}>N^{1/s}(\operatorname{ln}N)^{\delta}$ are contained in one of those intervals.
-Moreover taking ${\bar{\alpha}}$ such that $((r+1)s)^{-1}<{\bar{\alpha}}<\alpha,$ we see that among those intervals there are at most $r$ which contain $n^{\prime}$s with $X_{n}\geq N^{1/s}(\operatorname{ln}N)^{{\bar{\alpha}}}.$ We claim that the contribution of the remaining intervals is negligible for a.e.
-x.
-Indeed suppose there is an interval $I$ of length $2s(N)$ and ${\bar{n}}\in I$ such that
+By Lemma 2.5, the limit points of $W_{N}(\cdot)$ and $W_{N}^{\prime}(\cdot)$ are the same. Consider the targets
+$\Omega_{\rho_{N}}=\{y:\phi(y)>\rho_{N}\},$
+where $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{\delta}.$ By Proposition 2.3(c) and Lemma 2.2(b), for sufficiently for large $N,$ there exist at most $\left[(\delta s)^{-1}\right]$ intervals $I_{j}$ of length $2s(N)$ such that every $n$ satisfying $X_{n}>N^{1/s}(\operatorname{ln}N)^{\delta}$ is contained in one of those intervals.
+Moreover, by choosing $\bar{\alpha}$} such that $((r+1)s)^{-1}<{\bar{\alpha}}<\alpha,$ we see that among those intervals, at most $r$ contain $n$ for which
+$X_{n}\geq N^{1/s}(\operatorname{ln}N)^{{\bar{\alpha}}}.$
+We claim that the contribution of the remaining intervals is negligible for almost every $y.$ Indeed, suppose there is an interval $I$ of length $2s(N)$ and ${\bar{n}}\in I$ such that
 (2.10) $\max_{n\in I}X_{n}=X_{\bar{n}}\leq N^{1/s}(\operatorname{ln}N)^{\bar{\alpha}}\;\;\mathrm{and}\;\;\sum_{n\in I}X_{n}\geq{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}.$
@@
 $\sum_{n\in I}X_{n}\leq\sum_{n\in I^{\prime}}X_{n}+\sum_{n\in I^{\prime\prime}}X_{n}\leq N^{1/s}(\operatorname{ln}N)^{(\alpha+{\bar{\alpha}})/2}+2Y_{I}s(N).$
-By the definition of $I$ we have that $Y_{I}\geq{\bar{\varepsilon}}N^{1/s}(\operatorname{ln}N)^{\alpha-1}$ for some ${\bar{\varepsilon}}={\bar{\varepsilon}}({\varepsilon})>0.$ Accordingly if the interval $I$ satisfying (2.7), then there are two moments ${\bar{n}},{\hat{n}}\in[1,N]$ such that
+By the definition of $I,$ we have
+$Y_{I}\geq{\bar{\varepsilon}}N^{1/s}(\operatorname{ln}N)^{\alpha-1}$
+for some ${\bar{\varepsilon}}={\bar{\varepsilon}}({\varepsilon})>0.$ Accordingly, if the interval $I$ satisfies (2.10), then there exist two moments ${\bar{n}},{\hat{n}}\in[1,N]$ such that
 (2.11) $(\operatorname{ln}N)^{(\alpha-{\bar{\alpha}})/2}\leq|{\bar{n}}-{\hat{n}}|\leq 2s(N),\quad\min\left\{X_{\bar{n}},\,X_{\hat{n}}\right\}>{\bar{\varepsilon}}N^{1/s}(\operatorname{ln}N)^{\alpha-1}.$
-Thus the lemma below shows that the contribution of the intervals satisfying (2.7) is negligible. Its proof is available in Appendix B.
+Thus, the following lemma establishes that the contribution of the intervals satisfying (2.10) is negligible. Its proof is provided in Appendix B.
 Lemma 2.6.
-For a.e. $y$, (2.8) happens only for finitely many $N.$
+For almost every $y,$ (2.11) happens only for finitely many $N.$
 It then follows that the limit points of $W_{N}^{\prime}(\cdot)$ have at most $r$ positive jumps. That is, all limit points are in ${\mathbb{H}}_{r}.$
-To see that all points in ${\mathbb{H}}_{r}$ are realized, fix $r$-tuples $(c_{1},\dots,c_{r})$ and $(t_{1},\dots t_{r})$ such that $c_{j}>0$ for all $j$ and fix $0<{\varepsilon}<c_{j}.$ Consider the events
-$\Omega_{\rho_{N},j}=\left\{y:\frac{\phi(y)}{\rho_{N}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}]\right\}.$
-Taking $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{\alpha}$ and applying Proposition 2.3 and Lemma 2.2(d), we see that with probability $1$ there is a subsequence $N_{k}$ and times $n_{1}(k),n_{2}(k),\dots,n_{r}(k)$ such that
-$\frac{n_{j}(k)}{N_{k}}\in[t_{j}-{\varepsilon},t_{j}+{\varepsilon}]\quad\text{and}\quad\frac{X_{n_{j}(k)}(y)}{N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}].$
-It follows that for each $W\in{\mathbb{H}}_{r}$, $W$ belongs to the limit set with probability $1$ by taking a countable set of ${\varepsilon}_{l}$ converging to $0$ Recall the fact that two elements $W_{1}$ and $W_{2}$ of ${\mathbb{H}}_{r}$ with exactly $r$ jumps are $J_{1}$ close if and only if $\{c_{m}^{1}\}_{m=1}^{r}$ is close to $\{c_{m}^{2}\}_{m=1}^{r}$ and $\{t_{m}^{1}\}_{m=1}^{r}$ is close to $\{t_{m}^{2}\}_{m=1}^{r}$.
+To verify that all points in ${\mathbb{H}}_{r}$ are realizable, we fix an $r$-tuple $(c_{1},\dots,c_{r})$ such that $c_{j}>0$ for all $j,$ together with an increasing $r$-tuple $(t_{1},\dots,t_{r})$ satisfying $t_{1}<t_{2}<\cdots<t_{r}.$
+If $x$ is not a periodic point, we define the events
+$\Omega_{\rho_{N},j}=\left\{y:\frac{\phi(y)}{\rho_{N}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}]\right\},\,\,1\leq j\leq r$
+for $0<{\varepsilon}<c_{j}.$
+If $x$ is a periodic point of period $q,$ we instead choose $0<{\varepsilon}<\left(\gamma^{q/s}-1\right)c_{j}/\left(\gamma^{q/s}+1\right)$ and consider the events
+$\tilde{\Omega}_{\rho_{N},j}=\Omega_{\rho_{N},j}^{c}\mathrel{{\bigcap}}f^{-1}\Omega_{\rho_{N},j}^{c}\mathrel{{\bigcap}}\cdots\mathrel{{\bigcap}}f^{-(p_{0}-1)}\Omega_{\rho_{N},j}^{c}\mathrel{{\bigcap}}f^{-p_{0}}\Omega_{\rho_{N},j},\,\,1\leq j\leq r,$
+where $p_{0}=\operatorname{ln}\operatorname{ln}N.$
+Taking $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{\alpha}$ and applying Proposition 2.3 (d) and Lemma 2.2 (d), we conclude that with probability $1$ there is a subsequence $N_{k}$ and times $n_{1}(k),n_{2}(k),\dots,n_{r}(k)$ such that
+$\frac{n_{j}(k)}{N_{k}}\in[t_{j}-2{\varepsilon},t_{j}+2{\varepsilon}]\quad\text{and}\quad\frac{X_{n_{j}(k)}(y)}{N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}].$
+It follows that for each $\in{\mathbb{H}}_{r}, WW$ belongs to the limit set with probability $1$ by taking a countable set of ${\varepsilon}_{l}$ converging to $0.$ Recall the fact that two elements $W_{1}$ and $W_{2}$ of ${\mathbb{H}}_{r}$ with exactly $r$ jumps are $J_{1}$ close if and only if $\{c_{m}^{1}\}_{m=1}^{r}$ is close to $\{c_{m}^{2}\}_{m=1}^{r}$ and $\{t_{m}^{1}\}_{m=1}^{r}$ is close to $\{t_{m}^{2}\}_{m=1}^{r}$.
 Taking a countable dense subset in ${\mathbb{H}}_{r}$ completes the proof.
@@
 $\bar{X}_{k}=X_{k}{\mathbf{1}}_{\{X_{k}<N^{1/s}(\operatorname{ln}N)^{-D}\}}-\bar{b}_{N}\,\,\,\text{and}\,\,\,\bar{\bar{X}}_{k}=X_{k}{\mathbf{1}}_{\{N^{1/s}(\operatorname{ln}N)^{-D}\leq X_{k}<N^{1/s}(\operatorname{ln}N)^{\delta}\}}-\bar{\bar{b}}_{N},$
-where, as before, $\bar{b}_{N}=\bar{\bar{b}}_{N}=0$ for $s\leq 1$ and they are chosen to make the expectations of LHS’s zero for $s>1.$
+where, as before, $\bar{b}_{N}=\bar{\bar{b}}_{N}=0$ for $s\leq 1$ and they are chosen so that the expectations of the left-hand sides are zero for $s>1.$
 We will show that
-[…]
+(3.3) $\max_{1\leq n\leq N}\frac{|{\bar{S}}_{n}|}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\to 0\;\;\text{a.s.}$
 and
-[…]
-In § 3.2 we consider the case $s<1$ where the argument is much simpler since the sums ${\bar{S}}_{n}$ and ${\bar{{\bar{S}}}}_{n}$ are almost monotone.
-The case $21\leq s<$ 2 is more involved and requires the use of the reflection principle (see equations (3.14) and (3.23)).
-In § 3.3 and § 3.4 we prove (3.3) and (3.4) respectively for $21\leq s<$ 2 .
+(3.4) $\max_{1\leq n\leq N}\frac{|{\bar{{\bar{S}}}}_{n}|}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\to 0\;\;\text{a.s.}.$
+In § 3.2, we consider the case $s<1,$ where the argument is much simpler, since the sums ${\bar{S}}_{n}$ and ${\bar{{\bar{S}}}}_{n}$ are almost monotone.
+The case $1\leq s<2$ is more involved and requires the application of the reflection principle (see equations (3.14) and (3.23)).
+In § 3.3 and § 3.4, we establish (3.3) and (3.4), respectively, for $1\leq s<2.$
 3.2. The case of infinite expectation.
 Here we prove (3.3) and (3.4) in case $s<1.$
-We begin with (3.3). By (2.2) we have
-$E\left|\bar{S}_{N}\right|\leq N\int_{0}^{N^{1/s}(\operatorname{ln}N)^{-D}}\frac{a_{s}{\rm{d}}t}{t^{s}}+LN\leq CN^{1/s}(\operatorname{ln}N)^{-D(1-s)}.$
-Hence
+We begin by considering (3.3). By (2.1), we have
+$E\left|\bar{S}_{N}\right|\leq N\int_{0}^{N^{1/s}(\operatorname{ln}N)^{-D}}\frac{a_{s}{\rm{d}}t}{t^{s}}+LN\leq CN^{1/s}(\operatorname{ln}N)^{-D(1-s)},$
+for $L=\max_{x\in\mathbb{S}}\left|\psi(x)\right|.$ Hence
 $E\left|\frac{\bar{S}_{N}}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\right|\leq\frac{C}{(\operatorname{ln}N)^{D(1-s)+\alpha}}.$
-Take $N_{k}=2^{k}$, then for each $\varepsilon>0$,
+Let $N_{k}=2^{k}.$ For each $\varepsilon>0$, we have
 (3.5) $\sum_{k\geq 1}\mu\left(\left|\frac{\bar{S}_{N_{k}}}{N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}}\right|>\varepsilon\right)\leq\sum_{k\geq 1}\frac{C}{\varepsilon k^{D(1-s)+\alpha}}<\infty$
-where the convergence follows from (3.1). Hence Borel-Cantelli Lemma gives
-[…]
-Next, consider $2^{k}<N\leq 2^{k+1}.$ Since $|\bar{S}_{N}|\leq|\bar{S}_{N_{k+1}}|+2LN_{k+1}$ and $s>1$, (3.6) gives
-[…]
-proving (3.3).
-Next, we prove (3.4). Choose $D$ such that $D=q\delta$ where $q$ is a positive integer. For $j=0,-1,\cdots,-q,$ let
-$X_{k,j}=X_{k}{\mathbf{1}}_{\{X_{k}\in(N^{1/s}(\operatorname{ln}N)^{j\delta},N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}]\}},\quad{\bar{{\bar{S}}}}_{n,j}=\sum_{k=1}^{n}X_{k,j},$
+where the convergence follows from (3.1). By the classical Borel-Cantelli lemma
+(3.6) $\frac{|\bar{S}_{N_{k}}|}{N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}}\rightarrow 0\quad\text{a.s.}.$
+Next, consider the case $2^{k}<N\leq 2^{k+1}.$ Since $|\bar{S}_{N}|\leq|\bar{S}_{N_{k+1}}|+2LN_{k+1}$ and $s<1$, (3.6) yields
+(3.7) $\frac{|\bar{S}_{N}|}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\rightarrow 0\quad\text{ a.s.}$
+which completes the proof of (3.3).
+$\mu({\hat{B}}_{N,j})\sim 2a^{s}h(x)N^{-1}(\operatorname{ln}N)^{-j\delta s}.$
+Next, we prove (3.4). Let $D$ be chosen such that $D=q\delta,$ where $q$ is a positive integer. For $j=0,-1,\cdots,-q,$ define
+(3.8) ${\hat{B}}_{N,j}=\left\{y:\phi(y)\in\left(N^{1/s}(\operatorname{ln}N)^{j\delta},\;N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}\right]\right\},$
+and
+(3.9) ${\bar{{\bar{S}}}}_{n,j}=\sum_{k=1}^{n}\phi(f^{k}y){\mathbf{1}}_{{\hat{B}}_{N,j}}(f^{k}y),$
 where $\mathbf{1}$}. is the indicator function. It suffices to show that for each $j$,
 (3.10) $\max_{1\leq n\leq N}\frac{|{\bar{{\bar{S}}}}_{n,j}|}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\to 0.$
-Since $|{\bar{{\bar{S}}}}_{n,j}|\leq|{\bar{{\bar{S}}}}_{N,j}|+2LN$ it suffices to prove that
+Since $|{\bar{{\bar{S}}}}_{n,j}|\leq|{\bar{{\bar{S}}}}_{N,j}|+2LN$ and $s<1,$ it suffices to prove that
 (3.11) $\frac{|{\bar{{\bar{S}}}}_{N,j}|}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\to 0.$
-Next, we claim that for each $m$ there is a constant $C_{m}$ such that
-(3.10) $E\left({\bar{{\bar{S}}}}_{N,j}^{m}\right)\leq C_{m}N^{m/s}(\operatorname{ln}N)^{\delta m}.$
-Let ${\hat{B}}_{N,j}=\left\{y:\phi(y)\in\left(N^{1/s}(\operatorname{ln}N)^{j\delta},\;N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}\right]\right\},$ $j=0,-1,\cdots,-q.$ Note that
-$\frac{{\bar{{\bar{S}}}}_{N,j}}{N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}}\leq\sum_{n=1}^{N}{\mathbf{1}}_{{\hat{B}}_{N,j}}(f^{n}y).$
-Thus
-(3.11) $E\left(\left(\frac{\bar{\bar{S}}_{N,j}}{N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}}\right)^{m}\right)\leq\sum_{1\leq k_{1}\leq\cdots\leq k_{m}\leq N}\mu\left(\bigcap_{j=1}^{m}\left\{f^{k_{j}}y\in{\hat{B}}_{N,j}\right\}\right).$
 The next result is proven in Appendix B.
 Lemma 3.1.
-There is a constant $C$ such that
-$\sum_{1\leq k_{1}\leq\cdots\leq k_{m}\leq N}\mu\left(\bigcap_{j=1}^{m}\left\{f^{k_{j}}y\in{\hat{B}}_{N,j}\right\}\right)\leq C\left(N\mu\left({\hat{B}}_{N,j}\right)\right)^{m}.$
-Combining the lemma with (2.2) we see that (3.11) is less than $C(\operatorname{ln}N)^{-j\delta sm}.$ Since $j\leq 0$ and $s<1,$ (3.10) follows.
-By Markov inequality
-$\mu\left(|{\bar{{\bar{S}}}}_{N,j}|\geq{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\frac{C}{{\varepsilon}^{m}(\operatorname{ln}N)^{(\alpha-\delta)m}}$
+  • (a)
+    If $x$ is not a periodic point then ${\mathbb{E}}{\bar{{\bar{S}}}}_{N,j}^{m}\leq CN^{m/s}(\operatorname{ln}N)^{\delta m}.$
+  • (b)
+    If $x$ is periodic then ${\mathbb{E}}{\bar{{\bar{S}}}}_{N,j}^{m}\leq CN^{m/s}(\operatorname{ln}N)^{\delta m}\left(\operatorname{ln}\operatorname{ln}N\right)^{m}.$
+By Markov’s inequality
+$\mu\left(|{\bar{{\bar{S}}}}_{N,j}|\geq{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\frac{C\left(\operatorname{ln}\operatorname{ln}N\right)^{m}}{{\varepsilon}^{m}(\operatorname{ln}N)^{(\alpha-\delta)m}}$
 for even $m.$ Taking $m$ such that $m(\alpha-\delta)>1$ and arguing as in the proof of (3.7) we obtain
-[…]
+(3.12) $\frac{|\bar{\bar{S}}_{N,j}|}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\rightarrow 0\quad\text{a.s.}.$
 Since $j$ is arbitrary, we obtain (3.4) for $s<1.$
-3.3. Controlling small values of $X$ for $21\leq s<$ 2
-Here we prove (3.3) in the case $21\leq s<$ 2 .
-Given $y\in\mathbb{S}$, let $C_{n}(y)$ be the continuity component of $\phi 1_{\phi\leq N^{1/s}(\operatorname{ln}N)^{-D}}\circ f^{k},$ $k\leq n$ containing $y$ and $f^{n}|_{C_{n}(y)}$ be a diffeomorphism.
-Let $I_{n}(y)=f^{n}C_{n}(y).$ We say that $C_{n}(y)$ is short if the length of $I_{n}(y)$ is less than $N^{-1}(\operatorname{ln}N)^{-100}$ and we say that it is long otherwise.
-Fix $\varepsilon>0,$ we say that $C_{n}$ is extremal if there exists $z\in C_{n}$ such that $|{\bar{S}}_{n}(z)|\geq{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}$ and for each $k<n,$ $|{\bar{S}}_{k}(z)|<{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}$ on $C_{n}.$ Note that different extremal components are disjoint.
-In addition,
-(3.13) $|{\bar{S}}_{n}(y)|\geq\frac{3}{4}{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}\quad\text{for all }y\text{ in }C_{n}.$
+3.3. Controlling small values of $X$ for $1\leq s<2$
+We now prove (3.3) for $1\leq s<2.$
+Given $y\in\mathbb{S},$ let $C_{n}(y)$ be the continuity component of $\phi 1_{\phi\leq N^{1/s}(\operatorname{ln}N)^{-D}}\circ f^{k}$ for all $k\leq n,$ containing $y.$ Then $f^{n}|_{C_{n}(y)}$ be a diffeomorphism.
+Define $I_{n}(y)=f^{n}C_{n}(y).$ We say that $C_{n}(y)$ is short if the length of $I_{n}(y)$ is less than $N^{-1}(\operatorname{ln}N)^{-100},$ and long otherwise.
+Fix $\varepsilon>0.$ We say that $C_{n}$ is extremal if there exists $z\in C_{n}$ such that
+$\left|{\bar{S}}_{n}(z)\right|\geq{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}$
+and for each $k<n,$ $\left|{\bar{S}}_{k}(\cdot)\right|<{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}$ on $C_{n}.$ Note that different extremal components are disjoint. Moreover, we have
+(3.13) $\left|{\bar{S}}_{n}(y)\right|\geq\frac{3}{4}{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}\quad\text{for all }y\text{ in }C_{n}.$
 Indeed, fix a large $K$ and write
-$\left|{\bar{S}}_{n}(y)-{\bar{S}}_{n}(z)\right|\leq\sum_{k=1}^{n-K\operatorname{ln}N}|{\bar{X}}_{k}(y)-{\bar{X}}_{k}(z)|+|{\bar{S}}_{K\operatorname{ln}N}(f^{n-K\operatorname{ln}N}y)|+|{\bar{S}}_{K\operatorname{ln}N}(f^{n-K\operatorname{ln}N}z)|,$
-then the last two terms are smaller than $2KN^{1/s}(\operatorname{ln}N)^{1-D}$ due to the cutoff in the definition of ${\bar{S}}_{n}.$ To estimate the first term, we note that since $d(f^{n}y,f^{n}z)\leq 1$ and $f$ is expanding then for $k\leq n-K\operatorname{ln}N,$ $d(f^{k}y,f^{k}z)\leq\theta^{K\operatorname{ln}N}.$ Since $\phi{\mathbf{1}}_{\{\phi<N^{1/s}(\operatorname{ln}N)^{-D}\}}$ is Lipschitz on its continuity components with Lipschitz constant less than $CN^{1+1/s}(\operatorname{ln}N)^{-(1+s)D}$, it follows that if $K$ is sufficiently large then $|{\bar{X}}_{k}(y)-{\bar{X}}_{k}(z)|\leq N^{-100}$, proving (3.13).
+$\left|{\bar{S}}_{n}(y)-{\bar{S}}_{n}(z)\right|\leq\sum_{k=1}^{n-K\operatorname{ln}N}\left|{\bar{X}}_{k}(y)-{\bar{X}}_{k}(z)\right|+\left|{\bar{S}}_{K\operatorname{ln}N}(f^{n-K\operatorname{ln}N}y)\right|+\left|{\bar{S}}_{K\operatorname{ln}N}(f^{n-K\operatorname{ln}N}z)\right|.$
+The last two terms are smaller than $2KN^{1/s}(\operatorname{ln}N)^{1-D}$ due to the cutoff in the definition of ${\bar{S}}_{n}.$ To estimate the first term, observe that since $d(f^{n}y,f^{n}z)\leq 1$ and $f$ is expanding, for $k\leq n-K\operatorname{ln}N,$ we have $d(f^{k}y,f^{k}z)\leq\theta^{K\operatorname{ln}N}.$ Since $\phi{\mathbf{1}}_{\{\phi<N^{1/s}(\operatorname{ln}N)^{-D}\}}$ is Lipschitz on its continuity components with a Lipschitz constant less than $CN^{1+1/s}(\operatorname{ln}N)^{-(1+s)D}$, it follows that if $K$ is sufficiently large, then
+$\left|{\bar{X}}_{k}(y)-{\bar{X}}_{k}(z)\right|\leq N^{-100},$
+which proves (3.13).
 Let ${\mathcal{E}}_{N}$ be the set of points which belong to some extremal component for $n\leq N.$
 Lemma 3.2.
-If $D$ is large enough then $\mu({\mathcal{E}}_{N})\leq{C}{(\operatorname{ln}N)^{-2}}.$
-Lemma 3.2 implies (3.3).
-In fact, combining this lemma with Borel-Cantelli Lemma we conclude that with probability $1$, for large $k$, $|{\bar{S}}_{n}|<{\varepsilon}N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}$ for $n\leq N_{k},$ $N_{k}=2^{k}.$ Since ${\varepsilon}$ is arbitrary, (3.3) follows.
+If $D$ is is sufficiently large, then $\mu({\mathcal{E}}_{N})\leq{C}{(\operatorname{ln}N)^{-2}}.$
+Lemma 3.2 implies (3.3). In fact, combining this lemma with the classical Borel-Cantelli lemma we conclude that with probability $1$, for sufficiently large $k$, we have
+$\left|{\bar{S}}_{n}\right|<{\varepsilon}N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}\,\,\,\text{for}\,\,\,n\leq N_{k},\,\,\,N_{k}=2^{k}.$
+Since $\varepsilon$} is arbitrary, (3.3) follows.
 Proof of Lemma 3.2.
@@
 for some $C>0$ independent of $N.$
-To prove (3.15) we use Chebyshev inequality. By (2.2) we have
+To prove (3.15) we use Chebyshev’s inequality. By (2.1) we have
 (3.16) ${\rm Var}({\bar{X}}_{n})\leq CN^{2/s-1}(\operatorname{ln}N)^{{-D(2-s)}}.$
@@
 (3.18) ${\rm Var}({\bar{S}}_{N})\leq C\left(N^{2/s}(\operatorname{ln}N)^{{1-D(2-s)}}+1\right).$
-Since $D(2-s)+2\alpha>3$ due to (3.2), (3.15) follows by Chebyshev inequality.
-The proof of (3.14) follows similar ideas. Namely, in view of (3.13), it suffices to show that there exists $C<1$ such that
-(3.19) $\mu\left(y\in C_{n}:\left|\sum_{k=n+1}^{N}{\bar{X}}_{k}(y)\right|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq C\mu(C_{n}).$
+Since $D(2-s)+2\alpha>3$ due to (3.2), (3.15) follows by Chebyshev’s inequality.
+The proof of (3.14) follows similar ideas. Namely, in view of (3.13), it suffices to show that there exists $\tilde{C}<1$ such that
+(3.19) $\mu\left(y\in C_{n}:\left|\sum_{k=n+1}^{N}{\bar{X}}_{k}(y)\right|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\tilde{C}\mu(C_{n}).$
 Since $f^{n}$ has bounded distortion property, (3.19) follows from
-(3.20) $\mu\left(z\in I_{n}:|{\bar{S}}_{N-n}(z)|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq C\mu(I_{n}).$
+(3.20) $\mu\left(z\in I_{n}:|{\bar{S}}_{N-n}(z)|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\tilde{C}\mu(I_{n}).$
 Split ${\bar{S}}_{N-n}={\bar{S}}_{K\operatorname{ln}N}+({\bar{S}}_{N-n}-{\bar{S}}_{K\operatorname{ln}N}).$ The first term is bounded by
@@
 (3.22) $\mu\left(z\in I_{n}:\left|\sum_{k=K\operatorname{ln}N}^{N-n}{\bar{X}}_{k}(z)\right|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\frac{16C\mu(I_{n})}{{\varepsilon}^{2}}(\operatorname{ln}N)^{1-D(2-s)-2\alpha}.$
-Combining (3.21) and (3.22) we obtain (3.14). $\square$
-3.4. Controlling intermediate values of $X$ for $21\leq s<$ 2
-Here we prove (3.4) in case $21\leq s<$ 2 .
+Combining (3.21) and (3.22) we obtain (3.14). ∎
+3.4. Controlling intermediate values of $X$ for $1\leq s<2$
+Here we prove (3.4) in case $1\leq s<2.$
 The proof of (3.4) for $s\geq 1$ will proceed similarly to §3.3.
@@
 Namely it is not true that each summand in ${\bar{{\bar{S}}}}_{K\operatorname{ln}N}(y)$ is smaller than $N^{1/s}(\operatorname{ln}N)^{-D}.$ However, we claim that for almost every $y,$
-[…]
+(3.25) $\sharp\left\{k:\,1\leq k\leq K\operatorname{ln}N,\,\,{\bar{{\bar{X}}}}_{k}(y)\geq N^{1/s}(\operatorname{ln}N)^{-D}\right\}\leq(\operatorname{ln}N)^{(\alpha-\delta)/2}$
 for $N$ large enough. If (3.25) is not satisfied, then
@@
 which is sufficient to obtain (3.13).
-Equation (3.24) follows from the Markov inequality and the following estimate.
+Equation (3.24) follows from the Markov’s inequality and the following estimate.
 Lemma 3.3.
-Suppose $x$ is Diophantine. For each $m,$ we have
+Suppose $x$ is Diophantine. For each positive integer $m,$ we have
 (3.28) $E\left({\bar{{\bar{S}}}}_{N}^{m}\right)\leq C_{m}N^{m/s}(\operatorname{ln}N)^{\delta m}.$
@@
 $\bar{\bar{X}}_{n}=X_{n}{\mathbf{1}}_{\{N^{1/s}(\operatorname{ln}N)^{-D}\leq X_{n}<N^{1/s}(\operatorname{ln}N)^{\delta}\}}-\bar{\bar{b}}_{N},\quad\bar{\bar{b}}_{N}=E\left(\phi{\mathbf{1}}_{\{N^{1/s}(\operatorname{ln}N)^{-D}\leq\phi<N^{1/s}(\operatorname{ln}N)^{\delta}\}}\right)$
-for $21<s<$ 2 and
+for $1<s<2$ and
 $\bar{\bar{X}}_{n}=X_{n}{\mathbf{1}}_{\{N^{1/s}(\operatorname{ln}N)^{-D}\leq X_{n}<N^{1/s}(\operatorname{ln}N)^{\delta}\}}$
@@
 where $c_{m_{1},\dots,m_{r}}$ are combinatorial factors.
-Note that by (2.2)
-(3.30) ${\mathbb{E}}\left|\bar{\bar{X}}_{0}\right|^{l}\leq\left\{\begin{array}[]{ll}CN^{l/s-1}(\operatorname{ln}N)^{(s-l)D},&l<s,\\ CN^{l/s-1}(\operatorname{ln}N)^{(l-s)\delta},&l>s,\\ C\operatorname{ln}\operatorname{ln}N,&l=s,\\ \end{array}\right.$
+Note that by (2.1),
+[…]
 and
 $\left\|\bar{\bar{X}}^{l}_{0}\right\|_{BV}\leq CN^{l/s}(\operatorname{ln}N)^{l\delta}.$
-(i) Suppose that $\operatorname{Sep}(n_{1},\dots,n_{k})=k.$ Then we use the multiple mixing bound of Lemma 2.4(ii).
-We take $K$ in the definition of separation condition so large that the error term in Lemma 2.4(ii) is $O(N^{-k}).$ There are two further subcases.
-(a) $\min_{j}m_{j}>1.$ Using the second estimate in (3.30) we see that
+(i) Suppose that $\operatorname{Sep}_{N}(n_{1},\dots,n_{k})=k.$ In this case, we apply the multiple mixing bound from Lemma 2.4(ii).
+We choose $K$ in the definition of separation condition large enough so that the error term in Lemma 2.4(ii) is $O(N^{-k}).$ This leads to two further subcases.
+(a) $\min_{j}m_{j}>1.$ Using the second estimate in (3.30), we find that
 $\prod_{j}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)\leq CN^{m/s-k}(\operatorname{ln}N)^{m\delta},$
-while the error term is much smaller. Summing over all $n_{1},\dots,n_{k}$ we obtain this case gives the leading contribution of $CN^{m/s}(\operatorname{ln}N)^{m\delta}.$
-(b) $\min_{j}m_{j}=1.$ If $1<s<2,$ then
-$\prod_{j}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)=0,$
-so each term is $O(N^{-k})$ and the total contribution of all terms is $O(1).$
-If $s=1,$
+while the error term is much smaller. Summing over all $n_{1},\dots,n_{k},$ we see that this case gives the leading contribution of $CN^{m/s}(\operatorname{ln}N)^{m\delta}.$
+(b) $\min_{j}m_{j}=1.$ If $1<s<2,$ then $\prod_{j}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)=0,$ so each term is $O(N^{-k}),$ and the total contribution of all terms is $O(1).$
+If $s=1,$ we have
 $\prod_{j}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)=\left(\prod_{m_{j}\geq 2}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)\right)\left(\prod_{m_{j}=1}{\mathbb{E}}\bar{\bar{X}}_{n_{j}}\right)$
-$\leq$ $C\left(\prod_{m_{j}\geq 2}N^{m_{j}-1}\left(\operatorname{ln}N\right)^{(m_{j}-1)\delta}\right)\left(\operatorname{ln}\operatorname{ln}N\right)^{m-k}\leq CN^{m-k}\left(\operatorname{ln}N\right)^{(m-k)\delta}\left(\operatorname{ln}\operatorname{ln}N\right)^{m}$
+This leads to
+$\prod_{j}{\mathbb{E}}\left(\bar{\bar{X}}_{n_{j}}^{m_{j}}\right)\leq C\left(\prod_{m_{j}\geq 2}N^{m_{j}-1}\left(\operatorname{ln}N\right)^{(m_{j}-1)\delta}\right)\left(\operatorname{ln}\operatorname{ln}N\right)^{m-k}\leq CN^{m-k}\left(\operatorname{ln}N\right)^{(m-k)\delta}\left(\operatorname{ln}\operatorname{ln}N\right)^{m}$
 for $k<m$ and
@@
 for $k=m.$ It follows that the total contribution of all terms is bounded by $CN^{m}(\operatorname{ln}N)^{m\delta}$ when $\min_{j}m_{j}=1$ and $s=1.$
-(ii) Suppose that $\operatorname{Sep}(n_{1},\dots,n_{k})=\bar{k}<k.$ Given that $x$ is Diophantine, we consider the case where
+(ii) Suppose that $\operatorname{Sep}_{N}(n_{1},\dots,n_{k})=\bar{k}<k.$ Given that $x$ is Diophantine, we consider the case where
 $\min_{j}\{n_{j+1}-n_{j}\}\geq\epsilon\operatorname{ln}N$
-for some $\epsilon>0.$ Otherwise, the contribution to the sum is $0$ By (3.30),
+for some $\epsilon>0.$ Otherwise, the contribution to the sum is $0.$ By (3.30),
 ${\mathbb{E}}\left|\bar{\bar{X}}_{n_{j}}\right|^{m_{j}}\leq CN^{m_{j}/s-1}(\operatorname{ln}N)^{m_{j}D}.$
-In addition using Lemma 2.4(i), we see that $E(\bar{\bar{X}}_{n_{1}}^{m_{1}}\dots\bar{\bar{X}}_{n_{k}}^{m_{k}})$ is $O\left(N^{(m/s)-\bar{k}-\eta}(\operatorname{ln}N)^{Dm}\right),$ where $\eta=(k-\bar{k})\epsilon|\operatorname{ln}\theta|.$ Since the number of such terms in (3.29) is $O\left(N^{\bar{k}}(\operatorname{ln}N)^{k-\bar{k}}\right),$ we conclude that the contribution of these non-separated terms is $O\left(N^{m/s-\eta}(\operatorname{ln}N)^{(D+1)m}\right)$, which is much smaller than the RHS of (3.28).
-$\square$
-The proof of (3.23) is similar to the proof of (3.14). Namely, as in Section 3.3, it suffices to have the following analogue of (3.20):
+Additionally, using Lemma 2.4(i), we find that
+$E(\bar{\bar{X}}_{n_{1}}^{m_{1}}\dots\bar{\bar{X}}_{n_{k}}^{m_{k}})=O\left(N^{(m/s)-\bar{k}-\eta}(\operatorname{ln}N)^{Dm}\right),$
+where $\eta=(k-\bar{k})\epsilon|\operatorname{ln}\theta|.$ Since the number of such terms in (3.29) is $O\left(N^{\bar{k}}(\operatorname{ln}N)^{k-\bar{k}}\right),$ we conclude that the contribution of these non-separated terms is $O\left(N^{m/s-\eta}(\operatorname{ln}N)^{(D+1)m}\right),$ which is much smaller than the right-hand side of (3.28).
+∎
+The proof of (3.23) is similar to the proof of (3.14). Namely, as in §3.3, it suffices to establish the following analogue of (3.20):
 (3.31) $\mu\left(z\in I_{n}:|{\bar{{\bar{S}}}}_{N-n}(z)|>\frac{{\varepsilon}}{2}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq C\mu(I_{n}),$
-for some $C>0.$ Split ${\bar{{\bar{S}}}}_{N-n}={\bar{{\bar{S}}}}_{K\operatorname{ln}N}+\left({\bar{{\bar{S}}}}_{N-n}-{\bar{{\bar{S}}}}_{K\operatorname{ln}N}\right).$ The first term is estimated by (3.27).
-For the second term arguing in the same way as in the proof of (3.20) but replacing (3.18) by (3.28), we obtain
+for some constant $C>0.$ We split ${\bar{{\bar{S}}}}_{N-n}={\bar{{\bar{S}}}}_{K\operatorname{ln}N}+\left({\bar{{\bar{S}}}}_{N-n}-{\bar{{\bar{S}}}}_{K\operatorname{ln}N}\right).$ The first term is estimated using (3.27).
+For the second term, we proceed as in the proof of (3.20), but replace (3.18) with (3.28).
+This yields
 ${\mathbb{E}}\left(\left({\bar{{\bar{S}}}}_{N-n}-{\bar{{\bar{S}}}}_{K\operatorname{ln}N}\right)^{m}{\mathbf{1}}_{I_{n}}\right)\leq CN^{m/s}(\operatorname{ln}N)^{\delta m}\mu(I_{n}).$
-Therefore,
-(3.32) $\mu\left(z\in I_{n}:\left|{\bar{{\bar{S}}}}_{N-n}(z)-{\bar{{\bar{S}}}}_{K\operatorname{ln}N}(z)\right|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)$
-$\leq\frac{16C\mu(I_{n})}{{\varepsilon}^{2}}(\operatorname{ln}N)^{-m(\alpha-\delta)}.$
-Then we get (3.31) by combining (3.27) with (3.32).
+Therefore, we have
+(3.32) $\mu\left(z\in I_{n}:\left|{\bar{{\bar{S}}}}_{N-n}(z)-{\bar{{\bar{S}}}}_{K\operatorname{ln}N}(z)\right|>\frac{{\varepsilon}}{4}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\frac{16C\mu(I_{n})}{{\varepsilon}^{2}}(\operatorname{ln}N)^{-m(\alpha-\delta)}.$
+Finally, combining (3.27) with (3.32) yields (3.31).
 Appendix A Exponential mixing of all orders for expanding maps
 Proof of Lemma 2.4.
-We prove part(i) by induction. When $q=2$ the statement reduces to (2.4). For $q>2$ we apply (2.4) with $\psi_{1}=\varphi_{1},$ $\psi_{2}=\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}-k_{2}})$ and obtain
+We prove part (i) by induction.
+When $q=2$ the statement reduces to (2.7).
+For $q>2$ we apply (2.7) with $\psi_{1}=\varphi_{1},$ $\psi_{2}=\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}-k_{2}}),$ and obtain
 $\left\|\prod_{j=1}^{q}\left(\varphi_{j}\circ f^{k_{j}}\right)\right\|_{L^{1}}=\|\varphi_{1}\|_{L^{1}}\|\psi_{2}\|_{L^{1}}+O\left(\theta^{k_{2}-k_{1}}\|\varphi_{1}\|_{BV}\|\psi_{2}\|_{L^{1}}\right),$
-By induction
-$\|\psi_{2}\|_{L^{1}}\leq\left(\prod_{j=2}^{q-1}\left(||\varphi_{j}||_{L^{1}}+\theta^{k_{j}-k_{j-1}}\|\varphi_{j}\|_{BV}\right)\right)\|\varphi_{q}\|_{L^{1}}.$
+By the induction hypothesis, we have
+[…]
 It follows that
 $\left\|\prod_{j=1}^{q}\left(\varphi_{j}\circ f^{k_{j}}\right)\right\|_{L^{1}}\leq\left(\prod_{j=1}^{q-1}\left(\|\varphi_{j}\|_{L^{1}}+C\|\varphi_{j}\|_{BV}\theta^{k_{j+1}-k_{j}}\right)\right)\|\varphi_{q}\|_{L^{1}}.$
-Proceeding in a similar way, we can prove that
+Proceeding in a similar way, we can also prove that
 $\left\|\prod_{j=1}^{q}\left(\varphi_{j}\circ f^{k_{j}}\right)\right\|_{L^{1}}\geq\left(\prod_{j=1}^{q-1}\left(\|\varphi_{j}\|_{L^{1}}-C\|\varphi_{j}\|_{BV}\theta^{k_{j+1}-k_{j}}\right)\right)\|\varphi_{q}\|_{L^{1}}.$
-Now we turn to the proof of part (ii). When $q=2$ the statement also reduces to (2.4). Applying (2.4) with $\psi_{1}=\varphi_{1},$ $\psi_{2}=\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}-k_{2}}),$ we get
+Now we turn to the proof of part (ii).
+When $q=2,$ the statement reduces to (2.7).
+Applying (2.7) with $\psi_{1}=\varphi_{1}$ and $\psi_{2}=\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}-k_{2}}),$ we obtain
 $\left|{\mathbb{E}}\left(\prod_{j=1}^{q}(\varphi_{j}\circ f^{k_{j}})\right)-\left({\mathbb{E}}\varphi_{1}\right){\mathbb{E}}\left(\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}})\right)\right|$
 $\leq$ $C\theta^{k}\|\varphi_{1}\|_{BV}\left\|\left(\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}})\right)\right\|_{L_{1}}\leq C\theta^{k}\|\varphi_{1}\|_{BV}\left\|\left(\prod_{j=2}^{q}(|\varphi_{j}|\circ f^{k_{j}})\right)\right\|_{L_{1}}\leq C\theta^{k}\prod_{j=1}^{q}\|\varphi_{j}\|_{BV},$
-where the last inequality is due to part (i) of the lemma, which we have already proved.
-Applying inductive estimate to ${\mathbb{E}}\left(\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}})\right)$ we obtain (2.6).
-$\square$
+where the last inequality follows from part (i) of the lemma, which we have already proven.
+Applying the inductive estimate to ${\mathbb{E}}\left(\prod_{j=2}^{q}(\varphi_{j}\circ f^{k_{j}})\right)$ we obtain (2.9).
+∎
 Appendix B Good targets for expanding maps
 Proof of Proposition 2.3.
-We prove for $\Omega_{\rho_{N},j},$ and the proof for $\Omega_{\rho_{N}}$ follows in the same approach.
-Denote $B(x,r_{1},r_{2})=\{y\in\mathbb{S}:r_{1}\leq|y-x|\leq r_{2}\},$ and $\tilde{A}^{+}_{\rho_{N},j},$ $\tilde{A}^{-}_{\rho_{N},j}$ to be the characteristic functions of
+We use the notation $\Omega_{\rho_{N}}$ and $\Omega_{\rho_{N},j}$ as defined in (2.4) and (2.5), respectively.
+We define the annular set $B(x,r_{1},r_{2})=\{y\in\mathbb{S}:r_{1}\leq|y-x|\leq r_{2}\},$ and let $\tilde{A}^{+}_{\rho_{N},j}$, $\tilde{A}^{-}_{\rho_{N},j}$ be the characteristic functions of the sets
 $B\left(x,\left(\frac{a}{\rho_{N}(c_{j}+\varepsilon)+L}\right)^{s},\left(\frac{a}{\rho_{N}(c_{j}-\varepsilon)-L}\right)^{s}\right)$
@@
 $B\left(x,\left(\frac{a}{\rho_{N}(c_{j}+\varepsilon)-L}\right)^{s},\left(\frac{a}{\rho_{N}(c_{j}-\varepsilon)+L}\right)^{s}\right),$
-respectively. Then
-$\tilde{A}^{-}_{\rho_{N},j}\leq{\mathbf{1}}_{\Omega_{\rho_{N},j}}\leq\tilde{A}^{+}_{\rho_{N},j},\quad E\tilde{A}^{+}_{\rho_{N},j}-E\tilde{A}^{-}_{\rho_{N},j}\leq C\rho_{N}^{-(s+1)}\text{ and }\|\tilde{A}^{*}_{\rho_{N},j}\|_{BV}\leq 2\mu(\Omega_{\rho_{N},j})+4,$
-[…]
-First we prove $(M1)_{r}.$ If $0<k_{1}<k_{2}<\dots<k_{r}\leq N$ are such that $k_{i+1}-k_{i}\geq K\operatorname{ln}N,$ where $K$ is a sufficiently large constant, and each $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r},$
-$\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right)\leq{\mathbb{E}}\left(\prod_{j=1}^{r}\tilde{A}^{+}_{\rho_{N},i_{j}}\circ f^{k_{j}}\right)\leq$
-$\left(\prod_{j=1}^{r-1}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{j}}+C\theta^{K\operatorname{ln}N}(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{j}}+4)\right)\right)({\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{r}})\leq\left(\prod_{j=1}^{r}\sigma_{\rho_{N},i_{j}}\right)(1+{\varepsilon}_{N})$
-for ${\varepsilon}_{N}\rightarrow 0,$ which yields the RHS of $(M1)_{r}.$ The LHS is proved similarly replacing $\tilde{A}^{+}_{\rho_{N},i_{j}}$ by $\tilde{A}^{-}_{\rho_{N},i_{j}}$.
-Second we show $(M2)_{r}$ assuming that $x$ is Diophanitine.
-Let $\operatorname{Sep}(k_{1},\dots k_{r})=m<r.$ Since $x$ is Diophantine, there exists $\epsilon>0$ such that $B(x,c\rho_{N}^{-s})\cap f^{-k}B(x,c\rho_{N}^{-s})=\emptyset$ for $k\leq\epsilon\operatorname{ln}N$ and $c=2a^{s}/c_{j}^{s}.$ Thus we assume that $k_{j+1}-k_{j}\geq\epsilon\operatorname{ln}N$ for all $j.$ Then
-$\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right)$$\leq$ $C\left(\prod_{k_{j+1}-k_{j}>K\operatorname{ln}N}\sigma_{\rho_{N},i_{j}}\right)(\theta^{\epsilon\operatorname{ln}N})^{r-m}(1+{\varepsilon}_{N})\leq{\frac{C\sigma_{{\rho}_{N}}^{m}}{(\operatorname{ln}N)^{100r}}}.$
-Third we prove $(M3)_{r}.$ If $k_{1}<\dots<k_{r}<l_{1}<\dots<l_{r}$ are such that $2^{i}<k_{\alpha}\leq 2^{i+1},2^{j}<l_{\beta}\leq 2^{j+1}$ for $1\leq\alpha,\beta\leq r$, then for each $(i_{1},i_{2},\dots,i_{r}),\,(j_{1},j_{2}\dots j_{r})\in\{1,\dots,p\}^{r},$
-$\mu\left(\left[\bigcap_{\alpha=1}^{r}\Omega^{k_{\alpha}}_{{\rho}_{2^{i}},i_{\alpha}}\right]\bigcap\left[\bigcap_{\beta=1}^{r}\Omega^{l_{\beta}}_{{\rho}_{2^{j}},j_{\beta}}\right]\right)\leq{\mathbb{E}}\left(\left(\prod_{\alpha=1}^{r}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}\circ f^{k_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}\circ f^{k_{\beta}}\right)\right)$
-$\leq\left(\prod_{\alpha=1}^{r}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}+C\theta^{Ki}(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}+4)\right)\right)\left(\prod_{\beta=1}^{r-1}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}+C\theta^{Kj}(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}+4)\right)\right)({\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{r}})$
-$\leq\left(\prod_{\alpha=1}^{r}\sigma_{\rho_{2^{i}},i_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\sigma_{\rho_{2^{j}},j_{\beta}}\right)(1+{\varepsilon}_{i}).$
-Finally we prove $(M4)_{*}$ assuming that $x$ is non-periodic. The proof of $(M4)_{I_{1},\dots I_{r},i_{1},\dots,i_{r}}$ is similar and will be left to the reader. We denote by
-$\tau_{*}^{N}=\sum_{\overset{0<k_{1}<k_{2}<\dots<k_{r}\leq N}{{\operatorname{Sep}_{n}}(k_{1},\dots k_{r})=r}}\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N}}\right).$
-Using $(M1)_{r},$ which has been proved, we have $\lim_{N\rightarrow\infty}\tau_{*}^{N}/\left(N\sigma_{\rho_{N}}\right)^{r}=1.$ Therefore, it suffices to show that
-$\lim_{N\rightarrow\infty}\frac{\mu(A_{*}^{N})}{\tau_{*}^{N}}=1.$
-By Boole’s inequality $\mu(A_{*}^{N})\leq\tau_{*}^{N}.$ For the lower bound we note that
-(B.1) $\mu\left(A_{*}^{N}\right)\geq\tau_{*,b}^{N}-\sum^{*}\mu(\Omega^{k_{1}}_{{\rho}_{N}}\cap\dots\cap\Omega^{k_{r}}_{{\rho}_{N}}\cap\Omega^{k_{r+1}}_{{\rho}_{N}}),$
-where the sum in $\sum^{*}$ is taken over all tuples satisfying $\operatorname{Sep}(k_{1},\dots k_{r})=r.$ We now consider several cases depending on the size of $p=\min_{j=1,\dots,r}|k_{r+1}-k_{j}|$.
-(i) Since $x$ is non-periodic for each fixed $p_{0},$ we have that $\Omega_{{\rho}_{N}}\cap f^{-p}\Omega_{{\rho}_{N}}=\emptyset$ for all $p\leq p_{0}$ provided that $n$ is large enough.
-Hence the contribution of terms $p\leq p_{0}$ is zero.
-(ii) Next we suppose that $p_{0}<p\leq\varepsilon\operatorname{ln}N.$ Precisely we may assume that $k_{r+1}=k_{\bar{j}}+p$ for some ${\bar{j}}\leq r$ (the case where $k_{r+1}=k_{\bar{j}}-p$ is similar).
-Define $\bar{A}_{{\rho}_{N}}^{+}$ to be the characteristic function of $B\left(x,\left(\frac{a}{\rho_{N}-L}\right)^{s}\right).$ Then
-$\bar{A}^{+}_{\rho_{N}}\geq{\mathbf{1}}_{\Omega_{\rho_{N}}},\,{\mathbb{E}}\bar{A}^{+}_{\rho_{N}}-\mu(\Omega_{\rho_{N}})\leq C\rho_{N}^{-(s+1)}\text{ and }\|\bar{A}^{+}_{\rho_{N}}\|_{BV}\leq 2\mu(\Omega_{\rho_{N}})+4.$
+respectively, where $L=\max_{x\in\mathbb{S}}\left|\psi(x)\right|.$ Then
+$\tilde{A}^{-}_{\rho_{N},j}\leq{\mathbf{1}}_{\Omega_{\rho_{N},j}}\leq\tilde{A}^{+}_{\rho_{N},j},\quad{\mathbb{E}}\tilde{A}^{+}_{\rho_{N},j}-{\mathbb{E}}\tilde{A}^{-}_{\rho_{N},j}\leq C\rho_{N}^{-(s+1)}\text{ and }\left\|\tilde{A}^{\pm}_{\rho_{N},j}\right\|_{BV}\leq 2\mu(\Omega_{\rho_{N},j})+4.$
+We now verify the conditions $(M1)_{r},$ $(M2)_{r},$ $(M3)_{r},$ $(M4)_{*}$ and $(M4)_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}$ using Lemma 2.4.
+(a) We prove $(M1)_{r}$ and $(M3)_{r}$ for $\Omega_{\rho_{N},j},$ and the proof for $\Omega_{\rho_{N}}$ follows in the same approach.
+Suppose $0<k_{1}<k_{2}<\dots<k_{r}\leq N$ are such that $k_{i+1}-k_{i}\geq K\operatorname{ln}N,$ where $K$ is a sufficiently large constant, and each $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r}.$ Then we have
+$\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right)\leq{\mathbb{E}}\left(\prod_{j=1}^{r}\tilde{A}^{+}_{\rho_{N},i_{j}}\circ f^{k_{j}}\right)$
+$\leq$ $\left(\prod_{j=1}^{r-1}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{j}}+C\theta^{K\operatorname{ln}N}\left(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{j}}+4\right)\right)\right){\mathbb{E}}\tilde{A}^{+}_{\rho_{N},i_{r}}\leq\left(\prod_{j=1}^{r}\sigma_{\rho_{N},i_{j}}\right)(1+{\varepsilon}_{N})$
+where ${\varepsilon}_{N}\to 0$ as $N\to\infty.$ This establishes the right-hand side of $(M1)_{r}.$ The proof for the left-hand side follows similarly, replacing $\tilde{A}^{+}_{\rho_{N},i_{j}}$ by $\tilde{A}^{-}_{\rho_{N},i_{j}}$.
+Next we prove $(M3)_{r}.$ Suppose $k_{1}<\dots<k_{r}<l_{1}<\dots<l_{r}$ are such that $2^{i}<k_{\alpha}\leq 2^{i+1}$ and $2^{j}<l_{\beta}\leq 2^{j+1}$ for $1\leq\alpha,\beta\leq r.$ For each pairs $(i_{1},i_{2},\dots,i_{r})$ and $(j_{1},j_{2}\dots j_{r})$ in $\{1,\dots,p\}^{r},$ we have
+$\mu\left(\left(\bigcap_{\alpha=1}^{r}\Omega^{k_{\alpha}}_{{\rho}_{2^{i}},i_{\alpha}}\right)\bigcap\left(\bigcap_{\beta=1}^{r}\Omega^{l_{\beta}}_{{\rho}_{2^{j}},j_{\beta}}\right)\right)\leq{\mathbb{E}}\left(\left(\prod_{\alpha=1}^{r}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}\circ f^{k_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}\circ f^{k_{\beta}}\right)\right)$
+$\leq$ $\left(\prod_{\alpha=1}^{r}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}+C\theta^{Ki}\left(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{i}},i_{\alpha}}+4\right)\right)\right)\left(\prod_{\beta=1}^{r-1}\left({\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}+C\theta^{Kj}\left(2{\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{\beta}}+4\right)\right)\right){\mathbb{E}}\tilde{A}^{+}_{\rho_{2^{j}},j_{r}}$
+$\leq$ $\left(\prod_{\alpha=1}^{r}\sigma_{\rho_{2^{i}},i_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\sigma_{\rho_{2^{j}},j_{\beta}}\right)(1+{\varepsilon}_{i}),$
+which proves $(M3)_{r}.$
+(b) We prove $(M2)_{r}$ under the assumption that $x$ is Diophantine. Let $\operatorname{Sep}_{N}(k_{1},\dots k_{r})=m<r.$ Since $x$ is Diophantine, there exists $\epsilon>0$ such that
+$B\left(x,c\rho_{N}^{-s}\right)\cap f^{-k}B\left(x,c\rho_{N}^{-s}\right)=\emptyset$
+for positive integer $k\leq\epsilon\operatorname{ln}N,$ where $c=2a^{s}/c_{j}^{s}.$ Therefore, we assume that $k_{j+1}-k_{j}\geq\epsilon\operatorname{ln}N$ for all $j.$ Then we obtain
+$\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N}}\right)$ $\leq$ $C\left(\prod_{k_{j+1}-k_{j}>K\operatorname{ln}N}\sigma_{\rho_{N}}\right)(\theta^{\epsilon\operatorname{ln}N})^{r-m}\leq{\frac{C\sigma_{{\rho}_{N}}^{m}}{(\operatorname{ln}N)^{100r}}}.$
+(c) Recall that
+$A_{*}^{N}$ $=$ $\left\{\omega\in\Omega:\,\exists\,0<k_{1}<k_{2}<\dots<k_{r}\leq N\,\,\text{such that}\,\,\operatorname{Sep}_{N}(k_{1},\dots,k_{r})=r\,\,\text{and}\,\,\omega\in\Omega^{k_{j}}_{{\rho}_{N}}\right\}$
+$=$ $\bigcup_{\overset{0<k_{1}<\dots<k_{r}\leq N}{{\operatorname{Sep}_{N}}(k_{1},\dots,k_{r})=r}}\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N}}\right).$
+We define
+$\tau_{*}^{N}=\sum_{\overset{0<k_{1}<\dots<k_{r}\leq N}{{\operatorname{Sep}_{N}}(k_{1},\dots,k_{r})=r}}\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N}}\right).$
+Using $(M1)_{r}$ which has been established, we obtain
+$\lim_{N\rightarrow\infty}\frac{\tau_{*}^{N}}{\left(N\sigma_{\rho_{N}}\right)^{r}}=1.$
+In addition, since $\mu(A_{*}^{N})\leq\tau_{*}^{N},$ we have
+$\limsup_{N\rightarrow\infty}\frac{\mu(A_{*}^{N})}{\left(N\sigma_{\rho_{N}}\right)^{r}}\leq 1.$
+(d) We first assume that the point $x$ is not periodic and define
+(B.1) $\bar{\tau}_{*}^{N}=\sum_{k_{j}/N\in I_{j}}\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right).$
+Recall that
+$A_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}$ $=$ $\left\{\omega\in\Omega:\,\exists\,k_{j}\,\,\text{such that}\,\,k_{j}/N\in I_{j}\,\,\text{and }\,\,\omega\in\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right\}$
+$=$ $\bigcup_{k_{j}/N\in I_{j}}\left(\bigcap_{j=1}^{N}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right).$
+Using $(M1)_{r}$ which has been proved, we obtain
+$\lim_{N\rightarrow\infty}\frac{\bar{\tau}_{*}^{N}}{N^{r}\Pi_{j=1}^{r}\left(\sigma_{\rho_{N},i_{j}}|I_{j}|\right)}=1.$
+Therefore, it suffices to show that
+$\lim_{N\rightarrow\infty}\frac{\mu\left(A_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}\right)}{\bar{\tau}_{*}^{N}}=1.$
+By Boole’s inequality, we have
+$\mu\left(A_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}\right)\leq\bar{\tau}_{*}^{N}$
+and
+(B.2) $\mu\left(A_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}\right)\geq\bar{\tau}_{*}^{N}-\sum_{t=r+1}^{2r}\bar{\tau}_{t}^{N},$
+where
+$\bar{\tau}_{t}^{N}=\sum_{\overset{0<k_{1}<\dots<k_{r}\leq N}{{\operatorname{Sep}_{N}}(k_{1},\dots,k_{r})=r}}\mu\left(\Omega^{k_{1}}_{{\rho}_{N}}\cap\dots\cap\Omega^{k_{r}}_{{\rho}_{N}}\cap\Omega^{k_{r+1}}_{{\rho}_{N}}\cap\cdots\cap\Omega^{k_{t}}_{{\rho}_{N}}\right),$
+because $\Omega_{\rho_{N},i_{j}}\subset\Omega_{\rho_{N}}=\{y:\phi(y)>C\rho_{N}\}$ for all $i_{j},$ $1\leq j\leq r,$ and some $C>0.$
+We prove that
+$\frac{\bar{\tau}_{r+1}^{N}}{\bar{\tau}_{*}^{N}}\to 0\,\,\,\text{as}\,\,\,N\to\infty.$
+The proof of $\bar{\tau}_{t}^{N}/\bar{\tau}_{*}^{N}\to 0$ for other values of $t$ follows similarly. We consider several cases depending on the size of $p=\min_{j=1,\dots,r}|k_{r+1}-k_{j}|$.
+(i) Since $x$ is non-periodic, for each fixed $p_{0},$ we have $\Omega_{{\rho}_{N}}\mathrel{{$\bigcap$}}f^{-p}\Omega_{{\rho}_{N}}=\emptyset$ for all $1\leq p<p_{0},$ provided that $N$ is large enough.
+Therefore, the contribution from terms with $p<p_{0}$ is zero.
+(ii) Next, suppose that $p_{0}\leq p\leq\varepsilon\operatorname{ln}N.$ Specifically, we may assume that $k_{r+1}=k_{\bar{j}}+p$ for some ${\bar{j}}\leq r$ (the case where $k_{r+1}=k_{\bar{j}}-p$ is analogous).
+Define $\bar{A}_{{\rho}_{N}}^{+}$ as the characteristic function of the set
+$B\left(x,\left(\frac{a}{C\rho_{N}-L}\right)^{s}\right).$
+Then we have
+$\bar{A}^{+}_{\rho_{N}}\geq{\mathbf{1}}_{\Omega_{\rho_{N}}},\,{\mathbb{E}}\bar{A}^{+}_{\rho_{N}}-\mu(\Omega_{\rho_{N}})\leq C\rho_{N}^{-(s+1)}\text{ and }\left\|\bar{A}^{+}_{\rho_{N}}\right\|_{BV}\leq 2\mu(\Omega_{\rho_{N}})+4.$
 For $j=1,\dots,r,$ define
-${\hat{A}}_{j}=\begin{cases}\bar{A}_{{\rho}_{N}}^{+}&\text{if }j\neq{\bar{j}},\\ \bar{A}_{{\rho}_{N}}^{+}(\bar{A}_{{\rho}_{N}}^{+}\circ f^{p})&\text{if }j={\bar{j}}.\end{cases}$
-If we denote by $M=\max_{x\in\mathbb{S}}|f^{\prime}(x)|,$ then
+${\hat{A}}_{j}=\begin{cases}\bar{A}_{{\rho}_{N}}^{+},&\text{if }j\neq{\bar{j}};\\ \bar{A}_{{\rho}_{N}}^{+}\left(\bar{A}_{{\rho}_{N}}^{+}\circ f^{p}\right),&\text{if }j={\bar{j}}.\end{cases}$
+Let $\Lambda=\max_{x\in\mathbb{S}}|f^{\prime}(x)|.$ Then, we have
 $\mu\left(\prod_{j=1}^{r+1}\Omega_{{\rho}_{N}}^{k_{j}}\right)\leq{\mathbb{E}}\left(\prod_{j=1}^{r}{\hat{A}}_{j}\circ f^{k_{j}}\right)$
-$\leq$ $\left(\prod_{j\neq{\bar{j}}}\left({\mathbb{E}}{\hat{A}}_{j}+C\|\bar{A}_{{\rho}_{N}}^{+}\|_{BV}\theta^{K\operatorname{ln}N}\right)\right)\left({\mathbb{E}}{\hat{A}}_{\bar{j}}+C\|\bar{A}_{{\rho}_{N}}^{+}\|_{BV}^{2}M^{p}\theta^{K\operatorname{ln}N}\right)\leq\prod_{j=1}^{r}\left({\mathbb{E}}{\hat{A}}_{j}+n^{-10}\right)$
-if $s(n)=K\operatorname{ln}N$ and $K$ is large enough. Noting that
-${\mathbb{E}}({\hat{A}}_{\bar{j}})={\mathbb{E}}\left(\bar{A}_{{\rho}_{N}}^{+}(\bar{A}_{{\rho}_{N}}^{+}\circ f^{p})\right)\leq C\sigma_{{\rho}_{N}}\theta^{p}$
-by (2.4), we have that each term with $p$ as in the case (ii) is at most $C\sigma_{{\rho}_{N}}^{r}\theta^{p}.$ Summing over all $r$ tuples and over all $p\in(p_{0},\varepsilon\operatorname{ln}N]$ yields a total contribution of $CN^{r}\sigma_{\rho_{N}}^{r}\theta^{p_{0}}$ for some constant $C$ independent of $p_{0}.$
-(iii) The terms where ${\varepsilon}\operatorname{ln}N<p\leq K\operatorname{ln}N$ are $O(\sigma_{\rho_{N}}^{r+\eta})$ for some $\eta=\eta({\varepsilon})>0.$ Thus the total contribution of those terms are $O(N^{r}(\operatorname{ln}N)\;\sigma_{\rho_{N}}^{r+\eta}),$ which is also negligible.
-(iv) The terms where $p>K\operatorname{ln}N$ for sufficiently large $K$ is $O(\sigma_{\rho_{N}}^{r+1})$ similarly to the proof of $(M1)_{r}.$ Thus the total contribution of those terms are $O((N\sigma_{\rho_{N}})^{r+1})$ which is negligible compared to $\tau_{*}^{N}.$
-Summing the estimates of cases (i)–(iii) above we see that the second term in the RHS of (B.1) is negligible comparing to the first, which proves $(M4)_{*}.$square
+$\leq$ $\left(\prod_{j\neq{\bar{j}}}\left({\mathbb{E}}{\hat{A}}_{j}+C\left\|\bar{A}_{{\rho}_{N}}^{+}\right\|_{BV}\theta^{K\operatorname{ln}N}\right)\right)\left({\mathbb{E}}{\hat{A}}_{\bar{j}}+C\left\|\bar{A}_{{\rho}_{N}}^{+}\right\|_{BV}^{2}\Lambda^{p}\theta^{K\operatorname{ln}N}\right)\leq\prod_{j=1}^{r}\left({\mathbb{E}}{\hat{A}}_{j}+N^{-10}\right)$
+if $s(n)=K\operatorname{ln}N$ and $K$ is large enough.
+Noting that ${\mathbb{E}}{\hat{A}}_{\bar{j}}={\mathbb{E}}\left(\bar{A}_{{\rho}_{N}}^{+}(\bar{A}_{{\rho}_{N}}^{+}\circ f^{p})\right)\leq C\sigma_{{\rho}_{N}}\theta^{p}$ by (2.7), we conclude that each term for $p$ as in the case (ii) is at most $C\sigma_{{\rho}_{N}}^{r}\theta^{p}.$ Summing over all $r$ tuples and over all $p\in[p_{0},\varepsilon\operatorname{ln}N],$ we obtain a total contribution of $CN^{r}\sigma_{\rho_{N}}^{r}\theta^{p_{0}}$ for some constant $C$ independent of $p_{0}.$
+(iii) The terms where ${\varepsilon}\operatorname{ln}N<p\leq K\operatorname{ln}N$ contribute $O(\sigma_{\rho_{N}}^{r+\eta})$ for some $\eta=\eta({\varepsilon})>0.$ Therefore, the total contribution of those terms is $O\left(N^{r}(\operatorname{ln}N)\;\sigma_{\rho_{N}}^{r+\eta}\right).$
+(iv) The terms where $p>K\operatorname{ln}N$ for sufficiently large $K$ contribute $O(\sigma_{\rho_{N}}^{r+1}),$ similar to the proof of $(M1)_{r}.$ Hence, the total contribution of these terms is $O\left((N\sigma_{\rho_{N}})^{r+1}\right).$
+Summing the estimates from cases (i)–(iv), we observe that $\bar{\tau}_{r+1}^{N}$ is negligible in comparison to $\tau_{*}^{N}.$ This completes the proof of $(M4)_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}$ for non-periodic points.
+Next, we consider the case where the point $x$ is periodic. Suppose the period of $x$ is $q,$ i.e., $f^{q}(x)=x.$ We use the notation $\tilde{\Omega}_{\rho_{N},j}$ as defined in (2.6) and define
+$\tilde{A}_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}=\left\{\omega\in\Omega:\,\exists\,k_{j}\,\,\text{such that}\,\,k_{j}/N\in I_{j}\,\,\text{and }\,\,\omega\in\tilde{\Omega}^{k_{j}}_{{\rho}_{N},i_{j}}\right\}.$
+We now proceed to prove $(M4)_{I_{1},\dots I_{r};i_{1},\dots,i_{r}}$ for $\tilde{A}_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}.$
+Observe that
+$f^{-p_{0}}\Omega_{{\rho}_{N},j}=\left(f^{-(p_{0}-q)}\Omega_{{\rho}_{N},j}\mathrel{{\bigcap}}f^{-p_{0}}\Omega_{{\rho}_{N},j}\right)\mathrel{{\bigcup}}\tilde{\Omega}_{\rho_{N},j}.$
+Indeed if $f^{k}y\in\Omega_{{\rho}_{N},j},$ $f^{n}y\in\Omega_{{\rho}_{N},j}$ can only occur if $n=k+iq$ for $i\in\mathbb{Z}.$ Therefore, $(p_{0}-k)$ must be a multiple of $q$.
+In this case, $f^{p_{0}}y\in\Omega_{{\rho}_{N},j}$ implies $f^{p_{0}-q}y\in\Omega_{{\rho}_{N},j}$, since $f$ is expanding.
+In addition, we have
+(B.3) $f^{-(p_{0}-q)}\Omega_{{\rho}_{N},j}\mathrel{{$\bigcap$}}f^{-p_{0}}\Omega_{{\rho}_{N},j}=\emptyset$
+whenever $0<{\varepsilon}<\left(\gamma^{q/s}-1\right)c_{j}/\left(\gamma^{q/s}+1\right),$ where $\varepsilon$} is from the definition of
+$\Omega_{\rho_{N},j}=\left\{y:\frac{\phi(y)}{\rho_{N}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}]\right\},$
+and $\gamma=\min_{x\in\mathbb{S}}|f^{\prime}(x)|>1.$ If $f^{p_{0}-q}(y)\in\Omega_{{\rho}_{N},j},$ we have
+$\left|f^{p_{0}}(y)-x\right|\geq\gamma^{q}\left|f^{p_{0}-q}(y)-x\right|\geq\gamma^{q}\left(\frac{a}{\rho_{N}(c_{j}+\varepsilon)+L}\right)^{s}>\left(\frac{a}{\rho_{N}(c_{j}-\varepsilon)-L}\right)^{s}$
+where $L=\max_{x\in\mathbb{S}}\left|\psi(x)\right|.$ Therefore, we conclude that $f^{p_{0}}(y)\notin\Omega_{{\rho}_{N},j},$ and (B.3) follow.
+Hence $\tilde{\Omega}_{\rho_{N},j}=f^{-p_{0}}\Omega_{{\rho}_{N},j}$ for $0<{\varepsilon}<\frac{(\gamma^{q/s}-1)c_{j}}{\gamma^{q/s}+1}$ and $\mu\left(\tilde{\Omega}_{\rho_{N},j}\right)=\mu\left(f^{-p_{0}}\Omega_{{\rho}_{N},j}\right)=\sigma_{\rho_{N},j}.$
+We will show that
+$\lim_{N\rightarrow\infty}\frac{\mu\left(\tilde{A}_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}\right)}{N^{r}\Pi_{j=1}^{r}\left(\sigma_{\rho_{N},i_{j}}|I_{j}|\right)}=1.$
+Note that
+$\sum_{k_{j}/N\in I_{j}}\mu\left(\bigcap_{j=1}^{r}\tilde{\Omega}^{k_{j}}_{{\rho}_{N},i_{j}}\right)=\sum_{k_{j}/N\in I_{j}}\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right)=\bar{\tau}_{*}^{N},$
+which coincides with the non-periodic case. We prove $(M4)_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}$ using the inequality:
+(B.4) $\mu\left(\tilde{A}_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}\right)\geq\bar{\tau}_{*}^{N}-\sum_{t=r+1}^{2r}\tilde{\tau}_{t}^{N},$
+where
+$\tilde{\tau}_{t}^{N}=\sum_{k_{j}/N\in I_{j}}\mu\left(\tilde{\Omega}^{k_{1}}_{{\rho}_{N},i_{1}}\cap\dots\cap\tilde{\Omega}^{k_{r}}_{{\rho}_{N},i_{r}}\cap\tilde{\Omega}^{k_{r+1}}_{{\rho}_{N},i_{r+1}}\cap\cdots\cap\tilde{\Omega}^{k_{t}}_{{\rho}_{N},i_{t}}\right).$
+We now show that $\frac{\tilde{\tau}_{r+1}^{N}}{\bar{\tau}_{*}^{N}}\to 0\,\,\,\text{as}\,\,\,N\to\infty.$ The proof for other values of $t$ follows similarly.
+We consider several cases depending on the value of $p=\min_{j=1,\dots,r}|k_{r+1}-k_{j}|$.
+(1) If $p<p_{0}$ where $p_{0}$ is specified in the definition of $\tilde{\Omega}_{{\rho}_{N},j},$ we take $j_{0}$ such that $k_{r+1}=k_{j_{0}}+p$ (the case where $k_{r+1}=k_{j_{0}}-p$ is similar).
+Since $d(I_{i},I_{j})>{\varepsilon}$ for $i\neq j,$ both $k_{r+1}/N$ and $k_{j_{0}}/N$ belong to the same component $I_{j},$ thus $i_{r+1}=i_{j_{0}}.$ We then obtain
+$\tilde{\Omega}^{k_{j_{0}}}_{{\rho}_{N},i_{j_{0}}}\cap\tilde{\Omega}^{k_{r+1}}_{{\rho}_{N},i_{r+1}}\subset f^{-k_{j_{0}}}\left(f^{-p_{0}}\Omega_{{\rho}_{N},i_{j_{0}}}\mathrel{{\bigcap}}\left(f^{-p}\Omega_{{\rho}_{N},i_{j_{0}}}^{c}\cap\cdots\cap f^{-p-p_{0}+1}\Omega_{{\rho}_{N},i_{j_{0}}}^{c}\right)\right)=\emptyset.$
+(2) If $p\geq p_{0},$ we take $C>0$ such that
+$\Omega_{\rho_{N},i_{j}}\subset\Omega_{\rho_{N}}=\{y:\phi(y)>C\rho_{N}\}$
+for all $i_{j},$ $1\leq j\leq r.$ It follows that
+$\tilde{\Omega}_{{\rho}_{N},i_{j}}\subset f^{-p_{0}}\Omega_{{\rho}_{N},i_{j}}\subset f^{-p_{0}}\Omega_{\rho_{N}}$
+and hence
+$\mu\left(\mathrel{{\bigcap}}_{j=1}^{r+1}\tilde{\Omega}^{k_{j}}_{{\rho}_{N},i_{j}}\right)\leq\mu\left(\mathrel{{\bigcap}}_{j=1}^{r+1}f^{-p_{0}}\Omega^{k_{j}}_{{\rho}_{N}}\right)=\mu\left(\mathrel{{\bigcap}}_{j=1}^{r+1}\Omega^{k_{j}}_{{\rho}_{N}}\right).$
+Thus, by the same arguments as (ii)-(iv) in the non-periodic case, we conclude that $\tilde{\tau}_{r+1}^{N}$ is negligible compared to $\bar{\tau}_{*}^{N}.$ This completes the proof of $(M4)_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}$ for periodic points.
+∎
 Proof of Lemma 2.6.
@@
 $\tilde{H}_{k,{\bar{\varepsilon}}}=\left\{y:\exists\,{\bar{n}},\,{\hat{n}}\leq 2^{k+1},\,\,k^{(\alpha-{\bar{\alpha}})/2}\leq|{\bar{n}}-{\hat{n}}|\leq 2s(2^{k+1}),\,\,\min\{X_{\bar{n}},X_{\hat{n}}\}>{\bar{\varepsilon}}2^{k/s}k^{\alpha-1}\right\}.$
-Since $H_{N,{\bar{\varepsilon}}}\subset\tilde{H}_{k,{\bar{\varepsilon}}}$ for $2^{k}\leq N\leq 2^{k+1},$ it suffices to show that $\tilde{H}_{k,{\bar{\varepsilon}}}$ happens only for finitely many $k.$ Using (2.4) and similar method in the proof of Proposition 2.3 we conclude that
+Since $H_{N,{\bar{\varepsilon}}}\subset\tilde{H}_{k,{\bar{\varepsilon}}}$ for $2^{k}\leq N\leq 2^{k+1},$ it suffices to show that $\tilde{H}_{k,{\bar{\varepsilon}}}$ occurs only for finitely many $k.$ Using (2.7) and similar method to the proof of Proposition 2.3 we conclude that
 $\mu(\tilde{H}_{k,{\bar{\varepsilon}}})\leq\frac{C2^{k+1}(k+1)}{2^{k}k^{s(\alpha-1)}}\theta^{k(\alpha-{\bar{\alpha}})/2}=2Ck^{s(1-\alpha)+1}2^{k(\alpha-{\bar{\alpha}})\operatorname{ln}\theta/2}.$
-Hence $\sum_{k}\mu(\tilde{H}_{k,{\bar{\varepsilon}}})<\infty$ and the result follows from the classical Borel-Cantelli Lemma. $\square$
+Hence $\sum_{k}\mu(\tilde{H}_{k,{\bar{\varepsilon}}})<\infty$ and the result follows from the classical Borel-Cantelli lemma. ∎
 Proof of Lemma 3.1.
-Let $p=\min_{1\leq j\leq m-1}\{k_{j+1}-k_{j}\}.$ Since $x$ is non-periodic, for each fixed $p_{0},$ we have that ${\hat{B}}_{N,j}\cap f^{-p}{\hat{B}}_{N,j}=\emptyset$ for all $p\leq p_{0}$ provided that $N$ is large enough because $x$ is non-periodic.
-Thus the contribution of the terms with tuples satisfying $p\leq p_{0}$ is zero for large $N.$
-We suppose that $\operatorname{Sep}(k_{1},\dots k_{m})=l<m.$ If $p_{0}<p\leq\varepsilon\operatorname{ln}N,$ Arguing as the case $(M4)_{*}$(ii) in the proof of Proposition 2.3, we conclude that each term is smaller than $C\mu({\hat{B}}_{N,j})^{l}\theta^{p}$ for fixed $p.$ Since the total number of terms is $O\left(N^{l}p^{m-l}\right),$ we have that total contribution of the terms with $p={\hat{p}}$ is $O\left(N^{l}{\hat{p}}^{m-l}\mu({\hat{B}}_{N,j})^{l}\theta^{\hat{p}}\right).$ The sum over tuples with ${\hat{p}}\in[p_{0},{\varepsilon}\operatorname{ln}N]$ gives a total contribution of $CN^{l}{p_{0}}^{m-l}\mu({\hat{B}}_{N,j})^{l}\theta^{p_{0}}$ for fixed $l$ and $p_{0}$ large enough (independent of $N$).
-On the other hand if $\varepsilon\operatorname{ln}N<p\leq K\operatorname{ln}N,$ then similarly to the proof of the case $(M4)_{*}$(iii) we obtain that
-$\mu\left(\bigcap_{j=1}^{m}\left\{f^{n_{j}}y\in{\hat{B}}_{N,j}\right\}\right)\leq C\mu({\hat{B}}_{N,j})^{l+\eta}.$
-Summing over all $r$ tuples and over all $p\in(\varepsilon\operatorname{ln}N,K\operatorname{ln}N]$ yields a total contribution that is not greater than $C\left(N\mu({\hat{B}}_{N,j})\right)^{l}\mu({\hat{B}}_{N,j})^{\eta}\left(\operatorname{ln}N\right)^{m-l}.$
-Finally, if $\operatorname{Sep}(k_{1},\dots k_{m})=m,$ similarly to the case $(M1)_{r}$ in the proof of Proposition 2.3, we obtain that
-$\mu\left(\bigcap_{j=1}^{m}\left\{f^{k_{j}}y\in{\hat{B}}_{N,j}\right\}\right)\leq C\mu({\hat{B}}_{N,j})^{m}.$
-Summing over all tuples $(k_{1},\dots k_{m})$ and using that $(N\mu({\hat{B}}_{N,j}))^{l}\leq C(N\mu({\hat{B}}_{N,j}))^{m}$ for $l<m$, we obtain the result. $\square$
+(a) Let $\hat{T}_{N,j}=\sum_{n=1}^{N}{\mathbf{1}}_{{\hat{B}}_{N,j}}\circ f^{n},$ where ${\hat{B}}_{N,j}$} in given by (3.8).
+Since $\frac{{\bar{{\bar{S}}}}_{N,j}}{N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}}\leq\hat{T}_{N,j},$ we have
+${\mathbb{E}}\left(\left(\frac{\bar{\bar{S}}_{N,j}}{N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}}\right)^{m}\right)\leq{\mathbb{E}}\hat{T}_{N,j}^{m}.$
+Next, we consider
+${\mathbb{E}}\left(\binom{\hat{T}_{N,j}}{m}\right)=\sum_{1\leq k_{1}<\cdots<k_{m}\leq N}\mu\left(\bigcap_{j=1}^{m}f^{-k_{j}}{\hat{B}}_{N,j}\right)$
+and prove that
+(B.5) $\sum_{1\leq k_{1}<\cdots<k_{m}\leq N}\mu\left(\bigcap_{j=1}^{m}f^{-k_{j}}{\hat{B}}_{N,j}\right)\leq C\left(N\mu\left(\hat{B}_{N,j}\right)\right)^{m}$
+for non-periodic $x.$ Since $j\leq 0$ and $s<1$, we conclude that (B.5) is less than $\frac{C_{1}}{(\operatorname{ln}N)^{sj\delta m}}$.
+For $y\in{\hat{B}}_{N,j}$ we have $\phi(y)^{m}\left(N\mu\left(\hat{B}_{N,j}\right)\right)^{m}\leq C_{2}N^{m/s}(\operatorname{ln}N)^{\delta m},$ because on this set
+$N^{1/s}(\operatorname{ln}N)^{j\delta}<\phi(y)\leq N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}.$
+Thus combining (B.5) with the identity
+${\mathbb{E}}\hat{T}_{N,j}^{m}=\sum_{k=0}^{m}\left(\sum_{j=0}^{k}(-1)^{k-j}\binom{k}{j}k^{m}\right){\mathbb{E}}\left(\binom{\hat{T}_{N,j}}{k}\right)$,
+we obtain
+${\mathbb{E}}{\bar{{\bar{S}}}}_{N,j}^{m}\leq CN^{m/s}(\operatorname{ln}N)^{\delta m}.$
+We now prove (B.5). We consider several cases depending on the value of $p=\!\!\!\min_{1\leq j\leq m-1}(k_{j+1}\!-\!k_{j}).$
+(i) Since $x$ is non-periodic, for each fixed $p_{0},$ we have
+${\hat{B}}_{N,j}\cap f^{-p}{\hat{B}}_{N,j}=\emptyset\,\,\,\text{for all}\,\,\,p<p_{0}$
+whenever $N$ is sufficiently large because $x$ is non-periodic. Consequently, the contribution of terms corresponding to tuples with $p\leq p_{0}$ vanishes for large $N.$
+(ii) If $p_{0}\leq p\leq\varepsilon\operatorname{ln}N,$ suppose that $\operatorname{Sep}_{N}(k_{1},\dots,k_{m})=l<m.$ Arguing as in case $(M4)_{*}$(ii) of the proof of Proposition 2.3, we conclude that each term is smaller than $C\mu\left(\hat{B}_{N,j}\right)^{l}\theta^{p}$ for fixed $p.$ Since the total number of terms is $O\left(N^{l}p^{m-l}\right),$ it follows that the total contribution of the terms with $p={\hat{p}}$ is bounded above by $CN^{l}{\hat{p}}^{m-l}\mu\left(\hat{B}_{N,j}\right)^{l}\theta^{\hat{p}}$.
+Summing over tuples with ${\hat{p}}\in[p_{0},{\varepsilon}\operatorname{ln}N],$ we obtain a total contribution of
+$CN^{l}{p_{0}}^{m-l}\mu\left(\hat{B}_{N,j}\right)^{l}\theta^{p_{0}}$
+for some constant $C$ independent of $N,$ where $p_{0}$ is sufficiently large.
+(iii) If $\varepsilon\operatorname{ln}N<p\leq K\operatorname{ln}N$ and $\operatorname{Sep}_{N}(k_{1},\dots,k_{m})=l,$ then arguing similarly to case $(M4)_{*}$(iii) we obtain that
+$\mu\left(\bigcap_{j=1}^{m}f^{-k_{j}}{\hat{B}}_{N,j}\right)\leq C\mu\left(\hat{B}_{N,j}\right)^{l+\eta}.$
+Summing over all $r$ tuples and over all $p\in(\varepsilon\operatorname{ln}N,K\operatorname{ln}N],$ we conclude that the total contribution is at most $C\left(N\mu\left(\hat{B}_{N,j}\right)\right)^{l}\mu\left(\hat{B}_{N,j}\right)^{\eta}\left(\operatorname{ln}N\right)^{m-l}.$
+(iv) If $\operatorname{Sep}_{N}(k_{1},\dots,k_{m})=m,$ following the argument in case $(M1)_{r}$ of the proof of Proposition 2.3, we obtain that
+$\mu\left(\bigcap_{j=1}^{m}f^{-k_{j}}{\hat{B}}_{N,j}\right)\leq C\mu\left(\hat{B}_{N,j}\right)^{m}.$
+Summing over all tuples $(k_{1},\dots,k_{m})$ and noting that
+$\left(N\mu\left(\hat{B}_{N,j}\right)\right)^{l}\leq C\left(N\mu\left(\hat{B}_{N,j}\right)\right)^{m}\,\,\,\text{for}\,\,\,l<m,$
+we conclude that (B.5) holds, thus completing the proof of Lemma 3.1(a).
+(b) Suppose $x$ is a periodic point with period $q,$ i.e., $f^{q}(x)=x.$ If there exist $k,\,l\in\mathbb{Z}_{+}$ such that
+$\left\{f^{k}y,f^{k+q}y,\ldots,f^{k+(l-1)q}y\right\}\subset\hat{B}_{N,j}\,\,\,\text{and}\,\,\,f^{k+lq}y\notin\hat{B}_{N,j},$
+then for $p=\operatorname{ln}\operatorname{ln}N$ and sufficiently large $N,$ we have
+$f^{k+lq+i}y\notin\hat{B}_{N,j}\,\,\,\text{for all}\,\,\,i\leq p.$
+Furthermore, if there exists $n>k+lq+p$ such that $f^{n}y\in\hat{B}_{N,j},$ we show that the value $\phi(f^{n}y)$ constrains the sequence $\phi(f^{n+q}y),\ldots,\phi(f^{n+tq}y),$ provided that $f^{n+iq}y\in\hat{B}_{N,j}$ for all $0\leq i\leq t.$ Indeed, since
+$\left|f^{n+q}y-x\right|=\left|f^{n+q}y-f^{q}x\right|\geq\gamma^{q}\left|f^{n}y-x\right|,$
+where $\gamma=\min_{x\in\mathbb{S}}|f^{\prime}(x)|,$ we have
+(B.6) $\phi(f^{n+q}y)\leq\tilde{\gamma}\phi(f^{n}y)$
+for some constant $\tilde{\gamma}$ satisfying $\gamma^{q/s}<\tilde{\gamma}<1.$ Consequently, we obtain $\phi(f^{n+iq}y)\leq\tilde{\gamma}^{i}\phi(f^{n}y)$ for $1\leq i\leq t,$ which implies
+$\sum_{i=0}^{t}\phi(f^{n+iq}y)\leq\left(\sum_{i=0}^{t}\tilde{\gamma}^{i}\right)\phi(f^{n}y)\leq\mathsf{C}\phi(f^{n}y),$
+where $\mathsf{C}=\sum_{i=0}^{\infty}\tilde{\gamma}^{i}=\frac{1}{1-\tilde{\gamma}}$. Now, defining
+$\tilde{B}_{N,j}=\left\{y:fy\notin\hat{B}_{N,j},\cdots,f^{p-1}y\notin\hat{B}_{N,j},f^{p}y\in\hat{B}_{N,j}\right\},$
+we have
+$\sum_{i=0}^{tq}\phi(f^{n+i}y){\mathbf{1}}_{\hat{B}_{N,j}}(f^{n+i}y)=\sum_{i=0}^{t}\phi(f^{n+iq}y)\leq\mathsf{C}\phi(f^{n}y)=\mathsf{C}\phi(f^{n}y){\mathbf{1}}_{\tilde{B}_{N,j}}(f^{n-p}y).$
+Lettting $\tilde{S}_{N,j}(y)=\sum_{k=1}^{N}\phi(f^{k+p}y){\mathbf{1}}_{\tilde{B}_{N,j}}(f^{k}y),$ and summing over the orbit segments that enter $\hat{B}_{N,j},$ we obtain
+(B.7) ${\bar{{\bar{S}}}}_{N,j}(y)\leq\mathsf{C}\tilde{S}_{N,j}(y)+\sum_{i=0}^{k_{1}}\phi(f^{k_{0}+iq}y)$
+where $\{f^{k_{0}+iq}y\}_{i=0}^{k_{1}}$ represents the initial orbit segment entering $\hat{B}_{N,j}.$ Noting that $k_{1}\leq\frac{2\delta s}{-\operatorname{ln}\tilde{\gamma}}\operatorname{ln}\operatorname{ln}N,$ where $\tilde{\gamma}$ satisfies (B.6), it follows that
+$\frac{{\bar{{\bar{S}}}}_{N,j}}{N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}}\leq C\left(\sum_{k=1}^{N}I_{\tilde{B}_{N,j}}(f^{k}y)+\operatorname{ln}\operatorname{ln}N\right).$
+Define
+$\tilde{T}_{N,j}(y)=\sum_{k=1}^{N}{\mathbf{1}}_{\tilde{B}_{N,j}}(f^{k}y),$
+Then
+${\mathbb{E}}\left(\left(\frac{{\bar{{\bar{S}}}}_{N,j}}{N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}}\right)^{m}\right)\leq C^{m}{\mathbb{E}}\left(\left(\tilde{T}_{N,j}+\operatorname{ln}\operatorname{ln}N\right)^{m}\right)=\sum_{r=0}^{m}C^{m}\binom{m}{r}(\operatorname{ln}\operatorname{ln}N)^{m-r}{\mathbb{E}}\left(\tilde{T}_{N,j}\right)^{r}.$
+Next consider the expectation
+(B.8) $\mathbb{E}\left(\binom{\tilde{T}_{N,j}}{r}\right)$ $=\sum_{1\leq k_{1}<\cdots<k_{r}\leq N}\mu\left(\bigcap_{j=1}^{m}f^{-k_{j}}\tilde{B}_{N,j}\right)\leq\sum_{1\leq k_{1}<\cdots<k_{r}\leq N,\atop k_{i+1}-k_{i}\geq p}\mu\left(\bigcap_{j=1}^{m}f^{-k_{j}}\hat{B}_{N,j}\right),$
+where we used that
+$\bigcap_{j=1}^{m}f^{-k_{j}}\hat{B}_{N,j}=\emptyset$
+unless $k_{i+1}-k_{i}\geq p$ for all $0\leq i\leq r-1$. Using the same argument as non-periodic case, the right-hand side of (B.8) is bounded above by $C\left(N\mu(\hat{B}_{N,j})\right)^{r}.$
+Now arguing as in the non-periodic case we obtain
+$\mathbb{E}{\bar{{\bar{S}}}}_{N,j}^{m}\leq CN^{m/s}(\operatorname{ln}N)^{\delta m}(\operatorname{ln}\operatorname{ln}N)^{m},$
+completing the proof for the periodic $x$. ∎